### v1

Abstract
We consider moving boundary problems for biophysics and introduce a new computational framework to handle the complexity of the bulk-surface PDEs. In our framework, interpretability is maintained by adapting the fast, generalizable and accurate structure preservation scheme in [29]. We show that mesh distortion is mitigated by adopting the pioneering work of [36], which is tied to an Arbitrary Lagrangian Eulerian (ALE) framework. We test our algorithms accuracy on moving surfaces with boundary for the following PDEs: advection-diffusion-reaction equations, phase-field models of Cahn-Hilliard type, and Helfrich energy gradient flows. We performed convergence studies for all the schemes introduced to demonstrate accuracy. We use a staggered approach to achieve coupling and further verify the convergence of this coupling using numerical experiments. Finally, we demonstrate broad applicability of our work by simulating state-of-the-art tests of biophysical models that involve membrane deformation.
keywords:
Biophysics, moving boundary problems, bulk-surface coupling, structure preservation, mesh redistribution, multiphysics coupling
{MSCcodes}
35R37, 74K15, 65N30, 76T99.
1 Introduction
Many biophysical phenomena in cell biology can be characterized by the feedback between form and function. Cells routinely change their shape during cell motility, wound healing, and development. Such shape changes are a result of coordination of multiple biochemical and mechanical processes within the cell. Not surprisingly, these phenomena have been studied extensively using models based on partial differential equations (PDEs) with different levels of biophysical complexity [78]. In concert with experiments, such models have provided insight into the mechanisms underlying the change in cell shape [31, 95, 63]. However, an open challenge in the field of computational biophysics is the development of robust computational tools for numerical solutions of the governing PDEs. For example, cellular membranes are mechanical boundaries that are characterized by lateral dimensions that are large compared to their thickness, appearing in a variety of structures in cells [99]. It has been proven that the properties of these surfaces are crucial for the genesis and maintenance of cellular membrane systems [16, 86]. These restructuring phenomena strongly depend on the components of the membranes themselves and their surroundings [19, 10, 96, 17, 73, 104, 2, 5, 98], leading to PDE systems coupling complex multiphysics to free surfaces problems. The resulting systems typically exhibit highly non-linear dynamics and a general numerical framework to simulate these rich dynamics in a unified and stable setting is still missing. The goal of this article is to present a finite element based computational framework which addresses some of the key challenges arising in continuum-mechanics based computational models of cells.
1.1 State of the art
Numerous numerical methods have been proposed to deal with the challenges posed by a moving deformable membrane [61, 92, 1, 71, 93, 69, 90, 85, 7], including: immersed boundary methods, level-set method, mesh-free methods, particle methods, and the phase-field method. Despite numerous successes, even the most advanced schemes seem to lack the following features:
  1. P.1
    It is not rare, in the targeted applications, to incur advection-dominant (AD) equations possibly posed on a moving surface. Characteristic examples can be the transport of surfactants in two-phase flow or the interaction between actin’s barbed ends and the cellular membrane. There exists a large bulk of literature for parabolic advection-diffusion-reaction (ADR) PDEs on moving surfaces [38, 40, 18, 80, 39, 67, 68, 42, 25]. For what concerns AD equations, the bulk setting has been extensively explored and the surface case has seen recent developments  [91, 22, 97, 23, 33, 100, 101]. Surprisingly, most aforementioned approaches only consider the case of stationary surfaces, leaving the AD case largely unaddressed. A notable exception is the work in [53], where characteristic cut finite element methods on moving surfaces is proposed.
  2. P.2
    Physical structure preservation (SP) is still overlooked by the majority of the available articles. It is crucial, in order to maintain physical interpretability as well as numerical stability, to be able to guarantee constraints such as: mass and density positivity, phase bounds, energy decay, mass preservation, etc. A vast and diversified literature exist dedicated to preserve bounds, see [79, 74, 34, 75, 103, 102, 76, 28, 58, 56] and references therein. Unfortunately, this richness of options is lacking if we consider PDEs on surfaces, let alone moving surfaces.
  3. P.3
    Most articles assume idealized geometries, in which the complexity is still highly reduced with respect to realistic geometries resulting from imaging techniques [72]. This is especially crucial in this context since we expect the domain evolution to be dependent on the coupled equations and not known a priori. Steps are continuously being taken in this direction for what concerns geometric flows, which notoriously drive meshes towards bad conditioning, see [11, 81, 57, 36, 9, 51]. However, even in these cases, one is quite far from the structures considered in more applied articles such as [50].
The present work aims to propose a first step towards a cure of the missing pieces described above. It is meant to bridge the complexity gap between numerical accuracy and biophysical complexity in moving domains scenario. While there exist a plethora of numerical methods which each target a specific problem class or particular issue, we here employ and extend some of the most recent state-of-the-art methods and combine them into a holistic simulation framework which can handle coupled multiphysics and shape deformation problems arising in computational cell biology.
1.2 New contributions and outline
In Section 2, we provide the notation necessary for the formulation and discretization of PDEs posed on evolving domains. For purposes of generality, we keep the notation the same in the bulk and on the surface, only distinguishing between the two as needed. In Section 3, we adapt the structure-preserving scheme in [30] to moving bulk and surface problems as a solution to P.2. The model-agnostic, point-wise approach of [30] is fit to the first-order, piecewise-linear finite element framework here used. The result is a fast, generalizable and flexible scheme that can be readily adapted to multiple PDEs posed on moving bulk and surfaces. In this context, we use it to bound the PDE solution to physically-relevant values while at the same time preserving the total mass. In Section 4, we introduce a mesh redistribution algorithm based on the surface tangential motion of [36] and an Arbitrary Lagrangian Eulerian (ALE) approach. The two-step scheme in [36] is decoupled from the gradient flow dynamics and applied to general motions. This allows us to greatly alleviate the need of remeshing when simulating domains characterized by large deformations. At the same time, the scheme does not interfere with the shape evolution, maintaining accuracy. Both schemes in Section 3 and Section 4 take the form of a post-processing step and thus have the advantage of being a tunable feature of the code. In Section 5 we augment an ADR scheme for moving surfaces with a continuous interior penalty (CIP) stabilization to handle AD cases and solve P.1. The stabilization is proven to be convergent and so is its coupling with the structure-preserving algorithm of Section 3. The flexibility of the setting is further verified with an additional real-case example. In Section 6 we apply structure preservation to a Cahn-Hilliard type equation on moving surfaces. Numerical tests are presented that highlight how the algorithm in Section 3 is readily applicable to moving surfaces and true to the underlying physics. In Section 7, we focus on P.3, applying the mesh redistribution scheme of Section 4 to the parametric method of Barrett, Garke and Nürnberg [14] for Helfrich flow. The postprocessing step is extensively tested with state-of-the-art examples revealing good convergence properties and long-term stability. In Section 8, we couple the solvers in Section 5 through 7 together in a staggered approach. Convergence is tested with manufactured solutions [67] where the ADR algorithm and a simplified version of the Helfrich algorithm are coupled. In the same context a PDE system modeling tumor growth is also reproduced. The applicability of our approach to relevant cell biophysics is further explored simulating phase separation on a deformable membrane, where the Cahn-Hilliard and Helfrich schemes are coupled.
2 Model
This article explores numerical approaches to biophysical cell phenomena where the dynamics involve moving boundaries, such as cell membranes, that are assumed to have negligible thickness and thus are treated as moving hypersurfaces. As a result, we begin by outlining the essential notation required for both formulating and discretizing partial differential equations on evolving domains, with specific attention given to surface-bound PDEs [39, 15].
Let $M\subset\mathbb{R}^{d},\;d=2,3$, be an $m$-dimensional oriented, compact, $C^{2}$ manifold with boundary $\partial M$. The notation $\Omega\equiv M$ when $\dim{(M)}=d$ and $\Gamma\equiv M$ when $\dim{(M)}=d-1$ will also be used. In the case $\dim(M)=d-1$, we denote the tangent space with respect to ${\boldsymbol{p}}\in M$ with $T_{{\boldsymbol{p}}}(M)$ and the associated tangent bundle with $TM$. The unit normal vector ${\boldsymbol{n}}_{M}({\boldsymbol{p}})$ to $M$ at $\boldsymbol{p}$} is defined as the vector ${\boldsymbol{n}}_{M}({\boldsymbol{p}})\in\mathbb{R}^{d}$ such that ${\boldsymbol{n}}_{M}({\boldsymbol{p}})\perp T_{{\boldsymbol{p}}}(M)$, $\lVert{\boldsymbol{n}}_{M}({\boldsymbol{p}})\rVert=1$ and ${\boldsymbol{n}}_{M}({\boldsymbol{p}})$ agrees with the orientation given. The norm $\lVert\cdot\rVert$ is the standard Euclidean norm in $\mathbb{R}^{d}$. For $\dim{(M)}=d-1$ we define the tangential projection at $\boldsymbol{p}$} as
(1) $\mathbb{P}_{M}({\boldsymbol{p}})=\mathbb{I}-{\boldsymbol{n}}_{M}({\boldsymbol{p}})\otimes{\boldsymbol{n}}_{M}({\boldsymbol{p}}),$
where $\mathbb{I}$ is the identity matrix. For $\dim{(M)}=d$, we set $\mathbb{P}_{M}({\boldsymbol{p}})=\mathbb{I}$ and ${\boldsymbol{n}}_{M}({\boldsymbol{p}})=\mathbf{0}$. This allows us to define the tangential gradient of a differentiable function $f:M\to\mathbb{R}$ at $\boldsymbol{p}$} as
(2) $\nabla_{M}f({\boldsymbol{p}})=\nabla\bar{f}({\boldsymbol{p}})-\mathbb{P}_{M}({\boldsymbol{p}})\nabla\bar{f}({\boldsymbol{p}}),$
where $\bar{f}$ is a smooth extension of $f$ to a $d$-dimensional neighborhood of $M$ such that $\bar{f}\rvert_{M}=f$ and $\nabla$ is the Euclidean gradient in $\mathbb{R}^{d}$. Analogously, the tangential divergence is defined as $\nabla_{M}\cdot f({\boldsymbol{p}})=\mathrm{Tr}(\nabla_{M}f({\boldsymbol{p}}))$. This leads to the definition of the Laplace-Beltrami operator for a $C^{2}$-function $f:M\to\mathbb{R}$ as
(3) $\Delta_{M}f({\boldsymbol{p}})=\nabla_{M}\cdot\nabla_{M}f({\boldsymbol{p}}).$
For ${\boldsymbol{f}}:M\to\mathbb{R}^{d}$ and $\mathbb{F}:M\to\mathbb{R}^{d\times d}$ we define $(\nabla_{M}{\boldsymbol{f}})_{ij}=(\nabla({\boldsymbol{f}}\cdot{\boldsymbol{e}}_{i}))_{j}$ and $(\nabla_{M}\cdot\mathbb{F})_{i}=\nabla_{M}\cdot(\mathbb{F}^{T}{\boldsymbol{e}}_{i})=\mathrm{Tr}(\nabla_{M}(\mathbb{F}^{T}{\boldsymbol{e}}_{i}))$, where ${\boldsymbol{e}}_{i},~i\in\{1,\ldots,d\}$ is the canonical basis in $\mathbb{R}^{d}$. This allows us to define $\Delta_{M}{\boldsymbol{f}}=\nabla_{M}\cdot\nabla_{M}{\boldsymbol{f}}$. In order to analyze curvature properties of $M$, we introduce the extended Weingarten map $\mathbb{H}$ as
(4) $\mathbb{H}({\boldsymbol{p}})=-\nabla_{M}{\boldsymbol{n}}_{M}({\boldsymbol{p}}).$
It can be shown that $\mathbb{H}$ is symmetric, has an eigenvalue 0 in the direction of the normal, and restricts to the Weingarten map $\mathbb{W}({\boldsymbol{p}})$ on the tangent space $T_{{\boldsymbol{p}}}M$. For ${\boldsymbol{p}}\in M$, the mean curvature is defined as
(5) $\kappa({\boldsymbol{p}})=\mathrm{Tr}(\mathbb{H}({\boldsymbol{p}})).$
We recall that all of the above definitions are independent of the extension $\bar{f}$. A sketch of possible model setups is shown in Figure 1.
[…]
Figure 1: Sketch of an open and closed geometry for notation purposes. One is a 3D object and the other is a 2D object.
We consider a time interval J=[0,T],T>0J=[0,T],\;T>0 and a $C^{2}$-evolving manifold $\{M(t)\}_{t\in J}$ in $\mathbb{R}^{d}$. In this article an evolving manifold is modeled by a reference manifold $\widehat{M}$ together with a flow map
(6) ${\boldsymbol{\Phi}}:J\times\widehat{M}\to\mathbb{R}^{d},\quad{\boldsymbol{\Phi}},{\boldsymbol{\Phi}}^{-1}\in C^{1}(J;C^{2}(\mathbb{R}^{d};\mathbb{R}^{d})),$
such that
[…]
  • $^{\boldsymbol{\Phi}}_{0}=\mathbf{Id}_{\widehat{M}}$, where $\mathbf{Id}_{\widehat{M}}$ is the identity map on the reference manifold.
The mapping ${\boldsymbol{\Phi}}_{t}$ can be used to define pull-back and push-forward maps of functions [4, 3]
(7) ${\boldsymbol{\Phi}}_{-t}f=f\circ{\boldsymbol{\Phi}}_{t}:\widehat{M}\rightarrow\mathbb{R}\text{ for }f:M(t)\rightarrow\mathbb{R},$
${\boldsymbol{\Phi}}_{t}f=f\circ{\boldsymbol{\Phi}}_{t}^{-1}:M(t)\rightarrow\mathbb{R}\text{ for }f:\widehat{M}\rightarrow\mathbb{R}$.
Often we will call the space-time set $\mathcal{G}_{T}=\cup_{t\in J}(\{t\}\times M(t))$ a $C^{2}$-evolving manifold and identify it with $\{M(t)\}_{t\in J}$. In addition, we assume that there exists a velocity field ${\boldsymbol{v}}:J\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ with ${\boldsymbol{v}}\in C^{0}(J;C^{2}(\mathbb{R}^{d};\mathbb{R}^{d}))$ such that for any $t\in J$ and every $^{\boldsymbol{p}}\in\widehat{M}$
(8) $\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}{\boldsymbol{\Phi}}_{t}({\boldsymbol{p}})={\boldsymbol{v}}(t,{\boldsymbol{\Phi}}_{t}({\boldsymbol{p}})).$
For ${\boldsymbol{x}}\in M(t)$, the velocity $\boldsymbol{v}$} can be split into a tangential component ${\boldsymbol{v}}^{\top}(t,{\boldsymbol{x}})\in T_{{\boldsymbol{x}}}M(t)$ and a normal one ${\boldsymbol{v}}^{\perp}(t,{\boldsymbol{x}})={\boldsymbol{v}}(t,{\boldsymbol{x}})-{\boldsymbol{v}}^{\top}(t,{\boldsymbol{x}})$.
This allows us to define a normal derivative of a scalar function $f$ on $M(t)$ as
(9) $\partial^{\circ}f(t,{\boldsymbol{x}})=\partial_{t}\bar{f}(t,{\boldsymbol{x}})+{\boldsymbol{v}}^{\perp}(t,{\boldsymbol{x}})\cdot\nabla\bar{f}(t,{\boldsymbol{x}}),$
where $\bar{f}$ is a smooth extension of $f$ in a space-time neighborhood of $M(t)$. The material derivative is defined as
(10) $\partial^{\bullet}f(t,{\boldsymbol{x}})=[\partial^{\circ}f+{\boldsymbol{v}}^{\top}\cdot\nabla_{M(t)}\bar{f}](t,{\boldsymbol{x}})=[\partial_{t}\bar{f}+{\boldsymbol{v}}\cdot\nabla\bar{f}](t,{\boldsymbol{x}}).$
The definition of a flow map is not unique and the same evolving manifold can be described by different maps. The computational domain’s flow map can be chosen so to maintain good mesh properties while keeping the original map $\boldsymbol{\Phi}$} for the PDE evolution. This is achieved by introducing a second flow map ${\boldsymbol{\Phi}}^{\mathcal{A}}$ called Arbitrary Lagrangian Eulerian (ALE) map with corresponding velocity ${\boldsymbol{v}}^{\mathcal{A}}$ and ALE material derivative $\partial^{\mathcal{A}}$. It can be shown that ${\boldsymbol{v}}^{\mathcal{A}}(t,{\boldsymbol{x}})-{\boldsymbol{v}}(t,{\boldsymbol{x}})\in T_{{\boldsymbol{x}}}M(t)$ and that
(11) $\partial^{\mathcal{A}}f(t,{\boldsymbol{x}})=[\partial^{\bullet}f+({\boldsymbol{v}}^{\mathcal{A}}-{\boldsymbol{v}})\nabla_{M(t)}f](t,{\boldsymbol{x}}).$
It is worth noting that the normal velocity $\boldsymbol{v}$$^{\perp}$ is in any case uniquely determined, i.e. ${\boldsymbol{v}}^{\perp}(t,\cdot)={\boldsymbol{v}}^{\mathcal{A},\perp}(t,\cdot),\;\forall t\in J$, including the normal velocity of the boundary $\partial M(t)$.
2.1 Space discretization
The geometries are discretized using piecewise linear elements, where for surfaces we follow the setup of [39]. An $m$-dimensional manifold $M$ is approximated by a triangulated $m$-dimensional domain denoted by $M_{h}$. The elements of the discretization will be distinguished based on their codimension. We suppose $M_{h}$ is composed by a collection $\mathcal{T}_{h}$ of $m$-simplices which vertices ${\boldsymbol{y}}_{i},\;i=1,\ldots,N$ lie on $M$ and such that
(12) $\mathcal{T}_{h}=\{T\},\quad M_{h}=\bigcup T.$
For each $T\in\mathcal{T}_{h}$ we denote by $h(T)$ its diameter and define the mesh-size as
(13) $h=\max_{T\in\mathcal{T}_{h}}h(T).$
The boundary of every element $T$ is composed by $m+1$ facets of dimension m1m-1, forming a collection $\mathcal{F}=\{F\}$. The discrete boundary of $M_{h}$ is defined as the union of those facets which defining points lie on $\partial M$ and is denoted $\partial M_{h}$. It is useful to distinguish between internal and boundary facets, defined as
$\mathcal{F}_{h}^{i}$ $=\{F\in\mathcal{F}:\;F\cap\partial M_{h}=\emptyset\},$
$\mathcal{F}_{h}^{b}$ $=\{F\in\mathcal{F}:\;F\cap\partial M_{h}=F\}.$
Every element $T$ has a (constant) normal which we will denote by ${\boldsymbol{n}}_{T}$. In the case $\mathrm{dim}(M_{h})\equiv d$ we define ${\boldsymbol{n}}_{T}\equiv 0$. In the case $\mathrm{dim}(M_{h})\equiv d-1$ we define a discrete projection $\mathbb{P}_{T}=\mathbb{I}-{\boldsymbol{n}}_{T}\otimes{\boldsymbol{n}}_{T}$ and relative tangential gradient, understood in an element-wise sense. The facet normal, which we denote ${\boldsymbol{n}}_{F}$, is well-defined once we consider it on the boundary of an element $T$. For this reason for each $F\in\partial T$, ${\boldsymbol{n}}_{F}$ is the unit co-normal vector perpendicular to both ${\boldsymbol{n}}_{T}$ and $F$ and directed outward with respect to the element $T$. We will also use the notation ${\boldsymbol{n}}_{F}^{\pm}$, where given as $T^{+}$ and $T^{-}$ the neighboring elements to $F$, then ${\boldsymbol{n}}_{F}^{+}$ is the normal as seen from $T^{+}$ and ${\boldsymbol{n}}_{F}^{-}$ is the normal as seen from $T^{-}$. In the bulk case we simply have that ${\boldsymbol{n}}_{F}^{-}=-{\boldsymbol{n}}_{F}^{+}$, instead for surface cases the two normals are not co-planar. An illustration of these objects is given in Figure 2.
${\boldsymbol{n}}_{T^{+}}$$T^{+}$${\boldsymbol{n}}_{T^{-}}$$T^{-}$${\boldsymbol{n}}_{F^{+}}$${\boldsymbol{n}}_{F^{-}}$
Figure 2: Sketch of an discrete discrete mesh entities of a discrete surface $M_{h}$.
We will again use the notation $\Omega_{h}$ and $\Gamma_{h}$ when we need to distinguish between $d$-dimensional and (d1)(d-1)-dimensional discrete manifolds, respectively.
Throughout this work, we use a first-order fitted finite element method to discretize the PDEs in space [46]. In particular, for surfaces, we employ the surface finite element method (SFEM) as presented in [39]. The corresponding finite element space is defined as
(14) $V_{h}(M_{h}):=\{v_{h}\in C(M_{h})\;|\;v_{h}\rvert_{T}\text{ is linear for all }T\in\mathcal{T}_{h}\},$
which can be described as the span of $N$ piecewise linear continuous basis functions $\phi_{i},\;i=1,\ldots,N$ such that $\phi_{i}({\boldsymbol{y}}_{j})=\delta_{ij}$. Occasionally, we will also need the space
(15) $V_{h0}(M_{h}):=\{v_{h}\in C(M_{h})\;|\;v_{h}\rvert_{T}\text{ is linear for all }T\in\mathcal{T}_{h}\text{ and }v_{h}|_{\partial M_{h}}=0\}.$
Integration over mesh entities follows the same notation as for the continuous one, where the summation over every element of the collection is implicit, i.e. we write
(16) $(\cdot\,,\,\cdot)_{M_{h}}=\sum_{T\in M_{h}}(\cdot\,,\,\cdot)_{T},\quad(\cdot\,,\,\cdot)_{\mathcal{F}_{h}^{i}}=\sum_{F\in\mathcal{F}_{h}^{i}}(\cdot\,,\,\cdot)_{F},$
where (,)($\cdot\,$$\cdot$) is a certain inner product. The same element-wise summation is used in case of discrete norms.
2.2 Time discretization
To discretize the evolving manifold $M(t)$ we use the reference discretization $\widehat{M}_{h}$ and transport it using a suitable ALE map ${\boldsymbol{\Phi}}^{\mathcal{A}}$ [55, 60, 89, 52, 49, 8, 18, 68, 3]. The reference points $\widehat{{\boldsymbol{y}}}_{i}\in\widehat{M}_{h}$ defining the simplices of the reference triangulation $\mathcal{T}_{h}$ are evolved as ${\boldsymbol{y}}_{i}(t)={\boldsymbol{\Phi}}_{t}^{\mathcal{A}}(\widehat{{\boldsymbol{y}}}_{i})$. This allows to define evolved elements $T(t)$ and facets $F(t)$ such that
(17) $\mathcal{T}_{h}(t)=\{T(t)\},\quad\mathcal{F}_{h}(t)=\{F(t)\},\quad M_{h}(t)=\bigcup T(t),$
$\mathcal{F}_{h}^{i}(t)=\{F(t)\in\mathcal{F}_{h}(t):\;F(t)\cap\partial M_{h}(t)=\emptyset\},$
$\mathcal{F}_{h}^{b}(t)=\{F(t)\in\mathcal{F}_{h}(t):\;F(t)\cap\partial M_{h}(t)=F(t)\}.$
For the discretization of the function spaces, we start with the finite element space defined in (14), then transport the basis functions using the flow map $\phi_{i}(t)={\boldsymbol{\Phi}}_{t}^{\mathcal{A}}(\phi_{i})$ and define the evolved finite element space as their span
(18) $V_{h}(t)=\{\mathrm{span}\{\phi_{i}(t)\}:i=1,\ldots,N.\;\phi_{i}\in V_{h}(\widehat{M}_{h})\}.$
In an analogous way we can also define $V_{h0}(t)$. It follows from the definitions that $V_{h}(0)\equiv V_{h}(\widehat{M}_{h})$. Discrete material velocities are defined by
(19) ${\boldsymbol{v}}_{h}(t,\cdot)=\sum_{i=1}^{N}{\boldsymbol{v}}(t,{\boldsymbol{y}}_{i}(t))\,\phi_{i}(t),\quad{\boldsymbol{v}}^{\mathcal{A}}_{h}(t,\cdot)=\sum_{i=1}^{N}{\boldsymbol{v}}^{\mathcal{A}}(t,{\boldsymbol{y}}_{i}(t))\,\phi_{i}(t),$
with corresponding derivatives
(20) $\partial^{\bullet}_{h}\phi_{i}=\partial_{t}\bar{\phi_{i}}+{\boldsymbol{v}}_{h}\cdot\nabla\bar{\phi_{i}},\quad\partial^{\mathcal{A}}_{h}\phi_{i}=\partial_{t}\bar{\phi_{i}}+{\boldsymbol{v}}^{\mathcal{A}}_{h}\cdot\nabla\bar{\phi_{i}}.$
Using an ALE map, which describes the motion of the computational domain, we have that $\partial^{\mathcal{A}}_{h}\phi_{i}=0$.
In this work, time-dependent PDEs are discretized in a collection of time points $\{t_{0},\ldots,t_{N}\}$. We denote quantities on discrete time points by a superscript, where $(\cdot)^{n}$ refers to the previous time step and $(\cdot)^{n+1}$ refers to the current time step. The time step size is defined as $\tau^{n+1}=t_{n+1}-t_{n}$ and for the sake of simplicity we restrict in what follows to constant time step size $\tau$, with the understanding that every first-order-in-time scheme is amenable to adaptive time stepping.
Remark 2.1.
From now on, we will often omit the explicit dependence on time, indicating for example $M_{h}(t)$ as simply $M_{h}$ and leaving it to the context to indicate the considered time.
3 Structure preservation
In this section we describe how to deal with the challenges highlighted in P.2. In the effort of maintaining the interpretability of the considered systems, we focus on bounds and mass preservation. The method recently proposed in [30, 29] provides a flexible approach to achieve such goals for a wide variety of finite element formulations on stationary bulk domains. Here, we extend those advances to moving surfaces and moving bulk domains. The main advantages of the employed method are:
  • Bounds preservation is achieved in such a way that, in its simplest form, it reduces to the widely used cut-off approach.
  • Accuracy is theoretically retained for higher order methods in space and time, granted some underlying hypotheses are satisfied. This is not the case, for example, for discrete maximum principle preserving schemes.
  • The approach can easily be incorporated as a post-processing step without requiring explicit time integration. While explicit time integration as reviewed in [103] can be effective for hyperbolic equations, here we are also interested in parabolic equations.
  • The simplicity of the algorithm makes it suitable for implementation in legacy codes with negligible overhead computational time with respect to the original non-preserving discretization method. Nevertheless, the method enjoys good stability properties and even allows for error analysis for some specific cases, see [29].
  • No ad hoc problem reformulation is needed, but the scheme only acts on the nodal values of the solution through a Lagrange multiplier approach.
Following the presentation in [29], the derivation of the method starts from a nonlinear PDE in the form
(21) $u_{t}+\mathcal{L}u+\mathcal{N}(u)=0,$
where $\mathcal{L}$ is a linear or non-negative operator and $\mathcal{N}(u)$ is a semilinear or quasi-linear operator. The operators $\mathcal{L}$ and $\mathcal{N}(u)$ are acting on functions defined on the space-time manifold $\mathcal{G}_{T}$. As underlying assumption, the solution to the continuum problem lies in the interval $[a,b]$, i.e. given $a\leq u(0,{\boldsymbol{p}})\leq b$ for all $^{\boldsymbol{p}}\in\widehat{M}$ then $a\leq u(t,{\boldsymbol{x}})\leq b$ for all $(t,{\boldsymbol{x}})\in\mathcal{G}_{T}$. The corresponding generic spatial discretization is
(22) $\partial_{t}u_{h}+\mathcal{L}_{h}u_{h}+\mathcal{N}_{h}(u_{h})=0,$
where we assume $u_{h}\in V_{h}$. To make the scheme bound-preserving, a Lagrange multiplier $\lambda_{h}$ is introduced together with the quadratic function g(u)=(bu)(ua)g(u)=(b-u)(u-a). The problem is then reformulated as
(23a) $\partial_{t}u_{h}+\mathcal{L}_{h}u_{h}+\mathcal{N}_{h}(u_{h})=\lambda_{h}g^{\prime}(u_{h}),$
(23b) $\lambda_{h}\geq 0,\quad g(u_{h})\geq 0,\quad\lambda_{h}g(u_{h})=0,$
where the second line represents the usual Karush-Kuhn-Tucker conditions for constrained optimization [62, 70]. A core assumption of the scheme presented in [29] is that (23) is satisfied pointwise in a set of points $\Sigma_{h}$ that can be both mesh points or collocation points for $M_{h}$. We must then have:
(24a) $\partial_{t}u_{h}({\boldsymbol{p}})+\mathcal{L}_{h}u_{h}({\boldsymbol{p}})+\mathcal{N}_{h}(u_{h}({\boldsymbol{p}}))=\lambda_{h}({\boldsymbol{p}})g^{\prime}(u_{h}({\boldsymbol{p}})),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$
(24b) $\lambda_{h}({\boldsymbol{p}})\geq 0,\quad g(u_{h}({\boldsymbol{p}}))\geq 0,\quad\lambda_{h}({\boldsymbol{p}})g(u_{h}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h}.$
It is important to notice that $\Sigma_{h}$ does not include points where essential boundary conditions are applied. The next step is to discretize (24) in time, where we here focus on the case of Backward-Euler (BE or BDF-1) time integration in the interval $\{t^{n},t^{n+1}\}$. Higher order extensions are presented in [29] and can be adapted following similar arguments. The key idea is to apply an operator splitting approach to (23) dividing the problem in two steps.
In the predictor step we solve the unconstrained problem
(25) $\frac{\tilde{u}_{h}^{n+1}({\boldsymbol{p}})-u_{h}^{n}({\boldsymbol{p}})}{\tau}+\mathcal{L}_{h}\tilde{u}_{h}^{n+1}({\boldsymbol{p}})+\mathcal{N}_{h}(\tilde{u}_{h}^{n+1}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h}.$
In the corrector step we solve the constrained problem
(26a) $\frac{u_{h}^{n+1}({\boldsymbol{p}})-\tilde{u}_{h}^{n+1}({\boldsymbol{p}})}{\tau}=\lambda_{h}^{n+1}({\boldsymbol{p}})g^{\prime}(u_{h}^{n+1}({\boldsymbol{p}})),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$
(26b) $\lambda_{h}^{n+1}({\boldsymbol{p}})\geq 0,\quad g(u_{h}^{n+1}({\boldsymbol{p}}))\geq 0,\quad\lambda_{h}^{n+1}({\boldsymbol{p}})g(u_{h}^{n+1}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h}.$
If we are only interested in bound preservation, it turns out that the solution to (26) is the cutoff function
(27) $u_{h}^{n+1}=\text{cutoff}[\tilde{u}_{h}^{n+1}]=\begin{cases}\tilde{u}_{h}^{n+1}({\boldsymbol{p}}),&\text{ if }a<\tilde{u}_{h}^{n+1}({\boldsymbol{p}})<b,\\ a,&\text{ if }\tilde{u}_{h}^{n+1}({\boldsymbol{p}})<a,\\ b,&\text{ if }\tilde{u}_{h}^{n+1}({\boldsymbol{p}})>b.\end{cases}$
The above step is unfortunately not mass-preserving even if the solution is. To fix this, the authors in [30, 29] introduce an additional global space-independent Lagrange multiplier $\xi_{h}^{n+1}$. The corrector step (26) is modified to include the mass preservation constraint,
(28a) $\frac{1}{\tau}u_{h}^{n+1}({\boldsymbol{p}})=\lambda_{h}^{n+1}({\boldsymbol{p}})g^{\prime}(u_{h}^{n+1}({\boldsymbol{p}}))+\xi_{h}^{n+1}+\frac{1}{\tau}\tilde{u}_{h}^{n+1}({\boldsymbol{p}}),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$
(28b) $\lambda_{h}^{n+1}({\boldsymbol{p}})\geq 0,\quad g(u_{h}^{n+1}({\boldsymbol{p}}))\geq 0,\quad\lambda_{h}^{n+1}({\boldsymbol{p}})g(u_{h}^{n+1}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h},$
[…]
where $(\cdot\,,\,\cdot)^{h}$ is a suitably chosen discrete inner product. In our case this discrete inner product, that has to be expressed pointwise as $(u\,,\,v)^{h}=\sum_{{\boldsymbol{p}}\in\Sigma_{h}}\omega_{\boldsymbol{p}}u({\boldsymbol{p}})v({\boldsymbol{p}})$, is assumed to be equivalent to the classical $L^{2}$ inner product on $V_{h}$. For this reason, whenever required, we will choose $(\cdot\,,\,\cdot)^{h}$ to be the mass-lumped inner product. Rearranging (28a) as
(29) $\frac{1}{\tau}(u_{h}^{n+1}({\boldsymbol{p}})-(\tilde{u}_{h}^{n+1}({\boldsymbol{p}})+\tau\xi_{h}^{n+1}))=\lambda_{h}^{n+1}({\boldsymbol{p}})g^{\prime}(u_{h}^{n+1}({\boldsymbol{p}})),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$
the solution is then $u_{h}^{n+1}=\text{cutoff}[\tilde{u}_{h}^{n+1}+\tau\xi_{h}^{n+1}]$ where $\xi_{h}^{n+1}$ is the solution to the nonlinear equation
(30) $F^{n+1}(\xi_{h}^{n+1})=(\text{cutoff}[\tilde{u}_{h}^{n+1}+\tau\xi_{h}^{n+1}]\,,\,1)^{h}-(u_{h}^{n}\,,\,1)^{h}=0.$
Bounds on the energy norm are provided for a class of non-linear PDEs in [29] together with error analysis of a restricted class of problems.
Remark 3.1.
Since $(F^{n+1})^{\prime}$ might not exist, the authors of the scheme in [30, 29] suggest the use of a secant method to find the solution of $F^{n+1}(\xi)=0$ in (30) using
(31) $\xi_{k+1}=\xi_{k}-\frac{F^{n+1}(\xi_{k})(\xi_{k}-\xi_{k-1})}{F^{n+1}(\xi_{k})-F^{n+1}(\xi_{k-1})},$
together with initial guesses $\xi_{0}=0$ and $\xi_{1}=\mathcal{O}(\tau)$. From a practical standpoint, the algorithm converged in less than four iterations for each timestep for the case at hand.
The technique introduced is clearly directly applicable to a wide variety of equations. The authors themselves in [30, 29] have proposed experiments for Allen-Cahn, Cahn-Hilliard with variable mobility and Fokker-Planck equations on a fixed bulk domain. Inspired by the biophysical mechanisms driving cell reshaping, we will demonstrate their accuracy when modeling complex advection-diffusion-reaction equations and phase-separation phenomena. The structure preservation properties will allow to simultaneously maintain stability, interpretability and physical complexity.
4 Tangential grid control and ALE discretization
When dealing with evolving domains in a fitted framework, it is not uncommon to incur deformations that lead to highly distorted meshes. This is especially true when gradient flows are involved, since they only prescribe the velocity and hence the displacement in normal direction. Following 11, we introduce a particular ALE map tailored for gradient flow dynamics that works as follows.
Consider the time interval $\{t_{n},t_{n+1}\}$, and suppose the surface $\Gamma_{h}^{n}$ would be displaced, following the physics of the problem, to an updated surface $\widetilde{\Gamma}_{h}^{n+1}$ at time $t_{n+1}$ which is highly distorted. We define an artificial tangential motion aimed at maintaining the surface shape (and consequently energy) but redistributing the nodes more favorably to reduce mesh distortions. We choose the two-stage algorithm proposed in [36] for the following reasons:
  • The method has proven to be very effective for gradient flows such as the mean curvature flow and the surface diffusion flow.
  • Its two-stage nature makes it a tunable feature of the code, that can be turned on and off at will.
  • Being a two-stage process and not embedded in the gradient-flow algorithm, it can be applied on the mesh evolution independently of the presence of the gradient flow.
  • Beyond the scope of the current work, the method can potentially be applied to higher order in time integration schemes.
The algorithm presented in [36] is based on the assumption that the starting mesh is regular enough. The idea is then to impose an artificial tangential velocity that requires ${\boldsymbol{\Phi}}^{\mathcal{A}}(t,\cdot):\widehat{\Gamma}\to\Gamma$ to be an harmonic map, i.e.
(32) $\begin{cases}\frac{\partial{\boldsymbol{\Phi}}^{\mathcal{A}}}{\partial t}\cdot({\boldsymbol{n}}_{\Gamma}\circ{\boldsymbol{\Phi}}^{\mathcal{A}})=({\boldsymbol{v}}\cdot{\boldsymbol{n}})\circ{\boldsymbol{\Phi}}^{\mathcal{A}},\\ -\Delta_{\,\widehat{\Gamma}}{\boldsymbol{\Phi}}^{\mathcal{A}}=(\varkappa{\boldsymbol{n}})\circ{\boldsymbol{\Phi}}^{\mathcal{A}}.\end{cases}$
Equivalently, ${\boldsymbol{\Phi}}^{\mathcal{A}}$ is required to minimize the energy $\int_{\,\widehat{\Gamma}}|\nabla_{\widehat{M}}{\boldsymbol{\Phi}}^{\mathcal{A}}(t,\cdot)|^{2}$, under the constraint ${\boldsymbol{v}}\cdot{\boldsymbol{n}}_{\Gamma}=0$ on $\Gamma$, which is imposed using the scalar-valued Lagrange multiplier $\varkappa$. The fully discrete version of the scheme reads: given a surface $\widetilde{\Gamma}_{h}^{n+1}$ evolved from an initial surface $\widehat{\Gamma}_{h}$, find ${\boldsymbol{w}}_{h}^{n+1}\in[V_{h0}(\widetilde{\Gamma}^{n+1}_{h})]^{d}$ and $\varkappa_{h}^{n+1}\in V_{h0}(\widetilde{\Gamma}^{n+1}_{h})$ such that
$(\nabla_{\widehat{\Gamma}_{h}}({\boldsymbol{w}}_{h}^{n+1}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{t_{n+1}})\,,\,\nabla_{\widehat{\Gamma}_{h}}({\boldsymbol{\phi}}_{h}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{t_{n+1}}))_{\widehat{\Gamma}_{h}}-(\varkappa_{h}^{n+1}{\boldsymbol{n}}_{\widetilde{\Gamma}_{h}^{n+1}}\,,\,{\boldsymbol{\phi}}_{h})_{\widetilde{\Gamma}_{h}^{n+1}}$
(33a) $\qquad=-(\nabla_{\widehat{\Gamma}_{h}}(\mathbf{Id}_{\widetilde{\Gamma}_{h}^{n+1}}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{t_{n+1}})\,,\,\nabla_{\widehat{\Gamma}_{h}}({\boldsymbol{\phi}}_{h}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{t_{n+1}}))_{\widehat{\Gamma}_{h}}$
(33b) $-({\boldsymbol{w}}_{h}^{n+1}\cdot{\boldsymbol{n}}_{\widetilde{\Gamma}_{h}^{n+1}}\,,\,\psi_{h})_{\widetilde{\Gamma}_{h}^{n+1}}=0,$
for all ${\boldsymbol{\phi}}_{h}\in[V_{h0}(\widetilde{\Gamma}^{n+1}_{h})]^{d}$ and $\psi_{h}\in V_{h0}(\widetilde{\Gamma}^{n+1}_{h})$. The function ${\boldsymbol{w}}_{h}^{n+1}$ represents the displacement on $\widetilde{\Gamma}_{h}^{n+1}$ that moves tangentially the nodes maintaining good mesh properties. From this we have that $({\boldsymbol{v}}_{h}-{\boldsymbol{v}}_{h}^{\mathcal{A}})^{n+1}=-{\boldsymbol{w}}_{h}^{n+1}/\tau$.
After the surface has been displaced, a second step might be needed to extend the ALE motion in the bulk. The bulk mesh is advected through a continuous harmonic extension of ${\boldsymbol{w}}_{h}^{n+1}$. The Laplace problem reads correspondingly
(34) $\Delta{\boldsymbol{e}}_{h}^{n+1}$ $=0,\qquad\text{ in }\widetilde{\Omega}^{n+1}_{h},$
(35) ${\boldsymbol{e}}_{h}^{n+1}$ $={\boldsymbol{w}}_{h}^{n+1},\quad\text{ on }\partial\widetilde{\Omega}^{n+1}_{h},$
and we have that $({\boldsymbol{v}}_{h}-{\boldsymbol{v}}_{h}^{\mathcal{A}})^{n+1}=-{\boldsymbol{e}}_{h}^{n+1}/\tau$.
5 Advection-diffusion-reaction system
The model equation we consider reads
[…]
where $u:\mathcal{G}_{T}\to\mathbb{R}$ is the concentration, ${\boldsymbol{b}}_{M}$ is the advective velocity field, $\mathbb{A}_{M}$ is the diffusion matrix and $c_{M}>0$ is the reaction constant [52, 39, 42]. As in [39, p.304] we will assume that $(\mathbb{A}_{M})_{ij},\;({\boldsymbol{b}}_{M})_{i},\;\nabla_{M}\cdot{\boldsymbol{b}}_{M},\;c_{M}\in L^{\infty}(M),$ together with ${\boldsymbol{b}}_{M}({\boldsymbol{x}})\in T_{{\boldsymbol{x}}}M,\;\forall{\boldsymbol{x}}\in M$ and the requirement for $\mathbb{A}_{M}$ to be symmetric and to map the tangent space $T_{{\boldsymbol{x}}}M$ onto itself.
Extensive literature exists on finite element discretizations of such equations for what concerns the bulk case, see [46, 45]. For closed surfaces the SFEM discretization of parabolic PDEs in the form of (36) has been comprehensively reviewed in [39]. Convergence estimates for its continuous-in-time formulation using general order polynomials are found in e.g. [42]. For the sake of generality, here we focus on moving surfaces with boundary. The problem reads: Find $u_{h}\in V_{h}(\Gamma_{h})$ such that
(37) $\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}m_{h}(u_{h}\,,\,v_{h})+\overline{a}_{h}(u_{h}\,,\,v_{h})+(({\boldsymbol{b}}_{rel}u_{h}-\mathbb{A}_{\Gamma_{h}}\nabla_{\Gamma_{h}}u_{h})\cdot{\boldsymbol{n}}_{\partial\Gamma_{h}}\,,\,v_{h})_{\partial\Gamma_{h}}$
$=\overline{l}_{h}(v_{h})\;,\forall v_{h}\in V_{h}(\Gamma_{h}).$
where
(38) $m_{h}(u_{h},v_{h})=$ $~(u_{h}\,,\,v_{h})_{\Gamma_{h}},$
$\overline{a}_{h}(u_{h},v_{h})=$ $~(\mathbb{A}_{M_{h}}\nabla_{\Gamma_{h}}u_{h}\,,\,\nabla_{\Gamma_{h}}v_{h})_{\Gamma_{h}}+(-{\boldsymbol{b}}_{rel}u_{h}\,,\,\nabla_{\Gamma_{h}}v_{h})_{\Gamma_{h}}$
$+(c_{M_{h}}u_{h}\,,\,v_{h})_{\Gamma_{h}},$
$\overline{l}_{h}(v_{h})=$ $~(f\,,\,v_{h})_{\Gamma_{h}}.$
The coefficients $\mathbb{A}_{\Gamma_{h}},\;{\boldsymbol{b}}_{\Gamma_{h}},\;c_{\Gamma_{h}}$ are required to satisfy the necessary smoothness conditions element-wise on the discrete surface $\Gamma_{h}$. Recalling that ${\boldsymbol{v}}_{h}^{\mathcal{A}}$ is the discrete velocity associated with the ALE mapping, the relative velocity is defined as ${\boldsymbol{b}}_{rel}={\boldsymbol{b}}_{\Gamma_{h}}-{\boldsymbol{v}}^{\mathcal{A}}_{h}$. All boundary conditions are imposed weakly and the boundary is divided based on the type of boundary condition applied. We define as $\partial\Gamma_{h,D}$ the part of the boundary where Dirichlet conditions are applied and as $\partial\Gamma_{h,N}$ the part where Neumann boundary conditions are applied. It is required that
(39) $\partial\Gamma_{h,D}\cap\partial\Gamma_{h,N}=\emptyset,\quad\partial\Gamma_{h,D}\cup\partial\Gamma_{h,N}=\partial\Gamma_{h}.$
Nitsche’s boundary penalty method is used in order to impose the inhomogeneous Dirichlet boundary condition $u_{h}(t,{\boldsymbol{x}})|_{\partial\Gamma_{h,D}}=u(t,{\boldsymbol{x}})$ weakly for the diffusive term [87, 45]. For convenience, we define $v^{+}=\max\{v,0\}$ and $v^{-}=\min\{v,0\}$. The resulting bilinear forms are given by
(40) $a_{h}(u_{h}\,,\,v_{h})=$ $~\overline{a}_{h}(u_{h}\,,\,v_{h})-((\mathbb{A}_{\Gamma_{h}}\nabla_{\Gamma_{h}}u_{h})\cdot{\boldsymbol{n}}_{\partial\Gamma_{h}}\,,\,v_{h})_{\partial\Gamma_{h,D}}$
$-(u_{h}\,,\,(\mathbb{A}_{\Gamma_{h}}\nabla_{\Gamma_{h}}v_{h})\cdot{\boldsymbol{n}}_{\partial\Gamma_{h}})_{\partial\Gamma_{h,D}}+(\gamma_{\mathbb{A}}h^{-1}_{F}u_{h}\,,\,v_{h})_{\partial\Gamma_{h,D}}$
$+(({\boldsymbol{b}}_{rel}\cdot{\boldsymbol{n}}_{\partial\Gamma_{h}})^{+}u_{h}\,,\,v_{h})_{\partial\Gamma_{h}}$
$l_{h}(v_{h})=$ $~\overline{l}_{h}(v_{h})-(u\,,\,(\mathbb{A}_{\Gamma_{h}}\nabla_{\Gamma_{h}}v_{h})\cdot{\boldsymbol{n}}_{\partial\Gamma_{h}})_{\partial\Gamma_{h,D}}+(\gamma_{\mathbb{A}}h^{-1}_{F}u\,,\,v_{h})_{\partial\Gamma_{h,D}}$
$+((\mathbb{A}_{\Gamma_{h}}\nabla_{\Gamma_{h,N}}u)\cdot{\boldsymbol{n}}_{\partial\Gamma_{h}}\,,\,v_{h})_{\partial\Gamma_{h,N}}-(({\boldsymbol{b}}_{rel}\cdot{\boldsymbol{n}}_{\partial\Gamma_{h}})^{-}u\,,\,v_{h})_{\partial\Gamma_{h}},$
where $\gamma_{\mathbb{A}}>0$ is the penalty parameter enforcing $u_{h}=u$ on $\partial\Gamma_{h,D}$. Overall the problem for open surfaces reads: find $u_{h}\in V_{h}$ such that
(41) $\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}m_{h}(u_{h}\,,\,v_{h})+a_{h}(u_{h}\,,\,v_{h})=l_{h}(v_{h}),\;\forall v_{h}\in V_{h}(\Gamma_{h}).$
5.1 Numerical methods for advection-dominant problems
In biophysical cell models, as introduced in P.1, it is not uncommon to incur advection-dominant ADR equations. In our case, even purely parabolic equations might become advection-dominant due to the introduction of an ALE velocity ${\boldsymbol{v}}^{\mathcal{A}}$. In the present work we employ and extend the continuous interior penalty (CIP) stabilization to cope with the possibly dominant advection regime. The CIP method was proposed and analyzed in [21, 20] and then later extended to the case of stationary surfaces in [23]. Moreover, the CutFEM approach for moving surfaces developed in [53] applies a CIP-type stabilization for advection dominant problems on moving surfaces. We chose CIP for the following reasons:
  • It is a widely known, easily implementable technique whose implementation tools are usually shipped in classical finite element packages;
  • Although only weakly consistent, it has proven to be very successful in time-dependent problems given its commutativity with the time derivative [47];
  • Therefore, it has been shown to lead to convergent algorithms for discretizations of bulk, surface and implicitly described moving surfaces problems.
To formulate the CIP stabilization, we need to define averages and jumps of functions across edges and faces. For a piecewise discontinuous function $f$ defined on a surface or bulk mesh $\mathcal{T}_{h}(M_{h})$, we define its average and jump over an interior facet $F\in\mathcal{F}_{h}^{i}(M_{h})$ by
(42) $\{f\}|_{F}=\frac{1}{2}\left(f_{F}^{+}+f_{F}^{-}\right),\quad{\llbracket f\rrbracket|_{F}}=f_{F}^{+}-f_{F}^{-},$
respectively, where $f_{F}^{\pm}({\boldsymbol{x}})=\lim_{\delta\to 0^{+}}f({\boldsymbol{x}}_{F}^{\pm}-\delta{\boldsymbol{n}}_{F}^{\pm})$.
The main idea of the CIP approach is to penalize the jump of the streamline derivative across element interfaces. To this end, the CIP stabilization form is defined as follows,
(43) $s_{h}^{b}\left({u}_{h},{v}_{h}\right):=\gamma_{b}\sum_{F\in\mathcal{F}_{h}^{i}}\phi_{F}h_{F}\left(\llbracket{\boldsymbol{b}_{M_{h}}}\cdot\nabla_{M_{h}}{u}_{h}\rrbracket,\llbracket\boldsymbol{b}_{M_{h}}\cdot\nabla_{M_{h}}{v}_{h}\rrbracket\right)_{F},$
where $\gamma_{b}>0$ is a dimensionless stabilization constant and $\phi_{F}$ denotes a stabilization parameter defined by
(44) $\phi_{F}=\max(\phi_{T^{+}},\phi_{T^{-}}),\quad\phi_{T^{\pm}}=\min(\beta_{T^{\pm}}^{-1}h_{T^{\pm}},\mu_{T^{\pm}}^{-1}),\quad T^{\pm}\in\mathcal{T}_{h}.$
In the above $\beta_{T}$ is a local velocity scale and $\mu_{T}$ is the reciprocal of a time [47]. If ${\boldsymbol{b}}_{M_{h}}\in L^{\infty}(M_{h})\cap C^{0}(M_{h})$, then 43 can be simplified to
(45) $s_{h}^{b}({u}_{h},{v}_{h}):=\gamma_{b}\sum_{F\in\mathcal{F}_{h}^{i}}\beta_{F}h_{F}^{2}\left(\llbracket\nabla_{M_{h}}{u}_{h}\rrbracket,\llbracket\nabla_{M_{h}}{v}_{h}\rrbracket\right)_{F}.$
with $\beta_{F}=\max(\beta_{T^{+}},\beta_{T^{-}})$. We augment the system in (41) further to handle advection-dominant problems, which now reads: find $u_{h}\in V_{h}$ such that
(46) $\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}m_{h}(u_{h}\,,\,v_{h})+a_{h}(u_{h}\,,\,v_{h})+s_{h}^{b}(u_{h}\,,\,v_{h})=l_{h}(v_{h}),\;\forall v_{h}\in V_{h}(\Gamma_{h}).$
5.2 Numerical results for the ADR system
Given the novelty of the methods introduced, little to no theoretical convergence analysis is available in the literature and is left for future work. We thus proceed in reporting experimental convergence studies using select examples of increasing complexity to testify the accuracy of the framework.
We begin by considering a domain in 3D that moves under the linear transformation ${\boldsymbol{\Phi}}(t,{\boldsymbol{p}})={\boldsymbol{\Phi}}^{\mathcal{A}}(t,{\boldsymbol{p}})=\mathbb{A}(t){\boldsymbol{p}}+B(t)$ where
(47) $\mathbb{A}(t)=\begin{bmatrix}\cos(t)&-\sin(t)&0\\ \sin(t)&\cos(t)&0\\ 0&0&1\end{bmatrix},\quad B(t)=\begin{bmatrix}0\\ 0\\ 0\end{bmatrix}$.
We have that ${\boldsymbol{\Phi}}^{-1}(t,{\boldsymbol{x}})=[\mathbb{A}(t)]^{-1}({\boldsymbol{x}}-B(t))$ and that the domain velocity is ${\boldsymbol{v}}(t,{\boldsymbol{x}})=\mathbb{A}(t)^{\prime}{\boldsymbol{\Phi}}^{-1}(t,{\boldsymbol{x}})+B(t)^{\prime}\equiv{\boldsymbol{v}}^{\mathcal{A}}$. We construct manufactured solution for the problem
(48) $\partial_{t}u+\nabla_{\Gamma}\cdot({\boldsymbol{b}}_{\Gamma}u)+c_{\Gamma}u=f,$
with
(49) $u_{ex}$ $=\cos(2\pi x_{1})\cos(t),$
$c_{\Gamma}$ $=1+t^{2},$
${\boldsymbol{b}}_{\Gamma}$ $={\boldsymbol{v}}^{\mathcal{A}}+{\boldsymbol{b}}_{rel}={\boldsymbol{v}}^{\mathcal{A}}+\mathbb{P}_{\Gamma}(2,-{\boldsymbol{x}},0),$
where ${\boldsymbol{x}}=(x_{1},x_{2},x_{3})$. The right-hand side $f$ has been chosen so to satisfy the imposed solution. For $u_{ex}$ we have that $\int_{\Gamma}u(t)=0$ and that $u\in[-1,1]$. The geometry chosen is a half-sphere with unit radius that was cut along the $x_{1}$-$x_{2}$ plane. Under the above rotation the half-sphere has normal ${\boldsymbol{n}}(t,{\boldsymbol{x}})={\boldsymbol{x}}/\lVert{\boldsymbol{x}}\rVert$ . We test convergence properties for the following solvers:
  1. 1.
    adrSolver 1: CIP stabilized solver 46.
  2. 2.
    adrSolver 2: CIP stabilized solver with imposed bound preservation in the interval [1,1][-1,1].
  3. 3.
    adrSolver 3: CIP stabilized solver with imposed mass preservation using the lumped-mass inner product $(\cdot\,,\,\cdot)^{h}$.
  4. 4.
    adrSolver 4: CIP stabilized solver with both imposed bound preservation in the interval [1,1][-1,1] and mass preservation.
All solvers reveal the expected convergence in time and space. The results for adrSolver 4 with BDF-1 time stepping are shown in Figure 3. Analogous results are obtained for the other solvers.
Figure 3: Convergence studies for the solver adrSolver 4 in List 5.2. As expected, first and second order convergence are achieved in time and space, respectively.
To further test the effectiveness of the algorithm we now consider an ill-posed problem. Maintaining the half-sphere geometry and the transformation as in 47, we solve the pure transport problem
(50) $\partial_{t}u+\nabla_{\Gamma}({\boldsymbol{b}}_{\Gamma}u)=0,$
with
(51) ${\boldsymbol{b}}_{\Gamma}={\boldsymbol{v}}^{\mathcal{A}}+{\boldsymbol{b}}_{rel}={\boldsymbol{v}}^{\mathcal{A}}+(z,0,-x)\cdot(1-e^{-10z}),\quad u_{0}=e^{-3(x^{2}+y^{2})}.$
By construction the initial concentration is pushed towards the boundary where an incredibly sharp boundary layer forms, since zero flux boundary conditions are enforced by the fact that ${\boldsymbol{b}}_{rel}\cdot{\boldsymbol{n}}_{\partial\Gamma}=0$. Given the zero flux condition, the resulting solution is also mass-preserving. Parameters are set as follows
$h$ nvnv nene $T$ $\tau$
0.1 780 1495 1 0.01
where $nv$ is the number of vertices and $ne$ is the number of elements. We compare the results of the four different solvers of List 5.2 in Figure 5 and Figure 5. The bounds for adrSolver 2 and adrSolver 4 were set to $0,10^{5}]$ in order to guarantee positivity.
(a) Non-stabilized solver
(b) adrSolver 4
Figure 4: (a) Oscillating final solution of non-stabilized ADR solver at $t=1$. (b) Final solution for adrSolver 4 at $t=1$.
Figure 5: Left: relative absolute error in $\int_{\Gamma(t)}u(t)$ with respect to $\int_{\Gamma(0)}u(0)$ for the solvers in List 5.2. Right: Close-up of minimum value of $u$ for the different solvers in List 5.2.
It can be seen that the non-stabilized solver is unstable due to the pure advective regime, leading to oscillations that inevitably effect the whole domain 4(a). adrSolver 1 effectively controls the solution’s gradient thanks to the stabilization, generating a non-oscillating solution, but fails to maintain its positivity, see Figure 5 (right plot). This characteristic is lost along $\partial\Gamma$ where the species encounters the steep boundary layer. adrSolver 2 manages, thanks to the cutoff, to keep the solution positive but fails in maintaining mass-preservation as it can be seen in Figure 5 (left plot), where the relative absolute percentage mass error is plotted. adrSolver 4 successfully manages to deal with the boundary layer, maintain positivity, and avoid oscillations. The final solution is shown in 4(b).
6 Cahn-Hilliard system as a phase separation model
Phase field modeling plays a central role in biophysical systems to investigate phase transition processes. Common phase field models for stationary surfaces are derived from Allen-Cahn and Cahn-Hilliard dynamics [26, 44, 43, 35], where the motion follows the minimization of an energy functional. We will focus here on the functional
(52) $\mathcal{E}_{CH}(u)=\int_{M(t)}\sigma\Bigl(\underbrace{\frac{\varepsilon|\nabla_{\Gamma}u|^{2}}{2}}_{I}+\underbrace{\frac{1}{\varepsilon}F(u)}_{II}\Bigr),$
where $u$ is the phase field, $\varepsilon\in\mathbb{R}^{+}$ is a length scale parameter determining the interface length and $\sigma\in\mathbb{R}^{+}$ describes the interface tension between different phases. Term $I$ encodes the interface energy, penalizing inhomogeneities. Term $II$ encodes the local free energy and drives phase separation through the use of a double-well potential. We consider here two different double-well potentials
(53a) $F_{1}(u)$ $:=\frac{1}{4}((1-u)\log(1-u)+(1+u)\log(1+u))+\frac{1-u^{2}}{2},$
(53b) $F_{2}(u)$ $:=\frac{(1-u^{2})^{2}}{4},$
for which
(54) $F_{1}^{\prime}(u)=\log\Bigl(\frac{1+u}{1-u}\Bigr)-u,\text{ and }F_{2}^{\prime}(u)=u^{3}-u\text{ respectively}.$
$F_{1}$ is a Flory-Huggins-type mixing free energy density whose formulation finds its roots in configurational entropy [59, 48]. It can be seen that $F_{1}$ is only well posed for $u\in(-1,1)$. For this reason, a simpler polynomial potential $F_{2}$ is often used. The advantage of the latter is that it is well defined for values of $u$ outside the interval (1,1)(-1,1).
The extension of the Cahn-Hilliard equation to evolving surfaces has seen recent developments due to its applicability to biological membranes and alloys [105, 41]. The evolving surface Cahn-Hillard equation [24, 25] can be written as
(55) $\begin{cases}\partial^{\bullet}u+(\nabla_{\Gamma}\cdot{\boldsymbol{v}})u=\nabla_{\Gamma}(m\nabla_{\Gamma}w),\\ w=-\sigma\varepsilon\Delta_{\Gamma}u+\frac{\sigma}{\varepsilon}F^{\prime}(u),\end{cases}$
subject to the initial condition $u(0)=u_{0}$ for suitable initial data. $w$ represents the chemical potential and $m\in\mathbb{R}^{+}$ is the mobility. As explained in [77], models for evolving surfaces have similar energy functions to models on stationary surfaces, but important differences arise. Even if the evolution of the systems is driven by the minimization of (52), energy dissipation is no longer satisfied since the velocity is viewed as external force that adds to the system. Analogously, mass preservation depends on the velocity. To keep mass conservation, we must assume $\nabla_{\Gamma}\cdot{\boldsymbol{v}}=0$, which is equivalent to assume inextensibility of the cell membrane, and by consequence area preservation. This assumption also allows to consistently extend this one-phase model to $N$-phase models satisfying the hyperlink condition, as described in [77].
6.1 Numerical methods for Cahn-Hilliard system
For the discretization of (55), the evolutionary surface finite element method (ESFEM) is one of the most popular methods adopted and the one we consider here [39]. Its advantage lies in simple implementation, low memory consumption, ability to handle complex deformations and potential for efficient parallelization. We refer to [77] for a review of alternative methods. We employ a semi-implicit approach where the fully discrete formulation reads: given $u_{h}^{n}\in V_{h}(\Gamma_{h}^{n})$ find $u^{n+1}_{h},w^{n+1}_{h}\in V_{h}(\Gamma_{h}^{n})$ such that
(56a) $n_{h}\mathopen{}\left(\frac{u^{n+1}_{h}-u_{h}^{n}}{\tau}\,,\,\phi_{h}\mathclose{}\right)+(({\boldsymbol{v}}_{h}-{\boldsymbol{v}}^{\mathcal{A}}_{h})u^{n+1}_{h}\,,\,\nabla_{\Gamma^{n}_{h}}\phi_{h})_{\Gamma^{n}_{h}}+mb_{h}(w^{n+1}_{h}\,,\,\phi_{h})=0,$
(56b) $n_{h}(w^{n+1}_{h}\,,\,\psi_{h})-\sigma\varepsilon b_{h}(u^{n+1}_{h}\,,\,\psi_{h})=\frac{\sigma}{\varepsilon}n_{h}(F^{\prime}(u^{n}_{h})\,,\,\psi_{h}),$
for all $\phi_{h},\psi_{h}\in V_{h}(\Gamma_{h}^{n})$ with
(57) $n_{h}(u_{h}\,,\,\phi_{h})=(u_{h}\,,\,\phi_{h})_{\Gamma_{h}^{n}},\quad b_{h}(u_{h}\,,\,\psi_{h})=(\nabla_{\Gamma_{h}^{n}}u_{h}\,,\,\nabla_{\Gamma_{h}^{n}}\psi_{h})_{\Gamma_{h}^{n}}.$
Similar discretizations can be found in [43, 7, 85], where a Cahn-Hilliard solver is coupled with a domain evolution which is unknown. In those works, the potential $F_{2}$ was used. In our case, we can take advantage of the scheme presented in Section 3 to employ both $F_{1}$ and $F_{2}$, and apply mass conservation if needed. The scheme proceeds as follows.
In the predictor step we solve the unconstrained problem 56. Taking the non-linear term explicitly ensures $u_{h}^{n}$ lays inside the desired bound.
In the corrector step we solve the constrained problem imposing $u^{n+1}_{h}\in(-1,1)$.
Remark 6.1.
Another strength of the postprocessing technique of [29] evident here is that it can act directly on the quantity of interest per se without involving $w^{n+1}_{h}$, which is unbounded. We remind the reader that extensions for higher order BDFs methods are presented in the original articles.
6.2 Numerical results for Cahn-Hilliard system
We proceed in reporting experimental studies to assess the accuracy of the proposed framework. Consider a cylinder revolving around the $z$-axis with radius $R=1$ and height $l=2$. To satisfy the assumption $\nabla_{\Gamma}\cdot{\boldsymbol{v}}=0$, we transform the initial cylinder through the isometry ${\boldsymbol{\Phi}}(t,{\boldsymbol{p}})={\boldsymbol{\Phi}}^{\mathcal{A}}(t,{\boldsymbol{p}})=\mathbb{A}(t){\boldsymbol{p}}+B(t)$
(58) $\mathbb{A}(t)=\begin{bmatrix}\lambda(t)&0&0\\ 0&\lambda(t)&0\\ 0&0&1/\lambda(t)\end{bmatrix},\quad B(t)=\begin{bmatrix}0.2t\\ 0.1t\\ 0\end{bmatrix}$,
where $\lambda(t)=1+0.5\sin(\pi t)$. Convergence studies have already been proposed in the literature and we proceed by testing the quality of the structure preservation technique on realistic initial conditions. The verification is performed for the following solvers:
  1. 1.
    chSolver 1: Solver 56 with no postprocessing.
  2. 2.
    chSolver 2: Solver 56 with imposed bound preservation in the interval (1,1)(-1,1).
  3. 3.
    chSolver 3: Solver 56 with imposed mass preservation using the lumped-mass inner product $(\cdot\,,\,\cdot)^{h}$.
  4. 4.
    chSolver 4: Solver 56 with imposed bound preservation in the interval (1,1)(-1,1) and mass preservation.
Zero flux Neumann boundary conditions are imposed for both $u$ and $w$. A random uniform initial distribution in the interval [1,1][-1,1] is chosen for $u_{0}$. The parameters are set as
$h$ $m$ $\varepsilon$ $\sigma$ $T$ $\tau$
0.05 0.01 0.1 10 4 $2\cdot 10^{-3}$
Results for the potential $F_{1}$ are shown in Figure 7 and the same tests for $F_{2}$ are shown in Figure 7. In Figure 7, chSolver 1, 3 and 5 break after few steps due to the phase exceeding the bound (1,1)(-1,1), while the other solvers correctly reach the end of the simulation. Structure preservation properties are better appreciated in Figure 7. It can be seen that the energy landscape is not significantly modified by the use of either bound preservation and/or mass preservation. In both mass plots it is evident how the mass preservation property is lost for chSolver 2, that only has bound preservation, leading to a different overall dynamics for chSolver 2. Mass in instead correctly preserved by chSolver 4 as expected. Plotting the maximum and the minimum value in the bottom two pictures not only highlights the bound preservation but also visually testify that the phase separation dynamics is respected by the postprocessing step. This is particularly clear in the passages where the bound preservation is deactivated. For example around $t\approx 1$ in Figure 7, when the bound preservation is not needed, the maximum value of the bounded solvers naturally returns to align with the unbounded ones.
Figure 6: Experimental studies for potential $F_{1}(u)=\frac{1}{4}((1-u)\log(1-u)+(1+u)\log(1+u))+\frac{1-u^{2}}{2}$ and the different solvers in List 6.2.
Figure 7: Experimental studies for potential $F_{2}(u)=(1-u^{2})^{2}/4$ and the different solvers in List 6.2.
7 Membrane force system
In the realm of mathematical modeling, principal curvatures have long been used for the theory of elastic plates and shells, with the work of Kirchhoff [66] representing what may be the most famous case. Later in the ’70s, the work of Canham [27] and Helfrich [54] on the link between biological membranes shape and curvature functionals has de facto initiated an entirely new field of research focused on the use of geometry of hypersurfaces for biophysical modeling.
We will here focus on the Canham-Helfrich functional $\mathcal{E}_{B}(t)$ and its discretization. Considering membranes with boundaries, a commonly used expression for the energy functional is
(59) $\mathcal{E}_{B}(t)=\int_{\Gamma}\frac{\gamma_{W}}{2}(\kappa-\kappa_{0})^{2}+\int_{\Gamma}\gamma_{G}\kappa_{G}+\gamma_{\partial}\int_{\partial\Gamma}1,$
where $\gamma_{W},\;\gamma_{G}\in\mathbb{R}$ are bending rigidities, $\kappa:\mathcal{G}_{T}\to\mathbb{R}$ is the previously introduced mean curvature, $\kappa_{0}\in\mathbb{R}$ is the spontaneous curvature, and $\kappa_{G}:\mathcal{G}_{T}\to\mathbb{R}$ is the Gaussian curvature $\kappa_{G}=\det\mathbb{W}$. The parameter $\gamma_{\partial}$ takes into account the possible line energy of $\partial\Gamma$. For the sake of simplicity we will consider the case $\gamma_{G}=constant$, $\gamma_{W}=constant$, and assume that no topological change occur along the evolution. Taken into account these simplifications, the gradient flow of a membrane $\Gamma$ is a family of evolving surfaces with normal velocity $\boldsymbol{v}$$^{\perp}$ given by [88, 14]
(60) ${\boldsymbol{v}}^{\perp}=\gamma_{W}\Bigl(-\Delta_{\Gamma}\kappa+\frac{1}{2}\kappa(\kappa-\kappa_{0})^{2}-(\kappa-\kappa_{0})|\mathbb{W}|^{2}\Bigr){\boldsymbol{n}}_{\Gamma}.$
Note that such velocity is in the normal direction since tangential changes to the surface shape do not modify its energy. Equation (60) can be considered as a generalized Willmore flow equation. The mean curvature $\kappa$ is directly coupled to the surface $\Gamma$ through a second-order dependency. This can easily be seen in the case of surfaces whose configuration is described by a levelset equation. In that case, $\kappa$ is the trace of the extended Weingarten map $\mathbb{H}$ which in turn can be computed as the Hessian of the levelset for $\Gamma$. Equation (60) thus leads to a fourth-order nonlinear equation posed on a moving manifold.
Remark 7.1.
Other geometrical flows like the mean curvature flow and the surface diffusion flow are also of interest in this field. Being somewhat simpler, extensive research is already available for them, while less is known about the Willmore flow, which we focus on in this article. We refer to [38, 32, 15] for further details.
Remark 7.2.
Since this is a defining equation for the normal velocity, additional forces can be taken into account by introducing a right-hand side to (60).
7.1 Numerical methods for Helfrich system
We choose here the family of BGN-type algorithms as pioneered in [11] and later adapted for Willmore flow in [12, 14]. The schemes pivot around the two following ingredients:
  • Introduce the mean curvature vector ${\boldsymbol{\kappa}}(t)=\kappa(t){\boldsymbol{n}}_{\Gamma}(t)$ as independent variable and use the defining equation for the mean curvature vector ${\boldsymbol{\kappa}}=\Delta_{\Gamma}\mathbf{Id}_{\Gamma}$ in order to reduce the fourth order problem to a coupled second order one [11];
  • Use the weak formulation to rewrite and simplify the term $\kappa|\mathbb{W}|^{2}$ (more precisely $\kappa|\mathbb{H}|^{2}$ in the immersed setting) taking advantage of the newly introduced variable $\boldsymbol{\kappa}$}. This reformulation not only simplifies and linearizes the system but also leads to stable schemes as proven in [37].
Our scheme choice builds on the one derived in [14] for boundary value problems. Among the different types of boundary conditions considered in [14] we will restrict to the case of clamped boundary conditions, for which
[…]
where ${\boldsymbol{\mu}}(t)$ is a user-given function that dictates the evolution of ${\boldsymbol{n}}_{\partial\Gamma}(t)$. The weak form of (60) then reads: Find ${\boldsymbol{v}}^{\perp},~{\boldsymbol{y}},~{\boldsymbol{\kappa}}\in[H(t)]^{d}$ such that
(62a) $({\boldsymbol{v}}^{\perp}\,,\,{\boldsymbol{\phi}})=~(\nabla_{\Gamma}{\boldsymbol{y}}\,,\,\nabla_{\Gamma}{\boldsymbol{\phi}})+(\nabla_{\Gamma}\cdot{\boldsymbol{y}}\,,\,\nabla_{\Gamma}\cdot{\boldsymbol{\phi}})-((\nabla_{\Gamma}{\boldsymbol{y}})^{T}\,,\,\mathcal{D}({\boldsymbol{\phi}})\;\mathbb{P}_{\Gamma})$
$\qquad-\kappa_{0}({\boldsymbol{\kappa}}\,,\,(\nabla_{\Gamma}{\boldsymbol{\phi}})^{T}{\boldsymbol{n}}_{\Gamma})-\frac{1}{2}(\gamma_{W}|{\boldsymbol{\kappa}}-\kappa_{0}{\boldsymbol{n}}_{\Gamma}|^{2}\;\mathbb{P}\,,\,\nabla_{\Gamma}{\boldsymbol{\phi}})$
$\qquad+(({\boldsymbol{y}}\cdot{\boldsymbol{\kappa}})\;\mathbb{P}_{\Gamma}\,,\,\nabla_{\Gamma}{\boldsymbol{\phi}}),$
(62b) $({\boldsymbol{y}}\,,\,{\boldsymbol{\psi}})=~(\gamma_{W}({\boldsymbol{\kappa}}-\kappa_{0}{\boldsymbol{n}}_{\Gamma})\,,\,{\boldsymbol{\psi}}),$
(62c) $({\boldsymbol{\kappa}}\,,\,{\boldsymbol{\xi}})=~-(\mathbb{P}_{\Gamma}\,,\,\nabla_{\Gamma}{\boldsymbol{\xi}})+({\boldsymbol{\mu}}\,,\,{\boldsymbol{\xi}})_{\partial\Gamma},$
for all ${\boldsymbol{\phi}},{\boldsymbol{\psi}},{\boldsymbol{\xi}}\in[H(t)]^{d}$, where $H(t)$ is the Sobolev space $W^{1,2}(\Gamma(t))$. In the above $\mathcal{D}({\boldsymbol{\phi}})=\mathbb{P}_{\Gamma}(\nabla_{\Gamma}{\boldsymbol{\phi}}+\nabla_{\Gamma}{\boldsymbol{\phi}}^{T})\mathbb{P}_{\Gamma}$ and $\boldsymbol{y}$} is an auxiliary variable introduced to deal with the spontaneous curvature. For the discretization of the normal velocity in (62) we choose to introduce the displacement variable ${\boldsymbol{d}}_{h}^{n+1}\in[V_{h}(\Gamma_{h}^{n})]^{d}$ defined as
(63) ${\boldsymbol{v}}_{h}^{\perp}(t_{n})=\frac{1}{\tau}\sum_{i}({\boldsymbol{y}}_{i}(t_{n+1})-{\boldsymbol{y}}_{i}(t_{n})){\boldsymbol{\phi}}_{i}(t_{n})=\frac{1}{\tau}{\boldsymbol{d}}_{h}^{n+1},$
where we recall that ${\boldsymbol{y}}_{i}$ are the mesh points and $\{{\boldsymbol{\phi}}_{i}\}$ the associated vector-valued basis such that ${\boldsymbol{\phi}}_{i}({\boldsymbol{y}}_{j})=\delta_{ij}$ at all times. Moreover, (62c) and (62b) can be collapsed to
$\frac{1}{\gamma_{W}}({\boldsymbol{y}}^{n+1}\,,\,{\boldsymbol{\psi}})_{\Gamma^{n}}+(\nabla_{\Gamma^{n}}{\boldsymbol{d}}^{n+1}\,,\,\nabla_{\Gamma^{n}}{\boldsymbol{\psi}})_{\Gamma^{n}}=-(\mathbb{P}_{\Gamma^{n}}\,,\,\nabla_{\Gamma^{n}}{\boldsymbol{\psi}})_{\Gamma^{n}}$
(64) $+(\kappa_{0}{\boldsymbol{n}}_{\Gamma^{n}}\,,\,{\boldsymbol{\psi}})_{\Gamma^{n}}+({\boldsymbol{\mu}}\,,\,{\boldsymbol{\psi}})_{\partial\Gamma^{n}}.$
The fully discrete form of (62) reads: Given ${\boldsymbol{y}}^{n}_{h},{\boldsymbol{\kappa}}^{n}_{h}\in[V_{h}(\Gamma_{h}^{n})]^{d}$ find ${\boldsymbol{d}}^{n+1}_{h}\in[V_{h0}(\Gamma_{h}^{n})]^{d}$ and ${\boldsymbol{y}}^{n+1}_{h}\in[V_{h}(\Gamma_{h}^{n})]^{d}$ such that
(65a) $\mathopen{}\left(\frac{{\boldsymbol{d}}_{h}^{n+1}}{\tau}\,,\,{\boldsymbol{\phi}}_{h}\mathclose{}\right)_{\Gamma_{h}^{n}}^{h}-(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{y}}^{n+1}_{h}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}=~(\nabla_{\Gamma_{h}^{n}}\cdot{\boldsymbol{y}}^{n}_{h}\,,\,\nabla_{\Gamma_{h}^{n}}\cdot{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}$
$\qquad-((\nabla_{\Gamma_{h}^{n}}{\boldsymbol{y}}^{n}_{h})^{T}\,,\,\mathcal{D}({\boldsymbol{\phi}}_{h})\;\mathbb{P}_{\Gamma_{h}^{n}})_{\Gamma_{h}^{n}}-\kappa_{0}({\boldsymbol{\kappa}}^{n}_{h}\,,\,(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})^{T}{\boldsymbol{n}}_{\Gamma_{h}^{n}})_{\Gamma_{h}^{n}}^{h}$
$\qquad-\frac{1}{2}(\gamma_{W}|{\boldsymbol{\kappa}}^{n}_{h}-\kappa_{0}{\boldsymbol{n}}_{\Gamma_{h}^{n}}|^{2}\;\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}^{h}+(({\boldsymbol{y}}^{n}_{h}\cdot{\boldsymbol{\kappa}}^{n}_{h})\;\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}^{h},$
(65b) $\frac{1}{\gamma_{W}}({\boldsymbol{y}}^{n+1}_{h}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{d}}_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}$
$\qquad=-(\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}+(\kappa_{0}{\boldsymbol{n}}_{\Gamma_{h}^{n}}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+({\boldsymbol{\mu}}\,,\,{\boldsymbol{\psi}}_{h})_{\partial\Gamma^{n}},$
for all ${\boldsymbol{\phi}}_{h}\in[V_{h0}(\Gamma_{h}^{n})]^{d},~{\boldsymbol{\psi}}_{h}\in[V_{h}(\Gamma_{h}^{n})]^{d}$, where $(\cdot\,,\,\cdot)^{h}$ is the mass-lumped inner product.
Remark 7.3.
The mean curvature ${\boldsymbol{\kappa}}_{h}^{n+1}$ can be recovered a posteriori as ${\boldsymbol{\kappa}}_{h}^{n+1}=\pi^{h}(\gamma_{W}^{-1}{\boldsymbol{y}}_{h}^{n+1}+\kappa_{0}{\boldsymbol{n}}_{\Gamma_{h}^{n}})$ where $\pi^{h}(\cdot)$ is the standard interpolation operator.
Remark 7.4.
Note that (65) corresponds to the scheme presented in [12] for the choice of the parameter $\theta=1$, i.e. without tangential motion control.
7.2 Numerical results for Helfrich system
To test the accuracy of our method for gradient flows, we start by performing classical studies on Willmore flow such as the ones in [37, 12, 14]. For every test we compare
  1. 1.
    hfSolver 1: algorithm (65).
  2. 2.
    hfSolver 2: algorithm (65) together with the mesh redistribution described in (33). We highlight that it is of crucial importance that, when the algorithm in Section 4 is employed, geometrical quantities characteristic of the mesh are updated. In our case, we refer to the mean curvature ${\boldsymbol{\kappa}}_{h}$ and the auxiliary variable ${\boldsymbol{y}}_{h}$ of the algorithm in (65).
We start by testing the convergence using the example presented in [12, Section 5.3]. It is an evolving sphere with initial radius $R_{0}$ and spontaneous curvature $\kappa_{0}$. The radius evolution $R(t)$ is the solution to the ordinary differential equation
(66) $R_{t}=-\frac{\kappa_{0}}{R}\Bigl(\frac{2}{R}+\kappa_{0}\Bigr),\quad R(0)=R_{0}>0.$
The error norm chosen to verify convergence is the following
(67) $\lVert\Gamma-\Gamma_{h}\rVert_{L^{\infty}(L^{\infty})}=\max_{n}\max_{i}|\lVert{\boldsymbol{y}}_{i}(t_{n})\rVert-R(t_{n})|,$
where ${\boldsymbol{y}}_{i}$ are the vertices of the mesh simplices.
Figure 8: Convergence studies for Willmore flow of sphere under spontaneous mean curvature for the solvers in List 7.2.
Figure 9: Energy evolution for torus with $R=2$, $r=1$ for the solvers in List 7.2.
Results are presented in Figure 9. The convergence in space and time is clearly visible and in accordance with the optimal expected. Moreover, we can see that the use of the mesh redistribution actually improves the convergence rates. We believe this behavior is due to the automatic mesh adaptivity of Equation 33 (already mentioned in the original article [36]), which tends to accumulate nodes in certain regions providing better geometry description.
As a second test we check the energy evolution for a torus of major radius $R=2$ and minor radius $r=1$ as done in [12, Section 5.1]. The minimizer of the Willmore energy in this case is the Clifford torus with radius $R=\sqrt{2},\;r=1$, which has energy $\mathcal{E}=4\pi^{2}\approx 39.4784$. We use both solvers and plot the energy history for two different examples described by the following parameter sets:
$h$ nvnv nene $\tau$ $T$
0.2 2238 4476 $10^{-3}$ 2
$h$ nvnv nene $\tau$ $T$
0.1 8951 17902 $10^{-3}$ 2
The results are shown in Figure 9. We notice how both simulations using hfSolver 1 break before reaching the end time. This is due to instabilities that occur at regions of high curvature already observed in [12, Figure 2] for $\theta=1$. Applying the redistribution doesn’t impact the energy landscape and allows to proceed with the simulation until the stable state is reached. Moreover, refining the mesh shows how the energy converges toward the expected value.
7.2.1 Pinching tests for Helfrich system
We perform consolidated tests for the Willmore flow under the influence of spontaneous mean curvature as presented in [12, Section 5.3] and then repeated in [51], where a similar two-step procedure as the one adopted here is used. We consider two different tests for cigar-shaped surfaces:
  1. 1.
    hf-Test 1 with initial mesh as in 10(a) and the following sets of parameters
    $h$ nvnv nene $T$ $\kappa_{0}$
    0.25 601 1198 1 -2
    $h$ nvnv nene $T$ $\kappa_{0}$
    0.125 2321 4638 1 -2
  2. 2.
    hf-Test 2 with initial mesh as in 10(d) and the following sets of parameters
    $h$ nvnv nene $T$ $\kappa_{0}$
    0.25 845 1686 0.3 -3
    $h$ nvnv nene $T$ $\kappa_{0}$
    0.125 3282 6460 0.3 -3
A common timestep of $\tau=10^{-3}$ has been used for all the experiments. The idea is to test the ability of the algorithm to simulate pinching phenomena, since hf-Test 1 develops one neck and hf-Test 2 develops two necks. The results are plotted in Figure 10 for both hfSolver 1 and 2. It can be seen how the mesh adaptivity mentioned in [36] leads to a better resolution of the neck and overall better results for hfSolver 2. By comparison of 10(c) with [12, Figure 12] the shape is qualitatively in accordance with what is expected. Even more, while [12, Figure 12] requires a finer mesh to accurately resolve the three pearls, the mesh adaptivity of [36] alleviates that requirement for similar resolution. Figure 10 is also in accordance with [51, Figure 5.14], where a similar algorithm is employed for a cigar-shaped surface with different aspect ratio. As it happens in that article, it has to be noted that the refinement of the necks comes to the expense of a coarser mesh at the lobes. Looking at the energy evolution for hf-Test 2, the curve for hfSolver 1 and hfSolver 2 fall in the same place for all experiments, highlighting that the tangential mesh redistribution does not impact the energy landscape of the shape dynamics. For hf-Test 1 we have that the dynamics is stable for both examples when $h=0.2$, with a slight difference in energy when necking occurs. Refining the mesh leads to mesh breaking for hfSolver 1, while hfSolver 2 maintains its stability, asymptotically tending towards the same energy hfSolver 1 was tending to in the coarser case. Overall, the proposed technique appears stable in various regimes and accurate in resolving the energy evolution of the geometry.
(a) t=0
(b) t=0.5
(c) t=1
(d) t=0
(e) t=0.15
(f) t=0.3
(g) Willmore energy cigar hf-Test 1
(h) Willmore energy cigar hf-Test 2
Figure 10: (a)-(c): Mesh evolution for hf-Test 1 (coarser mesh). Top: hfSolver 1, bottom hfSolver 2. (d)-(f): Mesh evolution for hf-Test 2 (coraser mesh). Top: hfSolver 1, bottom hfSolver 2. (g) Energy evolution for hf-Test 1 and (h) Energy evolution for hf-Test 2 for the different solvers in List 7.2. We note that figures are rescaled.
8 Coupling
Simulating actual biological behavior involves complex coupling between different equations. Given we make heavy use of post-processing techniques, we couple the various solvers in a staggered way. We have then the option to choose between explicit and implicit staggering. Both approaches have shown to be effective in the literature [84, 65, 64, 51]. In [7, Appendix E], where fluid deformable membranes with phase separation are considered, both procedures show very similar results even in the most complex scenarios. Moreover, when implicit staggering is chosen, the authors of [7] claim that convergence is achieved in four steps on average. Strong of these results we proceed in coupling our solvers.
Remark 8.1.
An alternative option is to use a monolithic approach and couple the solvers in a unique system, avoiding staggering. In this way there is no need for sub-iterations and memory allocation is reduced. On the other side, the solution is usually only available for ad-hoc problems and not flexible. Together with this, eventual non-linearities have to be linearized. One can find an example of such a choice in [85], where wetting dynamics of liquid droplets on deformable membranes is simulated.
8.1 Numerical results for coupled mean curvature and ADR equations
We begin by testing the convergence properties of our coupled solver. To do so we pick the coupling example presented in [67]. The system which is solved is the following [67, p.686]:
(68a) $\partial^{\bullet}u+u\nabla_{\Gamma}\cdot{\boldsymbol{v}}-\Delta_{\Gamma}u$ $=f(t,{\boldsymbol{x}}),$
(68b) ${\boldsymbol{v}}-\alpha\Delta_{\Gamma}{\boldsymbol{v}}-\beta\Delta_{\Gamma}\mathbf{Id}_{\Gamma}$ $=(\delta u+g(t,{\boldsymbol{x}})){\boldsymbol{n}}_{\Gamma},$
with parameters $\alpha,\beta,\delta\in\mathbb{R}^{+}$ and ${\boldsymbol{x}}=(x_{1},x_{2},x_{3})$. Convergence studies are perfomed choosing $f,g$ such that the exact solution for $u$ is $u(t,{\boldsymbol{x}})=x_{1}x_{2}e^{-6t}$. The geometry is chosen to be a sphere whose radius evolves following the law
(69) $R(t)=\frac{r_{0}r_{K}}{r_{K}e^{-kt}+r_{0}(1-e^{-kt})},$
with parameters $r_{0},r_{K},k\in\mathbb{R}^{+}$. The parameters are chosen as follows
$T$ $\alpha$ $\beta$ $\delta$ $r_{0}$ $r_{K}$ $k$
1 0 1 0.4 1 2 0.5
To simulate (68a) we can use what has been introduced in Section 5 and add a right-had side coupling. Since we chose $\alpha=0$, we can simulate (68b) by simplifying (65). The result is a weighted mean curvature solver that, in this very case, takes the form
(70a) $\mathopen{}\left(\frac{{\boldsymbol{d}}_{h}^{n+1}}{\tau}\,,\,{\boldsymbol{\phi}}_{h}\mathclose{}\right)_{\Gamma_{h}^{n}}^{h}-\beta({\boldsymbol{\kappa}}^{n+1}_{h}\,,\,{\boldsymbol{\phi}}_{h})^{h}_{\Gamma_{h}^{n}}$ $=((\delta u+g(t,{\boldsymbol{x}})){\boldsymbol{n}}_{\Gamma_{h}^{n}}\,,\,{\boldsymbol{\phi}}_{h})^{h}_{\Gamma_{h}^{n}},$
(70b) $({\boldsymbol{\kappa}}^{n+1}_{h}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{d}}_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}$ $=-(\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}.$
We perform the test using implicit staggering. The results of the convergence studies are shown in Figure 11, where the same norm as in Equation 67 is used to study the results. One can see how the expected convergence is achieved in both space and time.
Figure 11: Convergence studies for mcadrSolver 4 and problem (68).
To qualitatively visualize the effect of the mesh redistribution and ADR stabilization, the example in [67, Sec. 11.2] is also reproduced. The simulation is derived from a proposed model for tumor growth. The coupled system is as follows:
(71a) $\partial^{\bullet}u+u\nabla_{\Gamma}\cdot{\boldsymbol{v}}-\Delta_{\Gamma}u$ $=f_{1}(u,w),$
(71b) $\partial^{\bullet}w+w\nabla_{\Gamma}\cdot{\boldsymbol{v}}-D_{c}\Delta_{\Gamma}w$ $=f_{2}(u,w),$
(71c) ${\boldsymbol{v}}-\alpha\Delta_{\Gamma}{\boldsymbol{v}}-\beta\nabla_{\Gamma}\mathbf{Id}_{\Gamma}$ $=\delta u{\boldsymbol{n}}_{\Gamma},$
with the non-linear couplings
(72) $f_{1}(u,w)=\gamma(a-u+u^{2}w),\quad f_{2}(u,w)=\gamma(b-u^{2}w).$
The parameters are set as
$T$ $\tau$ h $\alpha$ $\beta$ $\delta$ a b $\gamma$ $D_{c}$
8 $10^{-3}$ 0.07 0 0.01 0.4 0.1 0.9 100 10
As discussed in [67], the nonlinear system composed by (71a) and (71b) is solved until $t=5$ without coupling it to the mesh evolution, i.e. ignoring (71c). After that, the full system is evolved until $t=8$. Implicit staggering is used. We thus have the following solvers:
  1. 1.
    mcadrSolver 1: Implicit staggering of (41) and (70) and no mesh redistribution.
  2. 2.
    mcadrSolver 2: Implicit staggering of (46) and (70) and mesh redistribution.
In Figure 12 we compare the mesh quality at the final time $t=8$. It is evident that the redistribution step keeps the mesh well-behaved while maintaining solution accuracy.
(a) mcadrSolver 1
(b) mcadrSolver 2
Figure 12: Simulation at $t=8$ for the model (71) [67, Sec. 11.2] and the implicitly staggered schemes in List 8.1.
8.2 Numerical results for coupled Helfrich flow and Cahn-Hilliard euqations
We proceed now to test the coupling for more complicated systems and geometry evolution to explore the potential of our framework. The overdamped limit model for fluid deformable two-component membranes presented in [7, p.A41-33] is considered. The discrete scheme (65) is augmented with a Lagrange multiplier to maintain the inextensibility constraint $\nabla_{\Gamma_{h}^{n}}\cdot{\boldsymbol{v}}_{h}^{n}=0$. The modified version reads: Given ${\boldsymbol{y}}^{n}_{h},{\boldsymbol{\kappa}}^{n}_{h}\in[V_{h}(\Gamma_{h}^{n})]^{d}$ find ${\boldsymbol{d}}^{n+1}_{h}\in[V_{h0}(\Gamma_{h}^{n})]^{d}$, ${\boldsymbol{y}}^{n+1}_{h}\in[V_{h}(\Gamma_{h}^{n})]^{d}$ and $\lambda_{h}^{n+1}\in V_{h}(\Gamma_{h}^{n})$ such that
(73a) $\mathopen{}\left(\frac{{\boldsymbol{d}}_{h}^{n+1}}{\tau}\,,\,{\boldsymbol{\phi}}_{h}\mathclose{}\right)_{\Gamma_{h}^{n}}^{h}-(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{y}}^{n+1}_{h}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}+(\lambda_{h}^{n+1}{\boldsymbol{n}}_{\Gamma_{h}^{n}}\,,\,{\boldsymbol{\phi}}_{h})$
$\qquad=~p_{h}({\boldsymbol{y}}_{h}^{n}\,,\,{\boldsymbol{\phi}}_{h}),$
(73b) $\frac{1}{\gamma_{W}}({\boldsymbol{y}}^{n+1}_{h}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{d}}_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}$
$\qquad=-(\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}+(\kappa_{0}{\boldsymbol{n}}_{\Gamma_{h}^{n}}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+({\boldsymbol{\mu}}\,,\,{\boldsymbol{\psi}}_{h})_{\partial\Gamma^{n}},$
(73c) $(\nabla_{\Gamma_{h}^{n}}\lambda_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}\mu_{h})_{\Gamma_{h}^{n}}+(|{\boldsymbol{\kappa}}_{h}^{n}|^{2}\lambda_{h}^{n+1}\,,\,\mu_{h})_{\Gamma_{h}^{n}}$
$\qquad=((\nabla_{\Gamma_{h}^{n}}{\boldsymbol{y}}^{n}_{h}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\kappa}}^{n}_{h})_{\Gamma_{h}^{n}}\,,\,\mu_{h})+(p_{h}({\boldsymbol{y}}_{h}^{n}\,,\,{\boldsymbol{\kappa}}_{h}^{n})\,,\,\mu_{h}),$
for all ${\boldsymbol{\phi}}_{h}\in[V_{h0}(\Gamma_{h}^{n})]^{d},~{\boldsymbol{\psi}}_{h}\in[V_{h}(\Gamma_{h}^{n})]^{d},~\mu_{h}\in V_{h}(\Gamma_{h}^{n})$ where
$p_{h}({\boldsymbol{y}}_{h}^{n}\,,\,{\boldsymbol{\phi}}_{h})=(\nabla_{\Gamma_{h}^{n}}\cdot{\boldsymbol{y}}^{n}_{h}\,,\,\nabla_{\Gamma_{h}^{n}}\cdot{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}-((\nabla_{\Gamma_{h}^{n}}{\boldsymbol{y}}^{n}_{h})^{T}\,,\,\mathcal{D}({\boldsymbol{\phi}}_{h})\;\mathbb{P}_{\Gamma_{h}^{n}})_{\Gamma_{h}^{n}}$
$\qquad-\kappa_{0}({\boldsymbol{\kappa}}^{n}_{h}\,,\,(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})^{T}{\boldsymbol{n}}_{\Gamma_{h}^{n}})_{\Gamma_{h}^{n}}^{h}-\frac{1}{2}(\gamma_{W}|{\boldsymbol{\kappa}}^{n}_{h}-\kappa_{0}{\boldsymbol{n}}_{\Gamma_{h}^{n}}|^{2}\;\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}^{h}$
(74) $\qquad+(({\boldsymbol{y}}^{n}_{h}\cdot{\boldsymbol{\kappa}}^{n}_{h})\;\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}^{h}+({\boldsymbol{f}}_{h}^{n+1}\,,\,{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}^{h}.$
In (74), ${\boldsymbol{f}}_{h}^{n+1}$ is a general right-hand side. The system is composed of the following three steps:
  1. 1.
    Solver (73) with the right-hand side (recall (52)):
    (75) ${\boldsymbol{f}}_{h}^{n+1}=\mathcal{E}_{CH}(u_{h}^{n+1}){\boldsymbol{\kappa}}_{h}^{n}-\sigma\epsilon\nabla_{\Gamma_{h}^{n}}u_{h}^{n+1}\mathbb{H}_{h}^{n}\nabla_{\Gamma_{h}^{n}}u_{h}^{n+1},$
    where ${\boldsymbol{\omega}}_{\Gamma_{h}^{n}}$ is a certain vertex normal, see [11] for details, and $\mathbb{H}_{h}^{n}=-\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\omega}}_{\Gamma_{h}^{n}}$ is the discrete extended Weingarten map.
  2. 2.
    Solver (33) where the prescribed material velocity is set to
    (76) ${\boldsymbol{v}}_{h}^{\top}=-\nabla_{\Gamma_{h}^{n}}\lambda_{h}^{n+1},\quad{\boldsymbol{v}}_{h}^{\perp}={\boldsymbol{d}}_{h}^{n+1}/\tau.$
  3. 3.
    Solver (56) with tangential advective velocity ${\boldsymbol{v}}_{h}^{n+1}={\boldsymbol{v}}_{h}^{\top}$ and polynomial potential $F_{2}$ (53).
The test considers a sphere of radius one and the following initial parameters:
$T$ $h$ $\sigma$ $\varepsilon$ $m$
1 0.08 $1.5\sqrt{2}$ 0.1 0.001
The common initial condition for the Cahn-Hilliard solvers is
(77) $u_{0}({\boldsymbol{x}})=\cos(\pi x_{1})\cos(\pi x_{2})\cos(\pi x_{3}).$
The simulation is run with variable elasticity modulus $\gamma_{W}\in\{0.5,~0.1,~0.02\}$ following the parameters in [7]. The timestep is varied accordingly with values $\tau\in\{10^{-4},~10^{-4},2.5\cdot 10^{-5}\}$ respectively. The simulation is stopped once the area difference between initial and evolved surface differs by more than $10\%$. Results are presented in Figure 13, where it can be seen that the characteristic bulging dynamics is correctly recovered.
(b) $\gamma_{W}=0.5,~t=1$
(c) $\gamma_{W}=0.1,~t=0.75$
(d) $\gamma_{W}=0.02,~t=0.2$
Figure 13: (a) Initial setup for problem the problem 77. (b)-(d) Mesh for different elasticity modulus. The simulation is shown at end time $t=1$ or at last timestep before inextensibility threshold is violated.
9 Conclusions and outlook
Here, we presented a numerical framework to simulate moving boundary problems in biophysics. This work addresses a long-standing need for robust computational frameworks to tackle such problems. While other frameworks exist [85, 6, 7, 83, 13], our work targets flexibility of application while maintaining accuracy. Algorithms that enjoy stability and convergence properties are tied to non-invasive postprocessing techniques so to favor complex applications. The result is a stable pipeline with the capability of being scaled to more realistic scenarios.
Specifically, structure preservation is introduced on moving bulk and surface domains as a mean to allow for biophysically justified models and guarantee interpretability. We show that the scheme is readily implementable while being fast and accurate. The flexibility of the algorithm, which does not require ad hoc implementation, is proven by applying it to both advection-dominant ADR equations and phase-field models of the Cahn-Hilliard type. Convergence studies were performed for the bound- and mass-preserving scheme and the expected convergence rates were numerically verified. The framework also couples recent mesh redistribution techniques for gradient flows with the ALE method. We showed that this union is efficient in handling complex reshaping dynamics without the need for remeshing. As a result, the algorithm is tunable to the needs of the user and does not modify the underlying domain evolution. This is crucial in biophysical settings where the dynamics of the overall system is strongly coupled to the domains’ shape. Convergence studies were performed, demonstrating the accuracy of the scheme. Additionally, complex benchmark tests were conducted, showing good agreement with previous results in the literature. Building on previous results, we show that staggering is an effective method to simulate increasingly complex physics such as those that occur in biophysics.
The framework developed here has strong potential to simulate problems with increasing biological complexity. To achieve this goal, we foresee the need to incorporate large systems of ADR equations, that are not discussed in the present work [50, 83, 82]. Furthermore, fluid equations have also been shown to play an important role at certain timescales [94, 96, 2]. Including Navier-Stokes type equations on both interfaces and bulk [85, 7] will broaden the applicability of the computational schemes presented here.
In summary, we presented a finite element framework able to handle a rich set of reshaping dynamics arising from bulk-surface coupled PDEs in biophysics. This framework poses itself as a major stepping stone on the path to physical simulations of moving boundary problems in cell biology.
Acknowledgments
We acknowledge Dr. Emmet Francis for his valuable input on biologically relevant models for computational biology and for sharing the lessons learned from his programming experience [50].
Declaration of Interests
P.R. is a consultant for Simula Research Laboratories in Oslo, Norway and receives income. The terms of this arrangement have been reviewed and approved by the University of California, San Diego in accordance with its conflict-of-interest policies.

### v2

@@
 Abstract
-We consider moving boundary problems for biophysics and introduce a new computational framework to handle the complexity of the bulk-surface PDEs.
-In our framework, interpretability is maintained by adapting the fast, generalizable and accurate structure preservation scheme in [29].
-We show that mesh distortion is mitigated by adopting the pioneering work of [36], which is tied to an Arbitrary Lagrangian Eulerian (ALE) framework.
-We test our algorithms accuracy on moving surfaces with boundary for the following PDEs: advection-diffusion-reaction equations, phase-field models of Cahn-Hilliard type, and Helfrich energy gradient flows.
-We performed convergence studies for all the schemes introduced to demonstrate accuracy.
-We use a staggered approach to achieve coupling and further verify the convergence of this coupling using numerical experiments.
-Finally, we demonstrate broad applicability of our work by simulating state-of-the-art tests of biophysical models that involve membrane deformation.
+Cellular morphodynamics requires solving systems of coupled partial differential equations on moving bulk and surface domains, where advection-dominant transport, structure preservation, and severe mesh distortions make robust simulation difficult.
+We present a holistic finite element framework that jointly addresses these obstacles for biophysical applications by combining model-agnostic structure-preserving postprocessing, ALE-based mesh redistribution strategies driven by surface-tangential velocities, and stabilized discretization for advection-diffusion-reaction problems tailored to evolving domains.
+The methodology is modular and applies to advection-diffusion-reaction systems, Cahn–Hilliard phase separation, Helfrich-type geometric flows, as well as their staggered and potentially mixed-dimensional couplings.
+We provide a concise notation for evolving bulk and surface geometries, extend positivity-, bound-, and mass-preserving projections to moving meshes, and develop a two-step redistribution procedure that maintains element quality without remeshing.
+Convergence studies, manufactured solutions, and biologically motivated test cases —including tumor-growth surrogates and phase segregation on deformable membranes— demonstrate accuracy, stability, and versatility across the problem classes considered.
 keywords:
-Biophysics, moving boundary problems, bulk-surface coupling, structure preservation, mesh redistribution, multiphysics coupling
-{MSCcodes}
-35R37, 74K15, 65N30, 76T99.
+Biophysics , moving boundary problems , structure preservation , mesh redistribution , multiphysics coupling , bulk-surface coupling
+2008 MSC:
+35R37 , 74K15 , 65N30 , 76T99
+journal: Computer Methods in Applied Mechanics and Engineering
+[…]
+[…]
+{highlights}
+Fast and accurate structure preserving scheme for moving bulk and surface PDEs
+Flexible mesh-redistribution algorithm for biological moving boundary problems
+Stabilized advection-dominant transport equations on moving surfaces
+Spacetime convergence for structure preserving PDEs on moving surfaces
+Application to state-of-the-art coupled problems involving deformable membranes
 1 Introduction
+1.1 Background and state of the art
 Many biophysical phenomena in cell biology can be characterized by the feedback between form and function.
@@
 Such shape changes are a result of coordination of multiple biochemical and mechanical processes within the cell.
-Not surprisingly, these phenomena have been studied extensively using models based on partial differential equations (PDEs) with different levels of biophysical complexity [78].
-In concert with experiments, such models have provided insight into the mechanisms underlying the change in cell shape [31, 95, 63].
-However, an open challenge in the field of computational biophysics is the development of robust computational tools for numerical solutions of the governing PDEs.
-For example, cellular membranes are mechanical boundaries that are characterized by lateral dimensions that are large compared to their thickness, appearing in a variety of structures in cells [99].
-It has been proven that the properties of these surfaces are crucial for the genesis and maintenance of cellular membrane systems [16, 86].
-These restructuring phenomena strongly depend on the components of the membranes themselves and their surroundings [19, 10, 96, 17, 73, 104, 2, 5, 98], leading to PDE systems coupling complex multiphysics to free surfaces problems.
+Not surprisingly, these phenomena have been studied extensively using models based on partial differential equations (PDEs) with different levels of biophysical complexity Lomakin et al.
+[2015], Auddya et al.
+[2021].
+In concert with experiments, such models have provided insight into the mechanisms underlying the change in cell shape Dayel et al.
+[2009], Sens [2020], Keren et al.
+[2008].
+However, an open challenge in the field of computational cell biophysics is the development of robust computational tools for numerical solutions of the governing PDEs.
+For example, cellular membranes are mechanical boundaries that are characterized by lateral dimensions that are large compared to their thickness, appearing in a variety of structures in cells Steigmann [2018].
+It has been proven that the properties of these surfaces are crucial for the genesis and maintenance of cellular membrane systems Baumgart et al.
+[2003], Mullins [2005].
+These restructuring phenomena strongly depend on the components of the membranes themselves and their surroundings Brangwynne et al.
+[2009], Banani et al.
+[2017], Shin and Brangwynne [2017], Beutel et al.
+[2019], Lee et al.
+[2020b], Zhao and Zhang [2020], Alberti and Hyman [2021], Ambroggio et al.
+[2021], Snead et al.
+[2022], leading to PDE systems coupling complex multiphysics to free surfaces problems.
 The resulting systems typically exhibit highly non-linear dynamics and a general numerical framework to simulate these rich dynamics in a unified and stable setting is still missing.
 The goal of this article is to present a finite element based computational framework which addresses some of the key challenges arising in continuum-mechanics based computational models of cells.
-1.1 State of the art
-Numerous numerical methods have been proposed to deal with the challenges posed by a moving deformable membrane [61, 92, 1, 71, 93, 69, 90, 85, 7], including: immersed boundary methods, level-set method, mesh-free methods, particle methods, and the phase-field method.
-Despite numerous successes, even the most advanced schemes seem to lack the following features:
-  1. P.1
-    It is not rare, in the targeted applications, to incur advection-dominant (AD) equations possibly posed on a moving surface.
-Characteristic examples can be the transport of surfactants in two-phase flow or the interaction between actin’s barbed ends and the cellular membrane.
-There exists a large bulk of literature for parabolic advection-diffusion-reaction (ADR) PDEs on moving surfaces [38, 40, 18, 80, 39, 67, 68, 42, 25].
-For what concerns AD equations, the bulk setting has been extensively explored and the surface case has seen recent developments  [91, 22, 97, 23, 33, 100, 101].
-Surprisingly, most aforementioned approaches only consider the case of stationary surfaces, leaving the AD case largely unaddressed.
-A notable exception is the work in [53], where characteristic cut finite element methods on moving surfaces is proposed.
-  2. P.2
-    Physical structure preservation (SP) is still overlooked by the majority of the available articles.
-It is crucial, in order to maintain physical interpretability as well as numerical stability, to be able to guarantee constraints such as: mass and density positivity, phase bounds, energy decay, mass preservation, etc.
-A vast and diversified literature exist dedicated to preserve bounds, see [79, 74, 34, 75, 103, 102, 76, 28, 58, 56] and references therein.
-Unfortunately, this richness of options is lacking if we consider PDEs on surfaces, let alone moving surfaces.
-  3. P.3
-    Most articles assume idealized geometries, in which the complexity is still highly reduced with respect to realistic geometries resulting from imaging techniques [72].
-This is especially crucial in this context since we expect the domain evolution to be dependent on the coupled equations and not known a priori.
-Steps are continuously being taken in this direction for what concerns geometric flows, which notoriously drive meshes towards bad conditioning, see [11, 81, 57, 36, 9, 51].
-However, even in these cases, one is quite far from the structures considered in more applied articles such as [50].
-The present work aims to propose a first step towards a cure of the missing pieces described above.
+Numerous numerical methods have been proposed to deal with the challenges posed by a moving deformable membrane Iyer et al.
+[2023], Ong and Lai [2020], Aland et al.
+[2014], Laadhari et al.
+[2014], Rosolen et al.
+[2013], Krüger et al.
+[2011], Noguchi [2004], Mokbel et al.
+[2024], Bachini et al.
+[2023], including: parametric finite element methods, immersed boundary methods, level-set method, mesh-free methods, particle methods, and the phase-field method.
+We refer to Bonito and Nochetto [2020][Chapter 4-6] for a detailed review of some of these methods.
+However, even the most advanced schemes often do not simultaneously address the following features which are frequently crucial in the context of biophysical cell modeling.
+First, it is not rare, in the targeted applications, to incur advection-dominant equations possibly posed on a moving surface.
+Characteristic examples can be the transport of surfactants in two-phase flow Gross and Reusken [2011], Ganesan et al.
+[2017] or the interaction between actin’s barbed ends and the cellular membrane Bonilla-Quintana and Rangamani [2024], Doubrovinski and Kruse [2011].
+There exists a large bulk of literature for parabolic advection-diffusion-reaction (ADR) PDEs on moving surfaces Dziuk and Elliott [2007], Dziuk et al.
+[2012], Bonito et al.
+[2013], Lubich et al.
+[2013], Dziuk and Elliott [2013], Kovács et al.
+[2017], Kovács and Power Guerra [2018], Elliott and Ranner [2021], Caetano et al.
+[2023], but primarily with a focus on the diffusion-dominant regime.
+For what concerns advection-dominant equations, the bulk setting has been extensively explored and the surface case has seen recent developments  Olshanskii et al.
+[2014], Burman et al.
+[2020], Simon and Tobiska [2019], Burman et al.
+[2019], Dedner and Madhavan [2015], Ulfsby et al.
+[2023], Xiao et al.
+[2020].
+Surprisingly, most aforementioned approaches only consider the case of stationary surfaces, except for Hansbo et al.
+[2015], where a characteristic cut finite element method on moving surfaces is proposed.
+Second, the major part of the existing literature on surface PDEs pays little attention to structure preservation of physical quantities.
+In contrast, computational cell biophysics models often require the preservation of physical constraints —including mass conservation, density positivity, phase bounds, and energy decay— to maintain both physical interpretability and numerical stability of the computed solution.
+While there is a vast and diversified literature on bulk PDEs dedicated to preserve positivity, bounds and mass Lu et al.
+[2013], Li et al.
+[2020], Du et al.
+[2021], Li and Zhang [2020], Zhang and Shu [2011, 2010], Liu et al.
+[2018], Chen et al.
+[2019], Huang and Shen [2021], Hu and Huang [2020] this richness of options is unfortunately lacking if we consider PDEs on surfaces, let alone moving surfaces.
+Finally, in the context of cell shape dynamics, the domain evolution depends on a set of coupled PDEs and is not known a priori.
+For classical prototypical geometric flows, such as mean curvature flow, surface diffusion flow, and Helfrich flow, a large number of numerical methods have been proposed, see, e.g.
+Dziuk [2008], Barrett et al.
+[2007], M.
+Elliott and Fritz [2017], Hu and Li [2022], Duan and Li [2024], Bai et al.
+[2024], Garcke et al.
+[2025], assuming mostly idealized geometries.
+It is well-known that for most finite-element based algorithms, such geometric flows notoriously drive surface meshes towards bad conditioning.
+Consequently, it is crucial to design numerical methods that can maintain good mesh quality during the shape evolution also in the presence of more realistic geometries resulting e.g.
+from imaging techniques Lee et al.
+[2020a], Yang et al.
+[2016], Gulati and Rudraraju [2024], or when coupled to other bulk-surface PDEs Francis et al.
+[2024].
+1.2 New contributions and outline
+The present work aims to present a unified finite-element based computational framework which focuses on the three challenges outlined above: advection-dominant transport on moving surfaces, structure-preservation, and handling of large mesh deformations in problems where multi-physics, potential multi-dimensional problems are nonlinearly coupled to the shape evolution.
 It is meant to bridge the complexity gap between numerical accuracy and biophysical complexity in moving domains scenario.
 While there exist a plethora of numerical methods which each target a specific problem class or particular issue, we here employ and extend some of the most recent state-of-the-art methods and combine them into a holistic simulation framework which can handle coupled multiphysics and shape deformation problems arising in computational cell biology.
-1.2 New contributions and outline
-In Section 2, we provide the notation necessary for the formulation and discretization of PDEs posed on evolving domains.
-For purposes of generality, we keep the notation the same in the bulk and on the surface, only distinguishing between the two as needed.
-In Section 3, we adapt the structure-preserving scheme in [30] to moving bulk and surface problems as a solution to P.2.
-The model-agnostic, point-wise approach of [30] is fit to the first-order, piecewise-linear finite element framework here used.
-The result is a fast, generalizable and flexible scheme that can be readily adapted to multiple PDEs posed on moving bulk and surfaces.
-In this context, we use it to bound the PDE solution to physically-relevant values while at the same time preserving the total mass.
-In Section 4, we introduce a mesh redistribution algorithm based on the surface tangential motion of [36] and an Arbitrary Lagrangian Eulerian (ALE) approach.
-The two-step scheme in [36] is decoupled from the gradient flow dynamics and applied to general motions.
+We start in Section 2 by providing the notation necessary for the formulation and discretization of PDEs posed on evolving domains.
+For purposes of generality, the notation is kept general to cover both evolving surface and bulk domains, only distinguishing between the two as needed.
+In Section 3, we adapt the structure-preserving scheme in Cheng and Shen [2022b] to moving bulk and surface problems as a general approach to design positivity, bound and mass preserving schemes.
+The model-agnostic, pointwise approach of Cheng and Shen [2022b] is fit to the first-order, piecewise linear finite element framework used here.
+The result is a fast, generalizable and flexible approach that can be readily adapted to multiple PDEs posed on moving bulk and surfaces.
+In our context, it is used to bound PDE solution to physically-relevant values while at the same time preserving the total mass.
+Then, in Section 4, we present a mesh redistribution algorithm based on the surface tangential motion of Duan and Li [2024] and an Arbitrary Lagrangian Eulerian (ALE) approach.
+The two-step scheme in Duan and Li [2024] is decoupled from the gradient flow dynamics and applied to general motions.
 This allows us to greatly alleviate the need of remeshing when simulating domains characterized by large deformations.
 At the same time, the scheme does not interfere with the shape evolution, maintaining accuracy.
-Both schemes in Section 3 and Section 4 take the form of a post-processing step and thus have the advantage of being a tunable feature of the code.
-In Section 5 we augment an ADR scheme for moving surfaces with a continuous interior penalty (CIP) stabilization to handle AD cases and solve P.1.
-The stabilization is proven to be convergent and so is its coupling with the structure-preserving algorithm of Section 3.
+The main advantage of the schemes presented in Section 3 and Section 4 is that they can be realized as a post-processing step and thus represent tunable features of the simulation framework.
+In Section 5 we augment an ADR scheme based on the evolving surface finite element method with a continuous interior penalty (CIP) stabilization to handle advection-dominant problems.
+We demonstrate that stabilized ADR discretization, potentially combined with a structure preservation postprocessing step, yields optimal convergent schemes for the problem class.
 The flexibility of the setting is further verified with an additional real-case example.
-In Section 6 we apply structure preservation to a Cahn-Hilliard type equation on moving surfaces.
-Numerical tests are presented that highlight how the algorithm in Section 3 is readily applicable to moving surfaces and true to the underlying physics.
-In Section 7, we focus on P.3, applying the mesh redistribution scheme of Section 4 to the parametric method of Barrett, Garke and Nürnberg [14] for Helfrich flow.
+Furthermore, we apply structure preservation to a Cahn-Hilliard type equation on moving surfaces in Section 6.
+Numerical tests are presented that highlight how the algorithm in Section 3 is readily applicable to moving surfaces and preserves the underlying physics.
+In Section 7, we combine the efficient mesh redistribution scheme of Section 4 with the parametric method of Barrett, Garke and Nürnberg Barrett et al.
+[2017a] for Helfrich flow.
 The postprocessing step is extensively tested with state-of-the-art examples revealing good convergence properties and long-term stability.
-In Section 8, we couple the solvers in Section 5 through 7 together in a staggered approach.
-Convergence is tested with manufactured solutions [67] where the ADR algorithm and a simplified version of the Helfrich algorithm are coupled.
+Afterwards, in Section 8, we couple the solvers from Section 5 through 7 together in a staggered approach.
+Convergence is tested with manufactured solutions Kovács et al.
+[2017] where the ADR algorithm and a simplified version of the Helfrich algorithm are coupled.
 In the same context a PDE system modeling tumor growth is also reproduced.
-The applicability of our approach to relevant cell biophysics is further explored simulating phase separation on a deformable membrane, where the Cahn-Hilliard and Helfrich schemes are coupled.
+The applicability of our approach to relevant cell biophysics is further explored simulating phase separation on a deformable membrane, where the presented schemes for Cahn-Hilliard and Helfrich are coupled.
 2 Model
 This article explores numerical approaches to biophysical cell phenomena where the dynamics involve moving boundaries, such as cell membranes, that are assumed to have negligible thickness and thus are treated as moving hypersurfaces.
-As a result, we begin by outlining the essential notation required for both formulating and discretizing partial differential equations on evolving domains, with specific attention given to surface-bound PDEs [39, 15].
+As a result, we begin by outlining the essential notation required for both formulating and discretizing partial differential equations on evolving domains, with specific attention given to surface-bound PDEs Dziuk and Elliott [2013], Barrett et al.
+[2020].
 Let $M\subset\mathbb{R}^{d},\;d=2,3$, be an $m$-dimensional oriented, compact, $C^{2}$ manifold with boundary $\partial M$.
@@
 For $\dim{(M)}=d-1$ we define the tangential projection at $\boldsymbol{p}$} as
-(1) $\mathbb{P}_{M}({\boldsymbol{p}})=\mathbb{I}-{\boldsymbol{n}}_{M}({\boldsymbol{p}})\otimes{\boldsymbol{n}}_{M}({\boldsymbol{p}}),$
+$\mathbb{P}_{M}({\boldsymbol{p}})=\mathbb{I}-{\boldsymbol{n}}_{M}({\boldsymbol{p}})\otimes{\boldsymbol{n}}_{M}({\boldsymbol{p}}),$ (1)
 where $\mathbb{I}$ is the identity matrix.
@@
 This allows us to define the tangential gradient of a differentiable function $f:M\to\mathbb{R}$ at $\boldsymbol{p}$} as
-(2) $\nabla_{M}f({\boldsymbol{p}})=\nabla\bar{f}({\boldsymbol{p}})-\mathbb{P}_{M}({\boldsymbol{p}})\nabla\bar{f}({\boldsymbol{p}}),$
+$\nabla_{M}f({\boldsymbol{p}})=\nabla\bar{f}({\boldsymbol{p}})-\mathbb{P}_{M}({\boldsymbol{p}})\nabla\bar{f}({\boldsymbol{p}}),$ (2)
 where $\bar{f}$ is a smooth extension of $f$ to a $d$-dimensional neighborhood of $M$ such that $\bar{f}\rvert_{M}=f$ and $\nabla$ is the Euclidean gradient in $\mathbb{R}^{d}$.
@@
 This leads to the definition of the Laplace-Beltrami operator for a $C^{2}$-function $f:M\to\mathbb{R}$ as
-(3) $\Delta_{M}f({\boldsymbol{p}})=\nabla_{M}\cdot\nabla_{M}f({\boldsymbol{p}}).$
+$\Delta_{M}f({\boldsymbol{p}})=\nabla_{M}\cdot\nabla_{M}f({\boldsymbol{p}}).$ (3)
 For ${\boldsymbol{f}}:M\to\mathbb{R}^{d}$ and $\mathbb{F}:M\to\mathbb{R}^{d\times d}$ we define $(\nabla_{M}{\boldsymbol{f}})_{ij}=(\nabla({\boldsymbol{f}}\cdot{\boldsymbol{e}}_{i}))_{j}$ and $(\nabla_{M}\cdot\mathbb{F})_{i}=\nabla_{M}\cdot(\mathbb{F}^{T}{\boldsymbol{e}}_{i})=\mathrm{Tr}(\nabla_{M}(\mathbb{F}^{T}{\boldsymbol{e}}_{i}))$, where ${\boldsymbol{e}}_{i},~i\in\{1,\ldots,d\}$ is the canonical basis in $\mathbb{R}^{d}$.
@@
 In order to analyze curvature properties of $M$, we introduce the extended Weingarten map $\mathbb{H}$ as
-(4) $\mathbb{H}({\boldsymbol{p}})=-\nabla_{M}{\boldsymbol{n}}_{M}({\boldsymbol{p}}).$
+$\mathbb{H}({\boldsymbol{p}})=-\nabla_{M}{\boldsymbol{n}}_{M}({\boldsymbol{p}}).$ (4)
 It can be shown that $\mathbb{H}$ is symmetric, has an eigenvalue 0 in the direction of the normal, and restricts to the Weingarten map $\mathbb{W}({\boldsymbol{p}})$ on the tangent space $T_{{\boldsymbol{p}}}M$.
 For ${\boldsymbol{p}}\in M$, the mean curvature is defined as
-(5) $\kappa({\boldsymbol{p}})=\mathrm{Tr}(\mathbb{H}({\boldsymbol{p}})).$
+$\kappa({\boldsymbol{p}})=\mathrm{Tr}(\mathbb{H}({\boldsymbol{p}})).$ (5)
 We recall that all of the above definitions are independent of the extension $\bar{f}$. A sketch of possible model setups is shown in Figure 1.
@@
 In this article an evolving manifold is modeled by a reference manifold $\widehat{M}$ together with a flow map
-(6) ${\boldsymbol{\Phi}}:J\times\widehat{M}\to\mathbb{R}^{d},\quad{\boldsymbol{\Phi}},{\boldsymbol{\Phi}}^{-1}\in C^{1}(J;C^{2}(\mathbb{R}^{d};\mathbb{R}^{d})),$
+${\boldsymbol{\Phi}}:J\times\widehat{M}\to\mathbb{R}^{d},\quad{\boldsymbol{\Phi}},{\boldsymbol{\Phi}}^{-1}\in C^{1}(J;C^{2}(\mathbb{R}^{d};\mathbb{R}^{d})),$ (6)
 such that
@@
   • $^{\boldsymbol{\Phi}}_{0}=\mathbf{Id}_{\widehat{M}}$, where $\mathbf{Id}_{\widehat{M}}$ is the identity map on the reference manifold.
-The mapping ${\boldsymbol{\Phi}}_{t}$ can be used to define pull-back and push-forward maps of functions [4, 3]
-(7) ${\boldsymbol{\Phi}}_{-t}f=f\circ{\boldsymbol{\Phi}}_{t}:\widehat{M}\rightarrow\mathbb{R}\text{ for }f:M(t)\rightarrow\mathbb{R},$
+The mapping ${\boldsymbol{\Phi}}_{t}$ can be used to define pull-back and push-forward maps of functions Alphonse et al. [2015, 2023]
+${\boldsymbol{\Phi}}_{-t}f=f\circ{\boldsymbol{\Phi}}_{t}:\widehat{M}\rightarrow\mathbb{R}\text{ for }f:M(t)\rightarrow\mathbb{R},$ (7)
 ${\boldsymbol{\Phi}}_{t}f=f\circ{\boldsymbol{\Phi}}_{t}^{-1}:M(t)\rightarrow\mathbb{R}\text{ for }f:\widehat{M}\rightarrow\mathbb{R}$.
@@
 In addition, we assume that there exists a velocity field ${\boldsymbol{v}}:J\times\mathbb{R}^{d}\to\mathbb{R}^{d}$ with ${\boldsymbol{v}}\in C^{0}(J;C^{2}(\mathbb{R}^{d};\mathbb{R}^{d}))$ such that for any $t\in J$ and every $^{\boldsymbol{p}}\in\widehat{M}$
-(8) $\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}{\boldsymbol{\Phi}}_{t}({\boldsymbol{p}})={\boldsymbol{v}}(t,{\boldsymbol{\Phi}}_{t}({\boldsymbol{p}})).$
+$\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}{\boldsymbol{\Phi}}_{t}({\boldsymbol{p}})={\boldsymbol{v}}(t,{\boldsymbol{\Phi}}_{t}({\boldsymbol{p}})).$ (8)
 For ${\boldsymbol{x}}\in M(t)$, the velocity $\boldsymbol{v}$} can be split into a tangential component ${\boldsymbol{v}}^{\top}(t,{\boldsymbol{x}})\in T_{{\boldsymbol{x}}}M(t)$ and a normal one ${\boldsymbol{v}}^{\perp}(t,{\boldsymbol{x}})={\boldsymbol{v}}(t,{\boldsymbol{x}})-{\boldsymbol{v}}^{\top}(t,{\boldsymbol{x}})$.
 This allows us to define a normal derivative of a scalar function $f$ on $M(t)$ as
-(9) $\partial^{\circ}f(t,{\boldsymbol{x}})=\partial_{t}\bar{f}(t,{\boldsymbol{x}})+{\boldsymbol{v}}^{\perp}(t,{\boldsymbol{x}})\cdot\nabla\bar{f}(t,{\boldsymbol{x}}),$
+$\partial^{\circ}f(t,{\boldsymbol{x}})=\partial_{t}\bar{f}(t,{\boldsymbol{x}})+{\boldsymbol{v}}^{\perp}(t,{\boldsymbol{x}})\cdot\nabla\bar{f}(t,{\boldsymbol{x}}),$ (9)
 where $\bar{f}$ is a smooth extension of $f$ in a space-time neighborhood of $M(t)$. The material derivative is defined as
-(10) $\partial^{\bullet}f(t,{\boldsymbol{x}})=[\partial^{\circ}f+{\boldsymbol{v}}^{\top}\cdot\nabla_{M(t)}\bar{f}](t,{\boldsymbol{x}})=[\partial_{t}\bar{f}+{\boldsymbol{v}}\cdot\nabla\bar{f}](t,{\boldsymbol{x}}).$
+$\partial^{\bullet}f(t,{\boldsymbol{x}})=[\partial^{\circ}f+{\boldsymbol{v}}^{\top}\cdot\nabla_{M(t)}\bar{f}](t,{\boldsymbol{x}})=[\partial_{t}\bar{f}+{\boldsymbol{v}}\cdot\nabla\bar{f}](t,{\boldsymbol{x}}).$ (10)
 The definition of a flow map is not unique and the same evolving manifold can be described by different maps.
@@
 It can be shown that ${\boldsymbol{v}}^{\mathcal{A}}(t,{\boldsymbol{x}})-{\boldsymbol{v}}(t,{\boldsymbol{x}})\in T_{{\boldsymbol{x}}}M(t)$ and that
-(11) $\partial^{\mathcal{A}}f(t,{\boldsymbol{x}})=[\partial^{\bullet}f+({\boldsymbol{v}}^{\mathcal{A}}-{\boldsymbol{v}})\nabla_{M(t)}f](t,{\boldsymbol{x}}).$
+$\partial^{\mathcal{A}}f(t,{\boldsymbol{x}})=[\partial^{\bullet}f+({\boldsymbol{v}}^{\mathcal{A}}-{\boldsymbol{v}})\nabla_{M(t)}f](t,{\boldsymbol{x}}).$ (11)
 It is worth noting that the normal velocity $\boldsymbol{v}$$^{\perp}$ is in any case uniquely determined, i.e.
@@
 2.1 Space discretization
-The geometries are discretized using piecewise linear elements, where for surfaces we follow the setup of [39].
+The geometries are discretized using piecewise linear elements, where for surfaces we follow the setup of Dziuk and Elliott [2013].
 An $m$-dimensional manifold $M$ is approximated by a triangulated $m$-dimensional domain denoted by $M_{h}$.
@@
 We suppose $M_{h}$ is composed by a collection $\mathcal{T}_{h}$ of $m$-simplices which vertices ${\boldsymbol{y}}_{i},\;i=1,\ldots,N$ lie on $M$ and such that
-(12) $\mathcal{T}_{h}=\{T\},\quad M_{h}=\bigcup T.$
+$\mathcal{T}_{h}=\{T\},\quad M_{h}=\bigcup T.$ (12)
 For each $T\in\mathcal{T}_{h}$ we denote by $h(T)$ its diameter and define the mesh-size as
-(13) $h=\max_{T\in\mathcal{T}_{h}}h(T).$
+$h=\max_{T\in\mathcal{T}_{h}}h(T).$ (13)
 The boundary of every element $T$ is composed by $m+1$ facets of dimension m1m-1, forming a collection $\mathcal{F}=\{F\}$.
@@
 We will again use the notation $\Omega_{h}$ and $\Gamma_{h}$ when we need to distinguish between $d$-dimensional and (d1)(d-1)-dimensional discrete manifolds, respectively.
-Throughout this work, we use a first-order fitted finite element method to discretize the PDEs in space [46].
-In particular, for surfaces, we employ the surface finite element method (SFEM) as presented in [39].
+Throughout this work, we use a first-order fitted finite element method to discretize the PDEs in space Ern and Guermond [2021b].
+In particular, for surfaces, we employ the surface finite element method (SFEM) as presented in Dziuk and Elliott [2013].
 The corresponding finite element space is defined as
-(14) $V_{h}(M_{h}):=\{v_{h}\in C(M_{h})\;|\;v_{h}\rvert_{T}\text{ is linear for all }T\in\mathcal{T}_{h}\},$
+$V_{h}(M_{h}):=\{v_{h}\in C(M_{h})\;|\;v_{h}\rvert_{T}\text{ is linear for all }T\in\mathcal{T}_{h}\},$ (14)
 which can be described as the span of $N$ piecewise linear continuous basis functions $\phi_{i},\;i=1,\ldots,N$ such that $\phi_{i}({\boldsymbol{y}}_{j})=\delta_{ij}$.
 Occasionally, we will also need the space
-(15) $V_{h0}(M_{h}):=\{v_{h}\in C(M_{h})\;|\;v_{h}\rvert_{T}\text{ is linear for all }T\in\mathcal{T}_{h}\text{ and }v_{h}|_{\partial M_{h}}=0\}.$
+$V_{h0}(M_{h}):=\{v_{h}\in C(M_{h})\;|\;v_{h}\rvert_{T}\text{ is linear for all }T\in\mathcal{T}_{h}\text{ and }v_{h}|_{\partial M_{h}}=0\}.$ (15)
 Integration over mesh entities follows the same notation as for the continuous one, where the summation over every element of the collection is implicit, i.e. we write
-(16) $(\cdot\,,\,\cdot)_{M_{h}}=\sum_{T\in M_{h}}(\cdot\,,\,\cdot)_{T},\quad(\cdot\,,\,\cdot)_{\mathcal{F}_{h}^{i}}=\sum_{F\in\mathcal{F}_{h}^{i}}(\cdot\,,\,\cdot)_{F},$
+$(\cdot\,,\,\cdot)_{M_{h}}=\sum_{T\in M_{h}}(\cdot\,,\,\cdot)_{T},\quad(\cdot\,,\,\cdot)_{\mathcal{F}_{h}^{i}}=\sum_{F\in\mathcal{F}_{h}^{i}}(\cdot\,,\,\cdot)_{F},$ (16)
 where (,)($\cdot\,$$\cdot$) is a certain inner product. The same element-wise summation is used in case of discrete norms.
+Remark 1.
+Together with the fitted finite element discretization explained above, an unfitted approach can also be adopted.
+We refer to Burman et al.
+[2015], Palzhanov et al.
+[2021], Wang et al.
+[2022], Olshanskii et al.
+[2022], Neiva and Turlier [2025], Garg and Olshanskii [2025] for advances in this direction.
 2.2 Time discretization
-To discretize the evolving manifold $M(t)$ we use the reference discretization $\widehat{M}_{h}$ and transport it using a suitable ALE map ${\boldsymbol{\Phi}}^{\mathcal{A}}$ [55, 60, 89, 52, 49, 8, 18, 68, 3].
+The evolving manifold $M(t)$ is discretized using the reference discretization $\widehat{M}_{h}$ and transport it using a suitable ALE map ${\boldsymbol{\Phi}}^{\mathcal{A}}$ Hirt et al.
+[1974], Hughes et al.
+[1981], Nobile [2001], Gastaldi [2001], Formaggia and Nobile [2004], Badia and Codina [2006], Bonito et al.
+[2013], Kovács and Power Guerra [2018], Alphonse et al.
+[2023].
 The reference points $\widehat{{\boldsymbol{y}}}_{i}\in\widehat{M}_{h}$ defining the simplices of the reference triangulation $\mathcal{T}_{h}$ are evolved as ${\boldsymbol{y}}_{i}(t)={\boldsymbol{\Phi}}_{t}^{\mathcal{A}}(\widehat{{\boldsymbol{y}}}_{i})$.
 This allows to define evolved elements $T(t)$ and facets $F(t)$ such that
-(17) $\mathcal{T}_{h}(t)=\{T(t)\},\quad\mathcal{F}_{h}(t)=\{F(t)\},\quad M_{h}(t)=\bigcup T(t),$
+$\mathcal{T}_{h}(t)=\{T(t)\},\quad\mathcal{F}_{h}(t)=\{F(t)\},\quad M_{h}(t)=\bigcup T(t),$ (17)
 $\mathcal{F}_{h}^{i}(t)=\{F(t)\in\mathcal{F}_{h}(t):\;F(t)\cap\partial M_{h}(t)=\emptyset\},$
@@
 For the discretization of the function spaces, we start with the finite element space defined in (14), then transport the basis functions using the flow map $\phi_{i}(t)={\boldsymbol{\Phi}}_{t}^{\mathcal{A}}(\phi_{i})$ and define the evolved finite element space as their span
-(18) $V_{h}(t)=\{\mathrm{span}\{\phi_{i}(t)\}:i=1,\ldots,N.\;\phi_{i}\in V_{h}(\widehat{M}_{h})\}.$
-In an analogous way we can also define $V_{h0}(t)$. It follows from the definitions that $V_{h}(0)\equiv V_{h}(\widehat{M}_{h})$. Discrete material velocities are defined by
-(19) ${\boldsymbol{v}}_{h}(t,\cdot)=\sum_{i=1}^{N}{\boldsymbol{v}}(t,{\boldsymbol{y}}_{i}(t))\,\phi_{i}(t),\quad{\boldsymbol{v}}^{\mathcal{A}}_{h}(t,\cdot)=\sum_{i=1}^{N}{\boldsymbol{v}}^{\mathcal{A}}(t,{\boldsymbol{y}}_{i}(t))\,\phi_{i}(t),$
+$V_{h}(t)=\{\mathrm{span}\{\phi_{i}(t)\}:i=1,\ldots,N.\;\phi_{i}\in V_{h}(\widehat{M}_{h})\}.$ (18)
+The space $V_{h0}(t)$ is defined analogously. It follows from the definitions that $V_{h}(0)\equiv V_{h}(\widehat{M}_{h})$. Discrete material velocities are defined by
+${\boldsymbol{v}}_{h}(t,\cdot)=\sum_{i=1}^{N}{\boldsymbol{v}}(t,{\boldsymbol{y}}_{i}(t))\,\phi_{i}(t),\quad{\boldsymbol{v}}^{\mathcal{A}}_{h}(t,\cdot)=\sum_{i=1}^{N}{\boldsymbol{v}}^{\mathcal{A}}(t,{\boldsymbol{y}}_{i}(t))\,\phi_{i}(t),$ (19)
 with corresponding derivatives
-(20) $\partial^{\bullet}_{h}\phi_{i}=\partial_{t}\bar{\phi_{i}}+{\boldsymbol{v}}_{h}\cdot\nabla\bar{\phi_{i}},\quad\partial^{\mathcal{A}}_{h}\phi_{i}=\partial_{t}\bar{\phi_{i}}+{\boldsymbol{v}}^{\mathcal{A}}_{h}\cdot\nabla\bar{\phi_{i}}.$
-Using an ALE map, which describes the motion of the computational domain, we have that $\partial^{\mathcal{A}}_{h}\phi_{i}=0$.
+$\partial^{\bullet}_{h}\phi_{i}=\partial_{t}\bar{\phi_{i}}+{\boldsymbol{v}}_{h}\cdot\nabla\bar{\phi_{i}},\quad\partial^{\mathcal{A}}_{h}\phi_{i}=\partial_{t}\bar{\phi_{i}}+{\boldsymbol{v}}^{\mathcal{A}}_{h}\cdot\nabla\bar{\phi_{i}}.$ (20)
+Using an ALE map, which describes the motion of the computational domain, it holds that $\partial^{\mathcal{A}}_{h}\phi_{i}=0$.
 In this work, time-dependent PDEs are discretized in a collection of time points $\{t_{0},\ldots,t_{N}\}$.
@@
 The time step size is defined as $\tau^{n+1}=t_{n+1}-t_{n}$ and for the sake of simplicity we restrict in what follows to constant time step size $\tau$, with the understanding that every first-order-in-time scheme is amenable to adaptive time stepping.
-Remark 2.1.
+Remark 2.
 From now on, we will often omit the explicit dependence on time, indicating for example $M_{h}(t)$ as simply $M_{h}$ and leaving it to the context to indicate the considered time.
 3 Structure preservation
-In this section we describe how to deal with the challenges highlighted in P.2.
-In the effort of maintaining the interpretability of the considered systems, we focus on bounds and mass preservation.
-The method recently proposed in [30, 29] provides a flexible approach to achieve such goals for a wide variety of finite element formulations on stationary bulk domains.
+In this section we describe how to deal with the second one of the challenges highlighted in subsection 1.1, i.e.
+positivity, bounds and mass preservation in the interest of interpretability.
+The method recently proposed in Cheng and Shen [2022b, a] provides a flexible approach to achieve such goal for a wide variety of finite element formulations on stationary bulk domains.
 Here, we extend those advances to moving surfaces and moving bulk domains.
@@
   • The approach can easily be incorporated as a post-processing step without requiring explicit time integration.
-While explicit time integration as reviewed in [103] can be effective for hyperbolic equations, here we are also interested in parabolic equations.
+While explicit time integration as reviewed in Zhang and Shu [2011] can be effective for hyperbolic equations, here we are also interested in parabolic equations.
   • The simplicity of the algorithm makes it suitable for implementation in legacy codes with negligible overhead computational time with respect to the original non-preserving discretization method.
-Nevertheless, the method enjoys good stability properties and even allows for error analysis for some specific cases, see [29].
+Nevertheless, the method enjoys good stability properties and even allows for error analysis for some specific cases, see Cheng and Shen [2022a].
   • No ad hoc problem reformulation is needed, but the scheme only acts on the nodal values of the solution through a Lagrange multiplier approach.
-Following the presentation in [29], the derivation of the method starts from a nonlinear PDE in the form
-(21) $u_{t}+\mathcal{L}u+\mathcal{N}(u)=0,$
+Following the presentation in Cheng and Shen [2022a], the derivation of the method starts from a nonlinear PDE in the form
+$u_{t}+\mathcal{L}u+\mathcal{N}(u)=0,$ (21)
 where $\mathcal{L}$ is a linear or non-negative operator and $\mathcal{N}(u)$ is a semilinear or quasi-linear operator.
@@
 The corresponding generic spatial discretization is
-(22) $\partial_{t}u_{h}+\mathcal{L}_{h}u_{h}+\mathcal{N}_{h}(u_{h})=0,$
+$\partial_{t}u_{h}+\mathcal{L}_{h}u_{h}+\mathcal{N}_{h}(u_{h})=0,$ (22)
 where we assume $u_{h}\in V_{h}$.
@@
 The problem is then reformulated as
-(23a) $\partial_{t}u_{h}+\mathcal{L}_{h}u_{h}+\mathcal{N}_{h}(u_{h})=\lambda_{h}g^{\prime}(u_{h}),$
-(23b) $\lambda_{h}\geq 0,\quad g(u_{h})\geq 0,\quad\lambda_{h}g(u_{h})=0,$
-where the second line represents the usual Karush-Kuhn-Tucker conditions for constrained optimization [62, 70].
-A core assumption of the scheme presented in [29] is that (23) is satisfied pointwise in a set of points $\Sigma_{h}$ that can be both mesh points or collocation points for $M_{h}$.
-We must then have:
-(24a) $\partial_{t}u_{h}({\boldsymbol{p}})+\mathcal{L}_{h}u_{h}({\boldsymbol{p}})+\mathcal{N}_{h}(u_{h}({\boldsymbol{p}}))=\lambda_{h}({\boldsymbol{p}})g^{\prime}(u_{h}({\boldsymbol{p}})),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$
-(24b) $\lambda_{h}({\boldsymbol{p}})\geq 0,\quad g(u_{h}({\boldsymbol{p}}))\geq 0,\quad\lambda_{h}({\boldsymbol{p}})g(u_{h}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h}.$
+$\partial_{t}u_{h}+\mathcal{L}_{h}u_{h}+\mathcal{N}_{h}(u_{h})=\lambda_{h}g^{\prime}(u_{h}),$ (23a)
+$\lambda_{h}\geq 0,\quad g(u_{h})\geq 0,\quad\lambda_{h}g(u_{h})=0,$ (23b)
+where the second line represents the usual Karush-Kuhn-Tucker conditions for constrained optimization Karush [1939], Kuhn and Tucker [1951].
+A core assumption of the scheme presented in Cheng and Shen [2022a] is that (23) is satisfied pointwise in a set of points $\Sigma_{h}$ that can be both mesh points or collocation points for $M_{h}$.
+It must then hold:
+$\partial_{t}u_{h}({\boldsymbol{p}})+\mathcal{L}_{h}u_{h}({\boldsymbol{p}})+\mathcal{N}_{h}(u_{h}({\boldsymbol{p}}))=\lambda_{h}({\boldsymbol{p}})g^{\prime}(u_{h}({\boldsymbol{p}})),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$ (24a)
+$\lambda_{h}({\boldsymbol{p}})\geq 0,\quad g(u_{h}({\boldsymbol{p}}))\geq 0,\quad\lambda_{h}({\boldsymbol{p}})g(u_{h}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h}.$ (24b)
 It is important to notice that $\Sigma_{h}$ does not include points where essential boundary conditions are applied.
-The next step is to discretize (24) in time, where we here focus on the case of Backward-Euler (BE or BDF-1) time integration in the interval $\{t^{n},t^{n+1}\}$.
-Higher order extensions are presented in [29] and can be adapted following similar arguments.
+The next step is to discretize (24) in time, where we here focus on the case of Backward-Euler (BE or BDF-1) time integration in the interval $t^{n},t^{n+1}]$.
+Higher order extensions are presented in Cheng and Shen [2022a] and can be adapted following similar arguments.
 The key idea is to apply an operator splitting approach to (23) dividing the problem in two steps.
 In the predictor step we solve the unconstrained problem
-(25) $\frac{\tilde{u}_{h}^{n+1}({\boldsymbol{p}})-u_{h}^{n}({\boldsymbol{p}})}{\tau}+\mathcal{L}_{h}\tilde{u}_{h}^{n+1}({\boldsymbol{p}})+\mathcal{N}_{h}(\tilde{u}_{h}^{n+1}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h}.$
+$\frac{\tilde{u}_{h}^{n+1}({\boldsymbol{p}})-u_{h}^{n}({\boldsymbol{p}})}{\tau}+\mathcal{L}_{h}\tilde{u}_{h}^{n+1}({\boldsymbol{p}})+\mathcal{N}_{h}(\tilde{u}_{h}^{n+1}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h}.$ (25)
 In the corrector step we solve the constrained problem
-(26a) $\frac{u_{h}^{n+1}({\boldsymbol{p}})-\tilde{u}_{h}^{n+1}({\boldsymbol{p}})}{\tau}=\lambda_{h}^{n+1}({\boldsymbol{p}})g^{\prime}(u_{h}^{n+1}({\boldsymbol{p}})),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$
-(26b) $\lambda_{h}^{n+1}({\boldsymbol{p}})\geq 0,\quad g(u_{h}^{n+1}({\boldsymbol{p}}))\geq 0,\quad\lambda_{h}^{n+1}({\boldsymbol{p}})g(u_{h}^{n+1}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h}.$
-If we are only interested in bound preservation, it turns out that the solution to (26) is the cutoff function
-(27) $u_{h}^{n+1}=\text{cutoff}[\tilde{u}_{h}^{n+1}]=\begin{cases}\tilde{u}_{h}^{n+1}({\boldsymbol{p}}),&\text{ if }a<\tilde{u}_{h}^{n+1}({\boldsymbol{p}})<b,\\ a,&\text{ if }\tilde{u}_{h}^{n+1}({\boldsymbol{p}})<a,\\ b,&\text{ if }\tilde{u}_{h}^{n+1}({\boldsymbol{p}})>b.\end{cases}$
+$\frac{u_{h}^{n+1}({\boldsymbol{p}})-\tilde{u}_{h}^{n+1}({\boldsymbol{p}})}{\tau}=\lambda_{h}^{n+1}({\boldsymbol{p}})g^{\prime}(u_{h}^{n+1}({\boldsymbol{p}})),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$ (26a)
+$\lambda_{h}^{n+1}({\boldsymbol{p}})\geq 0,\quad g(u_{h}^{n+1}({\boldsymbol{p}}))\geq 0,\quad\lambda_{h}^{n+1}({\boldsymbol{p}})g(u_{h}^{n+1}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h}.$ (26b)
+If one is only interested in bound preservation, it turns out that the solution to (26) is the cutoff function
+$u_{h}^{n+1}=\text{cutoff}[\tilde{u}_{h}^{n+1}]=\begin{cases}\tilde{u}_{h}^{n+1}({\boldsymbol{p}}),&\text{ if }a<\tilde{u}_{h}^{n+1}({\boldsymbol{p}})<b,\\ a,&\text{ if }\tilde{u}_{h}^{n+1}({\boldsymbol{p}})<a,\\ b,&\text{ if }\tilde{u}_{h}^{n+1}({\boldsymbol{p}})>b.\end{cases}$ (27)
 The above step is unfortunately not mass-preserving even if the solution is.
-To fix this, the authors in [30, 29] introduce an additional global space-independent Lagrange multiplier $\xi_{h}^{n+1}$.
+To fix this, the authors in Cheng and Shen [2022b, a] introduce an additional global space-independent Lagrange multiplier $\xi_{h}^{n+1}$.
 The corrector step (26) is modified to include the mass preservation constraint,
-(28a) $\frac{1}{\tau}u_{h}^{n+1}({\boldsymbol{p}})=\lambda_{h}^{n+1}({\boldsymbol{p}})g^{\prime}(u_{h}^{n+1}({\boldsymbol{p}}))+\xi_{h}^{n+1}+\frac{1}{\tau}\tilde{u}_{h}^{n+1}({\boldsymbol{p}}),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$
-(28b) $\lambda_{h}^{n+1}({\boldsymbol{p}})\geq 0,\quad g(u_{h}^{n+1}({\boldsymbol{p}}))\geq 0,\quad\lambda_{h}^{n+1}({\boldsymbol{p}})g(u_{h}^{n+1}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h},$
+$\frac{1}{\tau}u_{h}^{n+1}({\boldsymbol{p}})=\lambda_{h}^{n+1}({\boldsymbol{p}})g^{\prime}(u_{h}^{n+1}({\boldsymbol{p}}))+\xi_{h}^{n+1}+\frac{1}{\tau}\tilde{u}_{h}^{n+1}({\boldsymbol{p}}),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$ (28a)
+$\lambda_{h}^{n+1}({\boldsymbol{p}})\geq 0,\quad g(u_{h}^{n+1}({\boldsymbol{p}}))\geq 0,\quad\lambda_{h}^{n+1}({\boldsymbol{p}})g(u_{h}^{n+1}({\boldsymbol{p}}))=0,\;\forall{\boldsymbol{p}}\in\Sigma_{h},$ (28b)
 […]
@@
 In our case this discrete inner product, that has to be expressed pointwise as $(u\,,\,v)^{h}=\sum_{{\boldsymbol{p}}\in\Sigma_{h}}\omega_{\boldsymbol{p}}u({\boldsymbol{p}})v({\boldsymbol{p}})$, is assumed to be equivalent to the classical $L^{2}$ inner product on $V_{h}$.
-For this reason, whenever required, we will choose $(\cdot\,,\,\cdot)^{h}$ to be the mass-lumped inner product.
+For this reason, whenever required, $(\cdot\,,\,\cdot)^{h}$ is chosen to be the mass-lumped inner product.
 Rearranging (28a) as
-(29) $\frac{1}{\tau}(u_{h}^{n+1}({\boldsymbol{p}})-(\tilde{u}_{h}^{n+1}({\boldsymbol{p}})+\tau\xi_{h}^{n+1}))=\lambda_{h}^{n+1}({\boldsymbol{p}})g^{\prime}(u_{h}^{n+1}({\boldsymbol{p}})),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$
+$\frac{1}{\tau}(u_{h}^{n+1}({\boldsymbol{p}})-(\tilde{u}_{h}^{n+1}({\boldsymbol{p}})+\tau\xi_{h}^{n+1}))=\lambda_{h}^{n+1}({\boldsymbol{p}})g^{\prime}(u_{h}^{n+1}({\boldsymbol{p}})),\;\forall{\boldsymbol{p}}\in\Sigma_{h},$ (29)
 the solution is then $u_{h}^{n+1}=\text{cutoff}[\tilde{u}_{h}^{n+1}+\tau\xi_{h}^{n+1}]$ where $\xi_{h}^{n+1}$ is the solution to the nonlinear equation
-(30) $F^{n+1}(\xi_{h}^{n+1})=(\text{cutoff}[\tilde{u}_{h}^{n+1}+\tau\xi_{h}^{n+1}]\,,\,1)^{h}-(u_{h}^{n}\,,\,1)^{h}=0.$
-Bounds on the energy norm are provided for a class of non-linear PDEs in [29] together with error analysis of a restricted class of problems.
-Remark 3.1.
-Since $(F^{n+1})^{\prime}$ might not exist, the authors of the scheme in [30, 29] suggest the use of a secant method to find the solution of $F^{n+1}(\xi)=0$ in (30) using
-(31) $\xi_{k+1}=\xi_{k}-\frac{F^{n+1}(\xi_{k})(\xi_{k}-\xi_{k-1})}{F^{n+1}(\xi_{k})-F^{n+1}(\xi_{k-1})},$
+$F^{n+1}(\xi_{h}^{n+1})=(\text{cutoff}[\tilde{u}_{h}^{n+1}+\tau\xi_{h}^{n+1}]\,,\,1)^{h}-(u_{h}^{n}\,,\,1)^{h}=0.$ (30)
+Bounds on the energy norm are provided for a class of non-linear PDEs in Cheng and Shen [2022a] together with error analysis of a restricted class of problems.
+Remark 3.
+Since $(F^{n+1})^{\prime}$ might not exist, the authors of the scheme in Cheng and Shen [2022b, a] suggest the use of a secant method to find the solution of $F^{n+1}(\xi)=0$ in (30) using
+$\xi_{k+1}=\xi_{k}-\frac{F^{n+1}(\xi_{k})(\xi_{k}-\xi_{k-1})}{F^{n+1}(\xi_{k})-F^{n+1}(\xi_{k-1})},$ (31)
 together with initial guesses $\xi_{0}=0$ and $\xi_{1}=\mathcal{O}(\tau)$. From a practical standpoint, the algorithm converged in less than four iterations for each timestep for the case at hand.
 The technique introduced is clearly directly applicable to a wide variety of equations.
-The authors themselves in [30, 29] have proposed experiments for Allen-Cahn, Cahn-Hilliard with variable mobility and Fokker-Planck equations on a fixed bulk domain.
+The authors themselves in Cheng and Shen [2022b, a] have proposed experiments for Allen-Cahn, Cahn-Hilliard with variable mobility and Fokker-Planck equations on a fixed bulk domain.
 Inspired by the biophysical mechanisms driving cell reshaping, we will demonstrate their accuracy when modeling complex advection-diffusion-reaction equations and phase-separation phenomena.
@@
 4 Tangential grid control and ALE discretization
-When dealing with evolving domains in a fitted framework, it is not uncommon to incur deformations that lead to highly distorted meshes.
+When dealing with evolving domains in a fitted framework, it is not uncommon to incur deformations that lead to highly distorted meshes, as mentioned in subsection 1.1.
 This is especially true when gradient flows are involved, since they only prescribe the velocity and hence the displacement in normal direction.
-Following 11, we introduce a particular ALE map tailored for gradient flow dynamics that works as follows.
-Consider the time interval $\{t_{n},t_{n+1}\}$, and suppose the surface $\Gamma_{h}^{n}$ would be displaced, following the physics of the problem, to an updated surface $\widetilde{\Gamma}_{h}^{n+1}$ at time $t_{n+1}$ which is highly distorted.
+Following (11), we introduce a particular ALE map tailored for gradient flow dynamics that works as follows.
+Consider the time interval $t_{n},t_{n+1}]$, and suppose the surface $\Gamma_{h}^{n}$ would be displaced, following the physics of the problem, to an updated surface $\widetilde{\Gamma}_{h}^{n+1}$ at time $t_{n+1}$ which is highly distorted.
 We define an artificial tangential motion aimed at maintaining the surface shape (and consequently energy) but redistributing the nodes more favorably to reduce mesh distortions.
-We choose the two-stage algorithm proposed in [36] for the following reasons:
+We choose the two-stage algorithm proposed in Duan and Li [2024] for the following reasons:
   • The method has proven to be very effective for gradient flows such as the mean curvature flow and the surface diffusion flow.
@@
   • Beyond the scope of the current work, the method can potentially be applied to higher order in time integration schemes.
-The algorithm presented in [36] is based on the assumption that the starting mesh is regular enough.
+The algorithm presented in Duan and Li [2024] is based on the assumption that the starting mesh is regular enough.
 The idea is then to impose an artificial tangential velocity that requires ${\boldsymbol{\Phi}}^{\mathcal{A}}(t,\cdot):\widehat{\Gamma}\to\Gamma$ to be an harmonic map, i.e.
-(32) $\begin{cases}\frac{\partial{\boldsymbol{\Phi}}^{\mathcal{A}}}{\partial t}\cdot({\boldsymbol{n}}_{\Gamma}\circ{\boldsymbol{\Phi}}^{\mathcal{A}})=({\boldsymbol{v}}\cdot{\boldsymbol{n}})\circ{\boldsymbol{\Phi}}^{\mathcal{A}},\\ -\Delta_{\,\widehat{\Gamma}}{\boldsymbol{\Phi}}^{\mathcal{A}}=(\varkappa{\boldsymbol{n}})\circ{\boldsymbol{\Phi}}^{\mathcal{A}}.\end{cases}$
+$\begin{cases}\frac{\partial{\boldsymbol{\Phi}}^{\mathcal{A}}}{\partial t}\cdot({\boldsymbol{n}}_{\Gamma}\circ{\boldsymbol{\Phi}}^{\mathcal{A}})=({\boldsymbol{v}}\cdot{\boldsymbol{n}})\circ{\boldsymbol{\Phi}}^{\mathcal{A}},\\ -\Delta_{\,\widehat{\Gamma}}{\boldsymbol{\Phi}}^{\mathcal{A}}=(\varkappa{\boldsymbol{n}})\circ{\boldsymbol{\Phi}}^{\mathcal{A}}.\end{cases}$ (32)
 Equivalently, ${\boldsymbol{\Phi}}^{\mathcal{A}}$ is required to minimize the energy $\int_{\,\widehat{\Gamma}}|\nabla_{\widehat{M}}{\boldsymbol{\Phi}}^{\mathcal{A}}(t,\cdot)|^{2}$, under the constraint ${\boldsymbol{v}}\cdot{\boldsymbol{n}}_{\Gamma}=0$ on $\Gamma$, which is imposed using the scalar-valued Lagrange multiplier $\varkappa$.
-The fully discrete version of the scheme reads: given a surface $\widetilde{\Gamma}_{h}^{n+1}$ evolved from an initial surface $\widehat{\Gamma}_{h}$, find ${\boldsymbol{w}}_{h}^{n+1}\in[V_{h0}(\widetilde{\Gamma}^{n+1}_{h})]^{d}$ and $\varkappa_{h}^{n+1}\in V_{h0}(\widetilde{\Gamma}^{n+1}_{h})$ such that
+The fully discrete version of the scheme reads: given a surface $\widetilde{\Gamma}_{h}^{n+1}$ evolved from an initial reference surface $\widehat{\Gamma}_{h}$, find ${\boldsymbol{w}}_{h}^{n+1}\in[V_{h0}(\widetilde{\Gamma}^{n+1}_{h})]^{d}$ and $\varkappa_{h}^{n+1}\in V_{h0}(\widetilde{\Gamma}^{n+1}_{h})$ such that
 $(\nabla_{\widehat{\Gamma}_{h}}({\boldsymbol{w}}_{h}^{n+1}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{t_{n+1}})\,,\,\nabla_{\widehat{\Gamma}_{h}}({\boldsymbol{\phi}}_{h}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{t_{n+1}}))_{\widehat{\Gamma}_{h}}-(\varkappa_{h}^{n+1}{\boldsymbol{n}}_{\widetilde{\Gamma}_{h}^{n+1}}\,,\,{\boldsymbol{\phi}}_{h})_{\widetilde{\Gamma}_{h}^{n+1}}$
-(33a) $\qquad=-(\nabla_{\widehat{\Gamma}_{h}}(\mathbf{Id}_{\widetilde{\Gamma}_{h}^{n+1}}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{t_{n+1}})\,,\,\nabla_{\widehat{\Gamma}_{h}}({\boldsymbol{\phi}}_{h}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{t_{n+1}}))_{\widehat{\Gamma}_{h}}$
-(33b) $-({\boldsymbol{w}}_{h}^{n+1}\cdot{\boldsymbol{n}}_{\widetilde{\Gamma}_{h}^{n+1}}\,,\,\psi_{h})_{\widetilde{\Gamma}_{h}^{n+1}}=0,$
+$\qquad=-(\nabla_{\widehat{\Gamma}_{h}}(\mathbf{Id}_{\widetilde{\Gamma}_{h}^{n+1}}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{t_{n+1}})\,,\,\nabla_{\widehat{\Gamma}_{h}}({\boldsymbol{\phi}}_{h}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{t_{n+1}}))_{\widehat{\Gamma}_{h}}$ (33a)
+$-({\boldsymbol{w}}_{h}^{n+1}\cdot{\boldsymbol{n}}_{\widetilde{\Gamma}_{h}^{n+1}}\,,\,\psi_{h})_{\widetilde{\Gamma}_{h}^{n+1}}=0,$ (33b)
 for all ${\boldsymbol{\phi}}_{h}\in[V_{h0}(\widetilde{\Gamma}^{n+1}_{h})]^{d}$ and $\psi_{h}\in V_{h0}(\widetilde{\Gamma}^{n+1}_{h})$.
@@
 From this we have that $({\boldsymbol{v}}_{h}-{\boldsymbol{v}}_{h}^{\mathcal{A}})^{n+1}=-{\boldsymbol{w}}_{h}^{n+1}/\tau$.
+Remark 4.
+The algorithm proposed in (33) does not constitute the only option available.
+Since (33) is a post-processing step, alternative schemes can be used as long as the mesh redistribution is purely tangential Sauer [2014], Hu and Li [2022].
+To highlight the flexibility of our framework we further introduce the scheme in Hu and Li [2022], which closely resembles the one presented in Duan and Li [2024].
+The fully discrete version of the modified scheme reads: given a surface $\widetilde{\Gamma}_{h}^{n+1}$ evolved from a surface $\Gamma^{n}_{h}$ at the previous timestep, find ${\boldsymbol{w}}_{h}^{n+1}\in[V_{h0}(\widetilde{\Gamma}^{n+1}_{h})]^{d}$ and $\varkappa_{h}^{n+1}\in V_{h0}(\widetilde{\Gamma}^{n+1}_{h})$ such that
+$\frac{1}{\tau}(\nabla_{\Gamma^{n}_{h}}({\boldsymbol{w}}_{h}^{n+1}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{\tau})\,,\,\nabla_{\Gamma^{n}_{h}}({\boldsymbol{\phi}}_{h}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{\tau}))_{\Gamma^{n}_{h}}-(\varkappa_{h}^{n+1}{\boldsymbol{n}}_{\widetilde{\Gamma}_{h}^{n+1}}\,,\,{\boldsymbol{\phi}}_{h})_{\widetilde{\Gamma}_{h}^{n+1}}$ $=0$ (34a)
+$-({\boldsymbol{w}}_{h}^{n+1}\cdot{\boldsymbol{n}}_{\widetilde{\Gamma}_{h}^{n+1}}\,,\,\psi_{h})_{\widetilde{\Gamma}_{h}^{n+1}}$ $=0,$ (34b)
+for all ${\boldsymbol{\phi}}_{h}\in[V_{h0}(\widetilde{\Gamma}^{n+1}_{h})]^{d}$ and $\psi_{h}\in V_{h0}(\widetilde{\Gamma}^{n+1}_{h})$, where ${\boldsymbol{\Phi}}^{\mathcal{A}}_{\tau}={\boldsymbol{\Phi}}^{\mathcal{A}}_{t_{n+1}}\circ{\boldsymbol{\Phi}}^{\mathcal{A}}_{-t_{n}}$.
 After the surface has been displaced, a second step might be needed to extend the ALE motion in the bulk.
@@
 The Laplace problem reads correspondingly
-(34) $\Delta{\boldsymbol{e}}_{h}^{n+1}$ $=0,\qquad\text{ in }\widetilde{\Omega}^{n+1}_{h},$
-(35) ${\boldsymbol{e}}_{h}^{n+1}$ $={\boldsymbol{w}}_{h}^{n+1},\quad\text{ on }\partial\widetilde{\Omega}^{n+1}_{h},$
+$\Delta{\boldsymbol{e}}_{h}^{n+1}$ $=0,\qquad\text{ in }\widetilde{\Omega}^{n+1}_{h},$ (35)
+${\boldsymbol{e}}_{h}^{n+1}$ $={\boldsymbol{w}}_{h}^{n+1},\quad\text{ on }\partial\widetilde{\Omega}^{n+1}_{h},$ (36)
 and we have that $({\boldsymbol{v}}_{h}-{\boldsymbol{v}}_{h}^{\mathcal{A}})^{n+1}=-{\boldsymbol{e}}_{h}^{n+1}/\tau$.
@@
 […]
-where $u:\mathcal{G}_{T}\to\mathbb{R}$ is the concentration, ${\boldsymbol{b}}_{M}$ is the advective velocity field, $\mathbb{A}_{M}$ is the diffusion matrix and $c_{M}>0$ is the reaction constant [52, 39, 42].
-As in [39, p.304] we will assume that $(\mathbb{A}_{M})_{ij},\;({\boldsymbol{b}}_{M})_{i},\;\nabla_{M}\cdot{\boldsymbol{b}}_{M},\;c_{M}\in L^{\infty}(M),$ together with ${\boldsymbol{b}}_{M}({\boldsymbol{x}})\in T_{{\boldsymbol{x}}}M,\;\forall{\boldsymbol{x}}\in M$ and the requirement for $\mathbb{A}_{M}$ to be symmetric and to map the tangent space $T_{{\boldsymbol{x}}}M$ onto itself.
-Extensive literature exists on finite element discretizations of such equations for what concerns the bulk case, see [46, 45].
-For closed surfaces the SFEM discretization of parabolic PDEs in the form of (36) has been comprehensively reviewed in [39].
+where $u:\mathcal{G}_{T}\to\mathbb{R}$ is the concentration, ${\boldsymbol{b}}_{M}$ is the advective velocity field, $\mathbb{A}_{M}$ is the diffusion matrix and $c_{M}>0$ is the reaction constant Gastaldi [2001], Dziuk and Elliott [2013], Elliott and Ranner [2021].
+As in [Dziuk and Elliott, 2013, p.304] it is assumed that $(\mathbb{A}_{M})_{ij},\;({\boldsymbol{b}}_{M})_{i},\;\nabla_{M}\cdot{\boldsymbol{b}}_{M},\;c_{M}\in L^{\infty}(M),$ together with ${\boldsymbol{b}}_{M}({\boldsymbol{x}})\in T_{{\boldsymbol{x}}}M,\;\forall{\boldsymbol{x}}\in M$ and the requirement for $\mathbb{A}_{M}$ to be symmetric and to map the tangent space $T_{{\boldsymbol{x}}}M$ onto itself.
+Extensive literature exists on finite element discretizations of such equations for what concerns the bulk case, see Ern and Guermond [2021b, a].
+For closed surfaces the SFEM discretization of parabolic PDEs in the form of (37) has been comprehensively reviewed in Dziuk and Elliott [2013].
 Convergence estimates for its continuous-in-time formulation using general order polynomials are found in e.g.
-[42].
+Elliott and Ranner [2021].
 For the sake of generality, here we focus on moving surfaces with boundary.
 The problem reads: Find $u_{h}\in V_{h}(\Gamma_{h})$ such that
-(37) $\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}m_{h}(u_{h}\,,\,v_{h})+\overline{a}_{h}(u_{h}\,,\,v_{h})+(({\boldsymbol{b}}_{rel}u_{h}-\mathbb{A}_{\Gamma_{h}}\nabla_{\Gamma_{h}}u_{h})\cdot{\boldsymbol{n}}_{\partial\Gamma_{h}}\,,\,v_{h})_{\partial\Gamma_{h}}$
+$\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}m_{h}(u_{h}\,,\,v_{h})+\overline{a}_{h}(u_{h}\,,\,v_{h})+(({\boldsymbol{b}}_{rel}u_{h}-\mathbb{A}_{\Gamma_{h}}\nabla_{\Gamma_{h}}u_{h})\cdot{\boldsymbol{n}}_{\partial\Gamma_{h}}\,,\,v_{h})_{\partial\Gamma_{h}}$ (38)
 $=\overline{l}_{h}(v_{h})\;,\forall v_{h}\in V_{h}(\Gamma_{h}).$
 where
-(38) $m_{h}(u_{h},v_{h})=$ $~(u_{h}\,,\,v_{h})_{\Gamma_{h}},$
+$m_{h}(u_{h},v_{h})=$ $~(u_{h}\,,\,v_{h})_{\Gamma_{h}},$ (39)
 $\overline{a}_{h}(u_{h},v_{h})=$ $~(\mathbb{A}_{M_{h}}\nabla_{\Gamma_{h}}u_{h}\,,\,\nabla_{\Gamma_{h}}v_{h})_{\Gamma_{h}}+(-{\boldsymbol{b}}_{rel}u_{h}\,,\,\nabla_{\Gamma_{h}}v_{h})_{\Gamma_{h}}$
@@
 It is required that
-(39) $\partial\Gamma_{h,D}\cap\partial\Gamma_{h,N}=\emptyset,\quad\partial\Gamma_{h,D}\cup\partial\Gamma_{h,N}=\partial\Gamma_{h}.$
-Nitsche’s boundary penalty method is used in order to impose the inhomogeneous Dirichlet boundary condition $u_{h}(t,{\boldsymbol{x}})|_{\partial\Gamma_{h,D}}=u(t,{\boldsymbol{x}})$ weakly for the diffusive term [87, 45].
+$\partial\Gamma_{h,D}\cap\partial\Gamma_{h,N}=\emptyset,\quad\partial\Gamma_{h,D}\cup\partial\Gamma_{h,N}=\partial\Gamma_{h}.$ (40)
+Nitsche’s boundary penalty method is used in order to impose the inhomogeneous Dirichlet boundary condition $u_{h}(t,{\boldsymbol{x}})|_{\partial\Gamma_{h,D}}=u(t,{\boldsymbol{x}})$ weakly for the diffusive term Nitsche [1971], Ern and Guermond [2021a].
 For convenience, we define $v^{+}=\max\{v,0\}$ and $v^{-}=\min\{v,0\}$.
 The resulting bilinear forms are given by
-(40) $a_{h}(u_{h}\,,\,v_{h})=$ $~\overline{a}_{h}(u_{h}\,,\,v_{h})-((\mathbb{A}_{\Gamma_{h}}\nabla_{\Gamma_{h}}u_{h})\cdot{\boldsymbol{n}}_{\partial\Gamma_{h}}\,,\,v_{h})_{\partial\Gamma_{h,D}}$
+$a_{h}(u_{h}\,,\,v_{h})=$ $~\overline{a}_{h}(u_{h}\,,\,v_{h})-((\mathbb{A}_{\Gamma_{h}}\nabla_{\Gamma_{h}}u_{h})\cdot{\boldsymbol{n}}_{\partial\Gamma_{h}}\,,\,v_{h})_{\partial\Gamma_{h,D}}$ (41)
 $-(u_{h}\,,\,(\mathbb{A}_{\Gamma_{h}}\nabla_{\Gamma_{h}}v_{h})\cdot{\boldsymbol{n}}_{\partial\Gamma_{h}})_{\partial\Gamma_{h,D}}+(\gamma_{\mathbb{A}}h^{-1}_{F}u_{h}\,,\,v_{h})_{\partial\Gamma_{h,D}}$
@@
 where $\gamma_{\mathbb{A}}>0$ is the penalty parameter enforcing $u_{h}=u$ on $\partial\Gamma_{h,D}$. Overall the problem for open surfaces reads: find $u_{h}\in V_{h}$ such that
-(41) $\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}m_{h}(u_{h}\,,\,v_{h})+a_{h}(u_{h}\,,\,v_{h})=l_{h}(v_{h}),\;\forall v_{h}\in V_{h}(\Gamma_{h}).$
+$\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}m_{h}(u_{h}\,,\,v_{h})+a_{h}(u_{h}\,,\,v_{h})=l_{h}(v_{h}),\;\forall v_{h}\in V_{h}(\Gamma_{h}).$ (42)
 5.1 Numerical methods for advection-dominant problems
-In biophysical cell models, as introduced in P.1, it is not uncommon to incur advection-dominant ADR equations.
+In biophysical cell models, as introduced in subsection 1.1, it is not uncommon to incur advection-dominant ADR equations.
 In our case, even purely parabolic equations might become advection-dominant due to the introduction of an ALE velocity ${\boldsymbol{v}}^{\mathcal{A}}$.
 In the present work we employ and extend the continuous interior penalty (CIP) stabilization to cope with the possibly dominant advection regime.
-The CIP method was proposed and analyzed in [21, 20] and then later extended to the case of stationary surfaces in [23].
-Moreover, the CutFEM approach for moving surfaces developed in [53] applies a CIP-type stabilization for advection dominant problems on moving surfaces.
+The CIP method was proposed and analyzed in Burman and Hansbo [2004], Burman and Fernández [2009] and then later extended to the case of stationary surfaces in Burman et al.
+[2019].
+Moreover, the CutFEM approach for moving surfaces developed in Hansbo et al.
+[2015] applies a CIP-type stabilization for advection dominant problems on moving surfaces.
 We chose CIP for the following reasons:
   • It is a widely known, easily implementable technique whose implementation tools are usually shipped in classical finite element packages;
-  • Although only weakly consistent, it has proven to be very successful in time-dependent problems given its commutativity with the time derivative [47];
+  • Although only weakly consistent, it has proven to be very successful in time-dependent problems given its commutativity with the time derivative Ern and Guermond [2021c];
   • Therefore, it has been shown to lead to convergent algorithms for discretizations of bulk, surface and implicitly described moving surfaces problems.
@@
 For a piecewise discontinuous function $f$ defined on a surface or bulk mesh $\mathcal{T}_{h}(M_{h})$, we define its average and jump over an interior facet $F\in\mathcal{F}_{h}^{i}(M_{h})$ by
-(42) $\{f\}|_{F}=\frac{1}{2}\left(f_{F}^{+}+f_{F}^{-}\right),\quad{\llbracket f\rrbracket|_{F}}=f_{F}^{+}-f_{F}^{-},$
+$\{f\}|_{F}=\frac{1}{2}\left(f_{F}^{+}+f_{F}^{-}\right),\quad{\llbracket f\rrbracket|_{F}}=f_{F}^{+}-f_{F}^{-},$ (43)
 respectively, where $f_{F}^{\pm}({\boldsymbol{x}})=\lim_{\delta\to 0^{+}}f({\boldsymbol{x}}_{F}^{\pm}-\delta{\boldsymbol{n}}_{F}^{\pm})$.
 The main idea of the CIP approach is to penalize the jump of the streamline derivative across element interfaces. To this end, the CIP stabilization form is defined as follows,
-(43) $s_{h}^{b}\left({u}_{h},{v}_{h}\right):=\gamma_{b}\sum_{F\in\mathcal{F}_{h}^{i}}\phi_{F}h_{F}\left(\llbracket{\boldsymbol{b}_{M_{h}}}\cdot\nabla_{M_{h}}{u}_{h}\rrbracket,\llbracket\boldsymbol{b}_{M_{h}}\cdot\nabla_{M_{h}}{v}_{h}\rrbracket\right)_{F},$
+$s_{h}^{b}\left({u}_{h},{v}_{h}\right):=\gamma_{b}\sum_{F\in\mathcal{F}_{h}^{i}}\phi_{F}h_{F}\left(\llbracket{\boldsymbol{b}_{M_{h}}}\cdot\nabla_{M_{h}}{u}_{h}\rrbracket,\llbracket\boldsymbol{b}_{M_{h}}\cdot\nabla_{M_{h}}{v}_{h}\rrbracket\right)_{F},$ (44)
 where $\gamma_{b}>0$ is a dimensionless stabilization constant and $\phi_{F}$ denotes a stabilization parameter defined by
-(44) $\phi_{F}=\max(\phi_{T^{+}},\phi_{T^{-}}),\quad\phi_{T^{\pm}}=\min(\beta_{T^{\pm}}^{-1}h_{T^{\pm}},\mu_{T^{\pm}}^{-1}),\quad T^{\pm}\in\mathcal{T}_{h}.$
-In the above $\beta_{T}$ is a local velocity scale and $\mu_{T}$ is the reciprocal of a time [47]. If ${\boldsymbol{b}}_{M_{h}}\in L^{\infty}(M_{h})\cap C^{0}(M_{h})$, then 43 can be simplified to
-(45) $s_{h}^{b}({u}_{h},{v}_{h}):=\gamma_{b}\sum_{F\in\mathcal{F}_{h}^{i}}\beta_{F}h_{F}^{2}\left(\llbracket\nabla_{M_{h}}{u}_{h}\rrbracket,\llbracket\nabla_{M_{h}}{v}_{h}\rrbracket\right)_{F}.$
+$\phi_{F}=\max(\phi_{T^{+}},\phi_{T^{-}}),\quad\phi_{T^{\pm}}=\min(\beta_{T^{\pm}}^{-1}h_{T^{\pm}},\mu_{T^{\pm}}^{-1}),\quad T^{\pm}\in\mathcal{T}_{h}.$ (45)
+In the above $\beta_{T}$ is a local velocity scale and $\mu_{T}$ is the reciprocal of a time Ern and Guermond [2021c].
+If ${\boldsymbol{b}}_{M_{h}}\in L^{\infty}(M_{h})\cap C^{0}(M_{h})$, then (44) can be simplified to
+$s_{h}^{b}({u}_{h},{v}_{h}):=\gamma_{b}\sum_{F\in\mathcal{F}_{h}^{i}}\beta_{F}h_{F}^{2}\left(\llbracket\nabla_{M_{h}}{u}_{h}\rrbracket,\llbracket\nabla_{M_{h}}{v}_{h}\rrbracket\right)_{F}.$ (46)
 with $\beta_{F}=\max(\beta_{T^{+}},\beta_{T^{-}})$. We augment the system in (42) further to handle advection-dominant problems, which now reads: find $u_{h}\in V_{h}$ such that
-(46) $\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}m_{h}(u_{h}\,,\,v_{h})+a_{h}(u_{h}\,,\,v_{h})+s_{h}^{b}(u_{h}\,,\,v_{h})=l_{h}(v_{h}),\;\forall v_{h}\in V_{h}(\Gamma_{h}).$
+$\frac{\mathop{}\!\textup{d}}{\mathop{}\!\textup{d}t}m_{h}(u_{h}\,,\,v_{h})+a_{h}(u_{h}\,,\,v_{h})+s_{h}^{b}(u_{h}\,,\,v_{h})=l_{h}(v_{h}),\;\forall v_{h}\in V_{h}(\Gamma_{h}).$ (47)
 5.2 Numerical results for the ADR system
@@
 We begin by considering a domain in 3D that moves under the linear transformation ${\boldsymbol{\Phi}}(t,{\boldsymbol{p}})={\boldsymbol{\Phi}}^{\mathcal{A}}(t,{\boldsymbol{p}})=\mathbb{A}(t){\boldsymbol{p}}+B(t)$ where
-(47) $\mathbb{A}(t)=\begin{bmatrix}\cos(t)&-\sin(t)&0\\ \sin(t)&\cos(t)&0\\ 0&0&1\end{bmatrix},\quad B(t)=\begin{bmatrix}0\\ 0\\ 0\end{bmatrix}$.
+$\mathbb{A}(t)=\begin{bmatrix}\cos(t)&-\sin(t)&0\\ \sin(t)&\cos(t)&0\\ 0&0&1\end{bmatrix},\quad B(t)=\begin{bmatrix}0\\ 0\\ 0\end{bmatrix}$. (48)
 We have that ${\boldsymbol{\Phi}}^{-1}(t,{\boldsymbol{x}})=[\mathbb{A}(t)]^{-1}({\boldsymbol{x}}-B(t))$ and that the domain velocity is ${\boldsymbol{v}}(t,{\boldsymbol{x}})=\mathbb{A}(t)^{\prime}{\boldsymbol{\Phi}}^{-1}(t,{\boldsymbol{x}})+B(t)^{\prime}\equiv{\boldsymbol{v}}^{\mathcal{A}}$.
 We construct manufactured solution for the problem
-(48) $\partial_{t}u+\nabla_{\Gamma}\cdot({\boldsymbol{b}}_{\Gamma}u)+c_{\Gamma}u=f,$
+$\partial_{t}u+\nabla_{\Gamma}\cdot({\boldsymbol{b}}_{\Gamma}u)+c_{\Gamma}u=f,$ (49)
 with
-(49) $u_{ex}$ $=\cos(2\pi x_{1})\cos(t),$
+$u_{ex}$ $=\cos(2\pi x_{1})\cos(t),$ (50)
 $c_{\Gamma}$ $=1+t^{2},$
-${\boldsymbol{b}}_{\Gamma}$ $={\boldsymbol{v}}^{\mathcal{A}}+{\boldsymbol{b}}_{rel}={\boldsymbol{v}}^{\mathcal{A}}+\mathbb{P}_{\Gamma}(2,-{\boldsymbol{x}},0),$
+${\boldsymbol{b}}_{\Gamma}$ $={\boldsymbol{v}}^{\mathcal{A}}+{\boldsymbol{b}}_{rel}={\boldsymbol{v}}^{\mathcal{A}}+\mathbb{P}_{\Gamma}(2,-x_{1},0),$
 where ${\boldsymbol{x}}=(x_{1},x_{2},x_{3})$.
@@
   1. 1.
-    adrSolver 1: CIP stabilized solver 46.
+    adrSolver 1: CIP stabilized solver (47).
   2. 2.
@@
 Figure 3: Convergence studies for the solver adrSolver 4 in List 5.2. As expected, first and second order convergence are achieved in time and space, respectively.
-To further test the effectiveness of the algorithm we now consider an ill-posed problem. Maintaining the half-sphere geometry and the transformation as in 47, we solve the pure transport problem
-(50) $\partial_{t}u+\nabla_{\Gamma}({\boldsymbol{b}}_{\Gamma}u)=0,$
+To further test the effectiveness of the algorithm we now consider an ill-posed problem. Maintaining the half-sphere geometry and the transformation as in (48), we solve the pure transport problem
+$\partial_{t}u+\nabla_{\Gamma}({\boldsymbol{b}}_{\Gamma}u)=0,$ (51)
 with
-(51) ${\boldsymbol{b}}_{\Gamma}={\boldsymbol{v}}^{\mathcal{A}}+{\boldsymbol{b}}_{rel}={\boldsymbol{v}}^{\mathcal{A}}+(z,0,-x)\cdot(1-e^{-10z}),\quad u_{0}=e^{-3(x^{2}+y^{2})}.$
+${\boldsymbol{b}}_{\Gamma}={\boldsymbol{v}}^{\mathcal{A}}+{\boldsymbol{b}}_{rel}={\boldsymbol{v}}^{\mathcal{A}}+(x_{3},0,-x_{1})\cdot(1-e^{-10x_{3}}),\quad u_{0}=e^{-3(x_{1}^{2}+x_{2}^{2})}.$ (52)
 By construction the initial concentration is pushed towards the boundary where an incredibly sharp boundary layer forms, since zero flux boundary conditions are enforced by the fact that ${\boldsymbol{b}}_{rel}\cdot{\boldsymbol{n}}_{\partial\Gamma}=0$.
@@
 Right: Close-up of minimum value of $u$ for the different solvers in List 5.2.
-It can be seen that the non-stabilized solver is unstable due to the pure advective regime, leading to oscillations that inevitably effect the whole domain 4(a).
+It can be seen that the non-stabilized solver is unstable due to the pure advective regime, leading to oscillations that inevitably effect the whole domain, see Figure 4(a).
 adrSolver 1 effectively controls the solution’s gradient thanks to the stabilization, generating a non-oscillating solution, but fails to maintain its positivity, see Figure 5 (right plot).
@@
 adrSolver 4 successfully manages to deal with the boundary layer, maintain positivity, and avoid oscillations.
-The final solution is shown in 4(b).
+The final solution is shown in Figure 4(b).
 6 Cahn-Hilliard system as a phase separation model
 Phase field modeling plays a central role in biophysical systems to investigate phase transition processes.
-Common phase field models for stationary surfaces are derived from Allen-Cahn and Cahn-Hilliard dynamics [26, 44, 43, 35], where the motion follows the minimization of an energy functional.
+Common phase field models for stationary surfaces are derived from Allen-Cahn and Cahn-Hilliard dynamics Cahn and Hilliard [1958], Erlebacher et al.
+[2001], Elliott and Stinner [2010], Du et al.
+[2011], where the motion follows the minimization of an energy functional.
 We will focus here on the functional
-(52) $\mathcal{E}_{CH}(u)=\int_{M(t)}\sigma\Bigl(\underbrace{\frac{\varepsilon|\nabla_{\Gamma}u|^{2}}{2}}_{I}+\underbrace{\frac{1}{\varepsilon}F(u)}_{II}\Bigr),$
+$\mathcal{E}_{CH}(u)=\int_{M(t)}\sigma\Bigl(\underbrace{\frac{\varepsilon|\nabla_{\Gamma}u|^{2}}{2}}_{I}+\underbrace{\frac{1}{\varepsilon}F(u)}_{II}\Bigr),$ (53)
 where $u$ is the phase field, $\varepsilon\in\mathbb{R}^{+}$ is a length scale parameter determining the interface length and $\sigma\in\mathbb{R}^{+}$ describes the interface tension between different phases.
@@
 We consider here two different double-well potentials
-(53a) $F_{1}(u)$ $:=\frac{1}{4}((1-u)\log(1-u)+(1+u)\log(1+u))+\frac{1-u^{2}}{2},$
-(53b) $F_{2}(u)$ $:=\frac{(1-u^{2})^{2}}{4},$
+$F_{1}(u)$ $:=\frac{1}{4}((1-u)\log(1-u)+(1+u)\log(1+u))+\frac{1-u^{2}}{2},$ (54a)
+$F_{2}(u)$ $:=\frac{(1-u^{2})^{2}}{4},$ (54b)
 for which
-(54) $F_{1}^{\prime}(u)=\log\Bigl(\frac{1+u}{1-u}\Bigr)-u,\text{ and }F_{2}^{\prime}(u)=u^{3}-u\text{ respectively}.$
-$F_{1}$ is a Flory-Huggins-type mixing free energy density whose formulation finds its roots in configurational entropy [59, 48].
+$F_{1}^{\prime}(u)=\frac{1}{4}\log\Bigl(\frac{1+u}{1-u}\Bigr)-u,\text{ and }F_{2}^{\prime}(u)=u^{3}-u\text{ respectively}.$ (55)
+$F_{1}$ is a Flory-Huggins-type mixing free energy density whose formulation finds its roots in configurational entropy Huggins [1941], Flory [1942].
 It can be seen that $F_{1}$ is only well posed for $u\in(-1,1)$.
@@
 The advantage of the latter is that it is well defined for values of $u$ outside the interval (1,1)(-1,1).
-The extension of the Cahn-Hilliard equation to evolving surfaces has seen recent developments due to its applicability to biological membranes and alloys [105, 41].
-The evolving surface Cahn-Hillard equation [24, 25] can be written as
-(55) $\begin{cases}\partial^{\bullet}u+(\nabla_{\Gamma}\cdot{\boldsymbol{v}})u=\nabla_{\Gamma}(m\nabla_{\Gamma}w),\\ w=-\sigma\varepsilon\Delta_{\Gamma}u+\frac{\sigma}{\varepsilon}F^{\prime}(u),\end{cases}$
+The extension of the Cahn-Hilliard equation to evolving surfaces has seen recent developments due to its applicability to biological membranes and alloys Zhiliakov et al.
+[2021], Eilks and Elliott [2008].
+The evolving surface Cahn-Hillard equation Caetano and Elliott [2021], Caetano et al.
+[2023] can be written as
+$\begin{cases}\partial^{\bullet}u+(\nabla_{\Gamma}\cdot{\boldsymbol{v}})u=\nabla_{\Gamma}(m\nabla_{\Gamma}w),\\ w=-\sigma\varepsilon\Delta_{\Gamma}u+\frac{\sigma}{\varepsilon}F^{\prime}(u),\end{cases}$ (56)
 subject to the initial condition $u(0)=u_{0}$ for suitable initial data.
 $w$ represents the chemical potential and $m\in\mathbb{R}^{+}$ is the mobility.
-As explained in [77], models for evolving surfaces have similar energy functions to models on stationary surfaces, but important differences arise.
+As explained in Liu et al.
+[2024], models for evolving surfaces have similar energy functions to models on stationary surfaces, but important differences arise.
 Even if the evolution of the systems is driven by the minimization of (53), energy dissipation is no longer satisfied since the velocity is viewed as external force that adds to the system.
@@
 To keep mass conservation, we must assume $\nabla_{\Gamma}\cdot{\boldsymbol{v}}=0$, which is equivalent to assume inextensibility of the cell membrane, and by consequence area preservation.
-This assumption also allows to consistently extend this one-phase model to $N$-phase models satisfying the hyperlink condition, as described in [77].
+This assumption also allows to consistently extend this one-phase model to $N$-phase models satisfying the hyperlink condition, as described in Liu et al.
+[2024].
 6.1 Numerical methods for Cahn-Hilliard system
-For the discretization of (55), the evolutionary surface finite element method (ESFEM) is one of the most popular methods adopted and the one we consider here [39].
+For the discretization of (56), the evolutionary surface finite element method (ESFEM) is one of the most popular methods adopted and the one we consider here Dziuk and Elliott [2013].
 Its advantage lies in simple implementation, low memory consumption, ability to handle complex deformations and potential for efficient parallelization.
-We refer to [77] for a review of alternative methods.
+We refer to Liu et al.
+[2024] for a review of alternative methods.
 We employ a semi-implicit approach where the fully discrete formulation reads: given $u_{h}^{n}\in V_{h}(\Gamma_{h}^{n})$ find $u^{n+1}_{h},w^{n+1}_{h}\in V_{h}(\Gamma_{h}^{n})$ such that
-(56a) $n_{h}\mathopen{}\left(\frac{u^{n+1}_{h}-u_{h}^{n}}{\tau}\,,\,\phi_{h}\mathclose{}\right)+(({\boldsymbol{v}}_{h}-{\boldsymbol{v}}^{\mathcal{A}}_{h})u^{n+1}_{h}\,,\,\nabla_{\Gamma^{n}_{h}}\phi_{h})_{\Gamma^{n}_{h}}+mb_{h}(w^{n+1}_{h}\,,\,\phi_{h})=0,$
-(56b) $n_{h}(w^{n+1}_{h}\,,\,\psi_{h})-\sigma\varepsilon b_{h}(u^{n+1}_{h}\,,\,\psi_{h})=\frac{\sigma}{\varepsilon}n_{h}(F^{\prime}(u^{n}_{h})\,,\,\psi_{h}),$
+$n_{h}\mathopen{}\left(\frac{u^{n+1}_{h}-u_{h}^{n}}{\tau}\,,\,\phi_{h}\mathclose{}\right)+(({\boldsymbol{v}}_{h}-{\boldsymbol{v}}^{\mathcal{A}}_{h})u^{n+1}_{h}\,,\,\nabla_{\Gamma^{n}_{h}}\phi_{h})_{\Gamma^{n}_{h}}+mb_{h}(w^{n+1}_{h}\,,\,\phi_{h})=0,$ (57a)
+$n_{h}(w^{n+1}_{h}\,,\,\psi_{h})-\sigma\varepsilon b_{h}(u^{n+1}_{h}\,,\,\psi_{h})=\frac{\sigma}{\varepsilon}n_{h}(F^{\prime}(u^{n}_{h})\,,\,\psi_{h}),$ (57b)
 for all $\phi_{h},\psi_{h}\in V_{h}(\Gamma_{h}^{n})$ with
-(57) $n_{h}(u_{h}\,,\,\phi_{h})=(u_{h}\,,\,\phi_{h})_{\Gamma_{h}^{n}},\quad b_{h}(u_{h}\,,\,\psi_{h})=(\nabla_{\Gamma_{h}^{n}}u_{h}\,,\,\nabla_{\Gamma_{h}^{n}}\psi_{h})_{\Gamma_{h}^{n}}.$
-Similar discretizations can be found in [43, 7, 85], where a Cahn-Hilliard solver is coupled with a domain evolution which is unknown.
+$n_{h}(u_{h}\,,\,\phi_{h})=(u_{h}\,,\,\phi_{h})_{\Gamma_{h}^{n}},\quad b_{h}(u_{h}\,,\,\psi_{h})=(\nabla_{\Gamma_{h}^{n}}u_{h}\,,\,\nabla_{\Gamma_{h}^{n}}\psi_{h})_{\Gamma_{h}^{n}}.$ (58)
+Similar discretizations can be found in Elliott and Stinner [2010], Bachini et al.
+[2023], Mokbel et al.
+[2024], where a Cahn-Hilliard solver is coupled with a domain evolution which is unknown.
 In those works, the potential $F_{2}$ was used.
@@
 The scheme proceeds as follows.
-In the predictor step we solve the unconstrained problem 56. Taking the non-linear term explicitly ensures $u_{h}^{n}$ lays inside the desired bound.
+In the predictor step we solve the unconstrained problem (57). Taking the non-linear term explicitly ensures $u_{h}^{n}$ lays inside the desired bound.
 In the corrector step we solve the constrained problem imposing $u^{n+1}_{h}\in(-1,1)$.
-Remark 6.1.
-Another strength of the postprocessing technique of [29] evident here is that it can act directly on the quantity of interest per se without involving $w^{n+1}_{h}$, which is unbounded.
-We remind the reader that extensions for higher order BDFs methods are presented in the original articles.
 6.2 Numerical results for Cahn-Hilliard system
 We proceed in reporting experimental studies to assess the accuracy of the proposed framework.
-Consider a cylinder revolving around the $z$-axis with radius $R=1$ and height $l=2$.
+Consider a cylinder revolving around the $x_{3}$-axis with radius $R=1$ and height $l=2$.
 To satisfy the assumption $\nabla_{\Gamma}\cdot{\boldsymbol{v}}=0$, we transform the initial cylinder through the isometry ${\boldsymbol{\Phi}}(t,{\boldsymbol{p}})={\boldsymbol{\Phi}}^{\mathcal{A}}(t,{\boldsymbol{p}})=\mathbb{A}(t){\boldsymbol{p}}+B(t)$
-(58) $\mathbb{A}(t)=\begin{bmatrix}\lambda(t)&0&0\\ 0&\lambda(t)&0\\ 0&0&1/\lambda(t)\end{bmatrix},\quad B(t)=\begin{bmatrix}0.2t\\ 0.1t\\ 0\end{bmatrix}$,
+$\mathbb{A}(t)=\begin{bmatrix}\lambda(t)&0&0\\ 0&\lambda(t)&0\\ 0&0&1/\lambda(t)\end{bmatrix},\quad B(t)=\begin{bmatrix}0.2t\\ 0.1t\\ 0\end{bmatrix}$, (59)
 where $\lambda(t)=1+0.5\sin(\pi t)$.
@@
   1. 1.
-    chSolver 1: Solver 56 with no postprocessing.
+    chSolver 1: Solver (57) with no postprocessing.
   2. 2.
-    chSolver 2: Solver 56 with imposed bound preservation in the interval (1,1)(-1,1).
+    chSolver 2: Solver (57) with imposed bound preservation in the interval (1,1)(-1,1).
   3. 3.
-    chSolver 3: Solver 56 with imposed mass preservation using the lumped-mass inner product $(\cdot\,,\,\cdot)^{h}$.
+    chSolver 3: Solver (57) with imposed mass preservation using the lumped-mass inner product $(\cdot\,,\,\cdot)^{h}$.
   4. 4.
-    chSolver 4: Solver 56 with imposed bound preservation in the interval (1,1)(-1,1) and mass preservation.
-Zero flux Neumann boundary conditions are imposed for both $u$ and $w$. A random uniform initial distribution in the interval [1,1][-1,1] is chosen for $u_{0}$. The parameters are set as
+    chSolver 4: Solver (57) with imposed bound preservation in the interval (1,1)(-1,1) and mass preservation.
+Zero flux Neumann boundary conditions are imposed for both $u$ and $w$. A random uniform initial distribution in the interval (1,1)(-1,1) is chosen for $u_{0}$. The parameters are set as
 $h$ $m$ $\varepsilon$ $\sigma$ $T$ $\tau$
@@
 7 Membrane force system
-In the realm of mathematical modeling, principal curvatures have long been used for the theory of elastic plates and shells, with the work of Kirchhoff [66] representing what may be the most famous case.
-Later in the ’70s, the work of Canham [27] and Helfrich [54] on the link between biological membranes shape and curvature functionals has de facto initiated an entirely new field of research focused on the use of geometry of hypersurfaces for biophysical modeling.
+In the realm of mathematical modeling, principal curvatures have long been used for the theory of elastic plates and shells, with the work of Kirchhoff Kirchhoff [1850] representing what may be the most famous case.
+Later in the ’70s, the work of Canham Canham [1970] and Helfrich Helfrich [1973] on the link between biological membranes shape and curvature functionals has de facto initiated an entirely new field of research focused on the use of geometry of hypersurfaces for biophysical modeling.
 We will here focus on the Canham-Helfrich functional $\mathcal{E}_{B}(t)$ and its discretization. Considering membranes with boundaries, a commonly used expression for the energy functional is
-(59) $\mathcal{E}_{B}(t)=\int_{\Gamma}\frac{\gamma_{W}}{2}(\kappa-\kappa_{0})^{2}+\int_{\Gamma}\gamma_{G}\kappa_{G}+\gamma_{\partial}\int_{\partial\Gamma}1,$
+$\mathcal{E}_{B}(t)=\int_{\Gamma}\frac{\gamma_{W}}{2}(\kappa-\kappa_{0})^{2}+\int_{\Gamma}\gamma_{G}\kappa_{G}+\gamma_{\partial}\int_{\partial\Gamma}1,$ (60)
 where $\gamma_{W},\;\gamma_{G}\in\mathbb{R}$ are bending rigidities, $\kappa:\mathcal{G}_{T}\to\mathbb{R}$ is the previously introduced mean curvature, $\kappa_{0}\in\mathbb{R}$ is the spontaneous curvature, and $\kappa_{G}:\mathcal{G}_{T}\to\mathbb{R}$ is the Gaussian curvature $\kappa_{G}=\det\mathbb{W}$.
 The parameter $\gamma_{\partial}$ takes into account the possible line energy of $\partial\Gamma$.
-For the sake of simplicity we will consider the case $\gamma_{G}=constant$, $\gamma_{W}=constant$, and assume that no topological change occur along the evolution.
-Taken into account these simplifications, the gradient flow of a membrane $\Gamma$ is a family of evolving surfaces with normal velocity $\boldsymbol{v}$$^{\perp}$ given by [88, 14]
-(60) ${\boldsymbol{v}}^{\perp}=\gamma_{W}\Bigl(-\Delta_{\Gamma}\kappa+\frac{1}{2}\kappa(\kappa-\kappa_{0})^{2}-(\kappa-\kappa_{0})|\mathbb{W}|^{2}\Bigr){\boldsymbol{n}}_{\Gamma}.$
+For the sake of simplicity we will consider the case $\gamma_{G}=constant$, $\gamma_{W}=constant$, and assume that no topological change occurs along the evolution.
+Taken into account these simplifications, the gradient flow of a membrane $\Gamma$ is a family of evolving surfaces with normal velocity $\boldsymbol{v}$$^{\perp}$ given by Nitsche [1993], Barrett et al.
+[2017a]
+${\boldsymbol{v}}^{\perp}=\gamma_{W}\Bigl(-\Delta_{\Gamma}\kappa+\frac{1}{2}\kappa(\kappa-\kappa_{0})^{2}-(\kappa-\kappa_{0})|\mathbb{W}|^{2}\Bigr){\boldsymbol{n}}_{\Gamma}.$ (61)
 Note that such velocity is in the normal direction since tangential changes to the surface shape do not modify its energy.
@@
 Equation (61) thus leads to a fourth-order nonlinear equation posed on a moving manifold.
-Remark 7.1.
+Remark 5.
 Other geometrical flows like the mean curvature flow and the surface diffusion flow are also of interest in this field.
 Being somewhat simpler, extensive research is already available for them, while less is known about the Willmore flow, which we focus on in this article.
-We refer to [38, 32, 15] for further details.
-Remark 7.2.
+We refer to Dziuk and Elliott [2007], Deckelnick et al.
+[2005], Barrett et al.
+[2020] for further details.
+Remark 6.
 Since this is a defining equation for the normal velocity, additional forces can be taken into account by introducing a right-hand side to (61).
 7.1 Numerical methods for Helfrich system
-We choose here the family of BGN-type algorithms as pioneered in [11] and later adapted for Willmore flow in [12, 14]. The schemes pivot around the two following ingredients:
-  • Introduce the mean curvature vector ${\boldsymbol{\kappa}}(t)=\kappa(t){\boldsymbol{n}}_{\Gamma}(t)$ as independent variable and use the defining equation for the mean curvature vector ${\boldsymbol{\kappa}}=\Delta_{\Gamma}\mathbf{Id}_{\Gamma}$ in order to reduce the fourth order problem to a coupled second order one [11];
+We choose here the family of BGN-type algorithms as pioneered in Barrett et al.
+[2007] and later adapted for Willmore flow in Barrett et al.
+[2016, 2017a].
+The schemes pivot around the two following ingredients:
+  • Introduce the mean curvature vector ${\boldsymbol{\kappa}}(t)=\kappa(t){\boldsymbol{n}}_{\Gamma}(t)$ as independent variable and use the defining equation for the mean curvature vector ${\boldsymbol{\kappa}}=\Delta_{\Gamma}\mathbf{Id}_{\Gamma}$ in order to reduce the fourth order problem to a coupled second order one Barrett et al.
+[2007];
   • Use the weak formulation to rewrite and simplify the term $\kappa|\mathbb{W}|^{2}$ (more precisely $\kappa|\mathbb{H}|^{2}$ in the immersed setting) taking advantage of the newly introduced variable $\boldsymbol{\kappa}$}.
-This reformulation not only simplifies and linearizes the system but also leads to stable schemes as proven in [37].
-Our scheme choice builds on the one derived in [14] for boundary value problems.
-Among the different types of boundary conditions considered in [14] we will restrict to the case of clamped boundary conditions, for which
+This reformulation not only simplifies and linearizes the system but also leads to stable schemes as proven in Dziuk [2008].
+Our scheme choice builds on the one derived in Barrett et al.
+[2017a] for boundary value problems.
+Among the different types of boundary conditions considered in Barrett et al.
+[2017a] we will restrict to the case of clamped boundary conditions, for which
 […]
@@
 The weak form of (61) then reads: Find ${\boldsymbol{v}}^{\perp},~{\boldsymbol{y}},~{\boldsymbol{\kappa}}\in[H(t)]^{d}$ such that
-(62a) $({\boldsymbol{v}}^{\perp}\,,\,{\boldsymbol{\phi}})=~(\nabla_{\Gamma}{\boldsymbol{y}}\,,\,\nabla_{\Gamma}{\boldsymbol{\phi}})+(\nabla_{\Gamma}\cdot{\boldsymbol{y}}\,,\,\nabla_{\Gamma}\cdot{\boldsymbol{\phi}})-((\nabla_{\Gamma}{\boldsymbol{y}})^{T}\,,\,\mathcal{D}({\boldsymbol{\phi}})\;\mathbb{P}_{\Gamma})$
+$({\boldsymbol{v}}^{\perp}\,,\,{\boldsymbol{\phi}})=~(\nabla_{\Gamma}{\boldsymbol{y}}\,,\,\nabla_{\Gamma}{\boldsymbol{\phi}})+(\nabla_{\Gamma}\cdot{\boldsymbol{y}}\,,\,\nabla_{\Gamma}\cdot{\boldsymbol{\phi}})-((\nabla_{\Gamma}{\boldsymbol{y}})^{T}\,,\,\mathcal{D}({\boldsymbol{\phi}})\;\mathbb{P}_{\Gamma})$ (63a)
 $\qquad-\kappa_{0}({\boldsymbol{\kappa}}\,,\,(\nabla_{\Gamma}{\boldsymbol{\phi}})^{T}{\boldsymbol{n}}_{\Gamma})-\frac{1}{2}(\gamma_{W}|{\boldsymbol{\kappa}}-\kappa_{0}{\boldsymbol{n}}_{\Gamma}|^{2}\;\mathbb{P}\,,\,\nabla_{\Gamma}{\boldsymbol{\phi}})$
 $\qquad+(({\boldsymbol{y}}\cdot{\boldsymbol{\kappa}})\;\mathbb{P}_{\Gamma}\,,\,\nabla_{\Gamma}{\boldsymbol{\phi}}),$
-(62b) $({\boldsymbol{y}}\,,\,{\boldsymbol{\psi}})=~(\gamma_{W}({\boldsymbol{\kappa}}-\kappa_{0}{\boldsymbol{n}}_{\Gamma})\,,\,{\boldsymbol{\psi}}),$
-(62c) $({\boldsymbol{\kappa}}\,,\,{\boldsymbol{\xi}})=~-(\mathbb{P}_{\Gamma}\,,\,\nabla_{\Gamma}{\boldsymbol{\xi}})+({\boldsymbol{\mu}}\,,\,{\boldsymbol{\xi}})_{\partial\Gamma},$
+$({\boldsymbol{y}}\,,\,{\boldsymbol{\psi}})=~(\gamma_{W}({\boldsymbol{\kappa}}-\kappa_{0}{\boldsymbol{n}}_{\Gamma})\,,\,{\boldsymbol{\psi}}),$ (63b)
+$({\boldsymbol{\kappa}}\,,\,{\boldsymbol{\xi}})=~-(\mathbb{P}_{\Gamma}\,,\,\nabla_{\Gamma}{\boldsymbol{\xi}})+({\boldsymbol{\mu}}\,,\,{\boldsymbol{\xi}})_{\partial\Gamma},$ (63c)
 for all ${\boldsymbol{\phi}},{\boldsymbol{\psi}},{\boldsymbol{\xi}}\in[H(t)]^{d}$, where $H(t)$ is the Sobolev space $W^{1,2}(\Gamma(t))$.
@@
 For the discretization of the normal velocity in (63) we choose to introduce the displacement variable ${\boldsymbol{d}}_{h}^{n+1}\in[V_{h}(\Gamma_{h}^{n})]^{d}$ defined as
-(63) ${\boldsymbol{v}}_{h}^{\perp}(t_{n})=\frac{1}{\tau}\sum_{i}({\boldsymbol{y}}_{i}(t_{n+1})-{\boldsymbol{y}}_{i}(t_{n})){\boldsymbol{\phi}}_{i}(t_{n})=\frac{1}{\tau}{\boldsymbol{d}}_{h}^{n+1},$
-where we recall that ${\boldsymbol{y}}_{i}$ are the mesh points and $\{{\boldsymbol{\phi}}_{i}\}$ the associated vector-valued basis such that ${\boldsymbol{\phi}}_{i}({\boldsymbol{y}}_{j})=\delta_{ij}$ at all times.
+${\boldsymbol{v}}_{h}^{\perp}(t_{n})=\frac{1}{\tau}\sum_{i}({\boldsymbol{y}}_{i}(t_{n+1})-{\boldsymbol{y}}_{i}(t_{n}))\phi_{i}(t_{n})=\frac{1}{\tau}{\boldsymbol{d}}_{h}^{n+1},$ (64)
+where we recall that ${\boldsymbol{y}}_{i}$ are the mesh points and $\{\phi_{i}\}$ the associated basis such that $\phi_{i}({\boldsymbol{y}}_{j})=\delta_{ij}$ at all times.
 Moreover, (63c) and (63b) can be collapsed to
 $\frac{1}{\gamma_{W}}({\boldsymbol{y}}^{n+1}\,,\,{\boldsymbol{\psi}})_{\Gamma^{n}}+(\nabla_{\Gamma^{n}}{\boldsymbol{d}}^{n+1}\,,\,\nabla_{\Gamma^{n}}{\boldsymbol{\psi}})_{\Gamma^{n}}=-(\mathbb{P}_{\Gamma^{n}}\,,\,\nabla_{\Gamma^{n}}{\boldsymbol{\psi}})_{\Gamma^{n}}$
-(64) $+(\kappa_{0}{\boldsymbol{n}}_{\Gamma^{n}}\,,\,{\boldsymbol{\psi}})_{\Gamma^{n}}+({\boldsymbol{\mu}}\,,\,{\boldsymbol{\psi}})_{\partial\Gamma^{n}}.$
+$+(\kappa_{0}{\boldsymbol{n}}_{\Gamma^{n}}\,,\,{\boldsymbol{\psi}})_{\Gamma^{n}}+({\boldsymbol{\mu}}\,,\,{\boldsymbol{\psi}})_{\partial\Gamma^{n}}.$ (65)
 The fully discrete form of (63) reads: Given ${\boldsymbol{y}}^{n}_{h},{\boldsymbol{\kappa}}^{n}_{h}\in[V_{h}(\Gamma_{h}^{n})]^{d}$ find ${\boldsymbol{d}}^{n+1}_{h}\in[V_{h0}(\Gamma_{h}^{n})]^{d}$ and ${\boldsymbol{y}}^{n+1}_{h}\in[V_{h}(\Gamma_{h}^{n})]^{d}$ such that
-(65a) $\mathopen{}\left(\frac{{\boldsymbol{d}}_{h}^{n+1}}{\tau}\,,\,{\boldsymbol{\phi}}_{h}\mathclose{}\right)_{\Gamma_{h}^{n}}^{h}-(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{y}}^{n+1}_{h}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}=~(\nabla_{\Gamma_{h}^{n}}\cdot{\boldsymbol{y}}^{n}_{h}\,,\,\nabla_{\Gamma_{h}^{n}}\cdot{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}$
+$\mathopen{}\left(\frac{{\boldsymbol{d}}_{h}^{n+1}}{\tau}\,,\,{\boldsymbol{\phi}}_{h}\mathclose{}\right)_{\Gamma_{h}^{n}}^{h}-(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{y}}^{n+1}_{h}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}=~(\nabla_{\Gamma_{h}^{n}}\cdot{\boldsymbol{y}}^{n}_{h}\,,\,\nabla_{\Gamma_{h}^{n}}\cdot{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}$ (66a)
 $\qquad-((\nabla_{\Gamma_{h}^{n}}{\boldsymbol{y}}^{n}_{h})^{T}\,,\,\mathcal{D}({\boldsymbol{\phi}}_{h})\;\mathbb{P}_{\Gamma_{h}^{n}})_{\Gamma_{h}^{n}}-\kappa_{0}({\boldsymbol{\kappa}}^{n}_{h}\,,\,(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})^{T}{\boldsymbol{n}}_{\Gamma_{h}^{n}})_{\Gamma_{h}^{n}}^{h}$
 $\qquad-\frac{1}{2}(\gamma_{W}|{\boldsymbol{\kappa}}^{n}_{h}-\kappa_{0}{\boldsymbol{n}}_{\Gamma_{h}^{n}}|^{2}\;\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}^{h}+(({\boldsymbol{y}}^{n}_{h}\cdot{\boldsymbol{\kappa}}^{n}_{h})\;\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}^{h},$
-(65b) $\frac{1}{\gamma_{W}}({\boldsymbol{y}}^{n+1}_{h}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{d}}_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}$
+$\frac{1}{\gamma_{W}}({\boldsymbol{y}}^{n+1}_{h}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{d}}_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}$ (66b)
 $\qquad=-(\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}+(\kappa_{0}{\boldsymbol{n}}_{\Gamma_{h}^{n}}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+({\boldsymbol{\mu}}\,,\,{\boldsymbol{\psi}}_{h})_{\partial\Gamma^{n}},$
 for all ${\boldsymbol{\phi}}_{h}\in[V_{h0}(\Gamma_{h}^{n})]^{d},~{\boldsymbol{\psi}}_{h}\in[V_{h}(\Gamma_{h}^{n})]^{d}$, where $(\cdot\,,\,\cdot)^{h}$ is the mass-lumped inner product.
-Remark 7.3.
+Remark 7.
 The mean curvature ${\boldsymbol{\kappa}}_{h}^{n+1}$ can be recovered a posteriori as ${\boldsymbol{\kappa}}_{h}^{n+1}=\pi^{h}(\gamma_{W}^{-1}{\boldsymbol{y}}_{h}^{n+1}+\kappa_{0}{\boldsymbol{n}}_{\Gamma_{h}^{n}})$ where $\pi^{h}(\cdot)$ is the standard interpolation operator.
-Remark 7.4.
-Note that (65) corresponds to the scheme presented in [12] for the choice of the parameter $\theta=1$, i.e. without tangential motion control.
+Remark 8.
+Note that (66) corresponds to the scheme presented in Barrett et al. [2016] for the choice of the parameter $\theta=1$, i.e. without tangential motion control.
 7.2 Numerical results for Helfrich system
-To test the accuracy of our method for gradient flows, we start by performing classical studies on Willmore flow such as the ones in [37, 12, 14]. For every test we compare
+To test the accuracy of our method for gradient flows, we start by performing classical studies on Willmore flow such as the ones in Dziuk [2008], Barrett et al.
+[2016, 2017a].
+For every test we compare
   1. 1.
@@
 We highlight that it is of crucial importance that, when the algorithm in Section 4 is employed, geometrical quantities characteristic of the mesh are updated.
-In our case, we refer to the mean curvature ${\boldsymbol{\kappa}}_{h}$ and the auxiliary variable ${\boldsymbol{y}}_{h}$ of the algorithm in (65).
-We start by testing the convergence using the example presented in [12, Section 5.3].
+In our case, we refer to the mean curvature ${\boldsymbol{\kappa}}^{n}_{h}$ and the auxiliary variable ${\boldsymbol{y}}^{n}_{h}$ of the algorithm in (66).
+We start by testing the convergence using the example presented in [Barrett et al., 2016, Section 5.3].
 It is an evolving sphere with initial radius $R_{0}$ and spontaneous curvature $\kappa_{0}$.
 The radius evolution $R(t)$ is the solution to the ordinary differential equation
-(66) $R_{t}=-\frac{\kappa_{0}}{R}\Bigl(\frac{2}{R}+\kappa_{0}\Bigr),\quad R(0)=R_{0}>0.$
+$\partial_{t}R=-\frac{\kappa_{0}}{R}\Bigl(\frac{2}{R}+\kappa_{0}\Bigr),\quad R(0)=R_{0}>0.$ (67)
 The error norm chosen to verify convergence is the following
-(67) $\lVert\Gamma-\Gamma_{h}\rVert_{L^{\infty}(L^{\infty})}=\max_{n}\max_{i}|\lVert{\boldsymbol{y}}_{i}(t_{n})\rVert-R(t_{n})|,$
+$\lVert\Gamma-\Gamma_{h}\rVert_{L^{\infty}(L^{\infty})}=\max_{n}\max_{i}|\lVert{\boldsymbol{y}}_{i}(t_{n})\rVert-R(t_{n})|,$ (68)
 where ${\boldsymbol{y}}_{i}$ are the vertices of the mesh simplices.
 Figure 8: Convergence studies for Willmore flow of sphere under spontaneous mean curvature for the solvers in List 7.2.
+As expected, first and second order convergence are achieved in time and space, respectively.
 Figure 9: Energy evolution for torus with $R=2$, $r=1$ for the solvers in List 7.2.
@@
 Moreover, we can see that the use of the mesh redistribution actually improves the convergence rates.
-We believe this behavior is due to the automatic mesh adaptivity of Equation 33 (already mentioned in the original article [36]), which tends to accumulate nodes in certain regions providing better geometry description.
-As a second test we check the energy evolution for a torus of major radius $R=2$ and minor radius $r=1$ as done in [12, Section 5.1].
+We believe this behavior is due to the automatic mesh adaptivity of (33) (already mentioned in the original article Duan and Li [2024]), which tends to accumulate nodes in certain regions providing better geometry description.
+As a second test we check the energy evolution for a torus of major radius $R=2$ and minor radius $r=1$ as done in [Barrett et al., 2016, Section 5.1].
 The minimizer of the Willmore energy in this case is the Clifford torus with radius $R=\sqrt{2},\;r=1$, which has energy $\mathcal{E}=4\pi^{2}\approx 39.4784$.
@@
 We notice how both simulations using hfSolver 1 break before reaching the end time.
-This is due to instabilities that occur at regions of high curvature already observed in [12, Figure 2] for $\theta=1$.
+This is due to instabilities that occur at regions of high curvature already observed in [Barrett et al., 2016, Figure 2] for $\theta=1$.
 Applying the redistribution doesn’t impact the energy landscape and allows to proceed with the simulation until the stable state is reached.
-Moreover, refining the mesh shows how the energy converges toward the expected value.
+Moreover, refining the mesh shows how the energy converges towards the expected value.
 7.2.1 Pinching tests for Helfrich system
-We perform consolidated tests for the Willmore flow under the influence of spontaneous mean curvature as presented in [12, Section 5.3] and then repeated in [51], where a similar two-step procedure as the one adopted here is used.
+We perform consolidated tests for the Willmore flow under the influence of spontaneous mean curvature as presented in [Barrett et al., 2016, Section 5.3] and then repeated in Garcke et al.
+[2025], where a similar two-step procedure as the one adopted here is used.
 We consider two different tests for cigar-shaped surfaces:
   1. 1.
-    hf-Test 1 with initial mesh as in 10(a) and the following sets of parameters
+    hf-Test 1 with initial mesh as in Figure 10(a) and the following sets of parameters
     $h$ nvnv nene $T$ $\kappa_{0}$
@@
   2. 2.
-    hf-Test 2 with initial mesh as in 10(d) and the following sets of parameters
+    hf-Test 2 with initial mesh as in Figure 10(d) and the following sets of parameters
     $h$ nvnv nene $T$ $\kappa_{0}$
@@
 The results are plotted in Figure 10 for both hfSolver 1 and 2.
-It can be seen how the mesh adaptivity mentioned in [36] leads to a better resolution of the neck and overall better results for hfSolver 2.
-By comparison of 10(c) with [12, Figure 12] the shape is qualitatively in accordance with what is expected.
-Even more, while [12, Figure 12] requires a finer mesh to accurately resolve the three pearls, the mesh adaptivity of [36] alleviates that requirement for similar resolution.
-Figure 10 is also in accordance with [51, Figure 5.14], where a similar algorithm is employed for a cigar-shaped surface with different aspect ratio.
+It can be seen how the mesh adaptivity mentioned in Duan and Li [2024] leads to a better resolution of the neck and overall better results for hfSolver 2.
+By comparison of Figure 10(c) with [Barrett et al., 2016, Figure 12] the shape is qualitatively in accordance with what is expected.
+Even more, while [Barrett et al., 2016, Figure 12] requires a finer mesh to accurately resolve the three pearls, the mesh adaptivity of Duan and Li [2024] alleviates that requirement for similar resolution.
+Figure 10 is also in accordance with [Garcke et al., 2025, Figure 5.14], where a similar algorithm is employed for a cigar-shaped surface with different aspect ratio.
 As it happens in that article, it has to be noted that the refinement of the necks comes to the expense of a coarser mesh at the lobes.
@@
 We have then the option to choose between explicit and implicit staggering.
-Both approaches have shown to be effective in the literature [84, 65, 64, 51].
-In [7, Appendix E], where fluid deformable membranes with phase separation are considered, both procedures show very similar results even in the most complex scenarios.
-Moreover, when implicit staggering is chosen, the authors of [7] claim that convergence is achieved in four steps on average.
+Both approaches have shown to be effective in the literature Mercker et al.
+[2013], Khanwale et al.
+[2023, 2020], Garcke et al.
+[2025].
+In [Bachini et al., 2023, Appendix E], where fluid deformable membranes with phase separation are considered, both procedures show very similar results even in the most complex scenarios.
+Moreover, when implicit staggering is chosen, the authors of Bachini et al.
+[2023] claim that convergence is achieved in four steps on average.
 Strong of these results we proceed in coupling our solvers.
-Remark 8.1.
+Remark 9.
 An alternative option is to use a monolithic approach and couple the solvers in a unique system, avoiding staggering.
@@
 Together with this, eventual non-linearities have to be linearized.
-One can find an example of such a choice in [85], where wetting dynamics of liquid droplets on deformable membranes is simulated.
+One can find an example of such a choice in Mokbel et al.
+[2024], where wetting dynamics of liquid droplets on deformable membranes is simulated.
 8.1 Numerical results for coupled mean curvature and ADR equations
-We begin by testing the convergence properties of our coupled solver. To do so we pick the coupling example presented in [67]. The system which is solved is the following [67, p.686]:
-(68a) $\partial^{\bullet}u+u\nabla_{\Gamma}\cdot{\boldsymbol{v}}-\Delta_{\Gamma}u$ $=f(t,{\boldsymbol{x}}),$
-(68b) ${\boldsymbol{v}}-\alpha\Delta_{\Gamma}{\boldsymbol{v}}-\beta\Delta_{\Gamma}\mathbf{Id}_{\Gamma}$ $=(\delta u+g(t,{\boldsymbol{x}})){\boldsymbol{n}}_{\Gamma},$
+We begin by testing the convergence properties of our coupled solver.
+To do so we pick the coupling example presented in Barreira et al.
+[2011], Kovács et al.
+[2017].
+The system which is solved is the following [Kovács et al., 2017, p.686]:
+$\partial^{\bullet}u+u\nabla_{\Gamma}\cdot{\boldsymbol{v}}-\Delta_{\Gamma}u$ $=f(t,{\boldsymbol{x}}),$ (69a)
+${\boldsymbol{v}}-\alpha\Delta_{\Gamma}{\boldsymbol{v}}-\beta\Delta_{\Gamma}\mathbf{Id}_{\Gamma}$ $=(\delta u+g(t,{\boldsymbol{x}})){\boldsymbol{n}}_{\Gamma},$ (69b)
 with parameters $\alpha,\beta,\delta\in\mathbb{R}^{+}$ and ${\boldsymbol{x}}=(x_{1},x_{2},x_{3})$.
-Convergence studies are perfomed choosing $f,g$ such that the exact solution for $u$ is $u(t,{\boldsymbol{x}})=x_{1}x_{2}e^{-6t}$.
+Convergence studies are performed choosing $f,g$ such that the exact solution for $u$ is $u(t,{\boldsymbol{x}})=x_{1}x_{2}e^{-6t}$.
 The geometry is chosen to be a sphere whose radius evolves following the law
-(69) $R(t)=\frac{r_{0}r_{K}}{r_{K}e^{-kt}+r_{0}(1-e^{-kt})},$
+$R(t)=\frac{r_{0}r_{K}}{r_{K}e^{-kt}+r_{0}(1-e^{-kt})},$ (70)
 with parameters $r_{0},r_{K},k\in\mathbb{R}^{+}$. The parameters are chosen as follows
@@
 1 0 1 0.4 1 2 0.5
-To simulate (68a) we can use what has been introduced in Section 5 and add a right-had side coupling.
+To simulate (69a) we can use what has been introduced in Section 5 and add a right-hand side coupling.
 Since we chose $\alpha=0$, we can simulate (69b) by simplifying (66).
 The result is a weighted mean curvature solver that, in this very case, takes the form
-(70a) $\mathopen{}\left(\frac{{\boldsymbol{d}}_{h}^{n+1}}{\tau}\,,\,{\boldsymbol{\phi}}_{h}\mathclose{}\right)_{\Gamma_{h}^{n}}^{h}-\beta({\boldsymbol{\kappa}}^{n+1}_{h}\,,\,{\boldsymbol{\phi}}_{h})^{h}_{\Gamma_{h}^{n}}$ $=((\delta u+g(t,{\boldsymbol{x}})){\boldsymbol{n}}_{\Gamma_{h}^{n}}\,,\,{\boldsymbol{\phi}}_{h})^{h}_{\Gamma_{h}^{n}},$
-(70b) $({\boldsymbol{\kappa}}^{n+1}_{h}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{d}}_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}$ $=-(\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}.$
+$\mathopen{}\left(\frac{{\boldsymbol{d}}_{h}^{n+1}}{\tau}\,,\,{\boldsymbol{\phi}}_{h}\mathclose{}\right)_{\Gamma_{h}^{n}}^{h}-\beta({\boldsymbol{\kappa}}^{n+1}_{h}\,,\,{\boldsymbol{\phi}}_{h})^{h}_{\Gamma_{h}^{n}}$ $=((\delta u+g(t,{\boldsymbol{x}})){\boldsymbol{n}}_{\Gamma_{h}^{n}}\,,\,{\boldsymbol{\phi}}_{h})^{h}_{\Gamma_{h}^{n}},$ (71a)
+$({\boldsymbol{\kappa}}^{n+1}_{h}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{d}}_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}$ $=-(\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}.$ (71b)
 We perform the test using implicit staggering.
-The results of the convergence studies are shown in Figure 11, where the same norm as in Equation 67 is used to study the results.
+The results of the convergence studies are shown in Figure 11, where the same norm as in (68) is used to study the results.
 One can see how the expected convergence is achieved in both space and time.
-Figure 11: Convergence studies for mcadrSolver 4 and problem (68).
-To qualitatively visualize the effect of the mesh redistribution and ADR stabilization, the example in [67, Sec.
+Figure 11: Convergence studies for mcadrSolver 4 and problem (69). As expected, first and second order convergence are achieved in time and space, respectively.
+To qualitatively visualize the effect of the mesh redistribution and ADR stabilization, the example in [Kovács et al., 2017, Sec.
 11.2] is also reproduced.
@@
 The coupled system is as follows:
-(71a) $\partial^{\bullet}u+u\nabla_{\Gamma}\cdot{\boldsymbol{v}}-\Delta_{\Gamma}u$ $=f_{1}(u,w),$
-(71b) $\partial^{\bullet}w+w\nabla_{\Gamma}\cdot{\boldsymbol{v}}-D_{c}\Delta_{\Gamma}w$ $=f_{2}(u,w),$
-(71c) ${\boldsymbol{v}}-\alpha\Delta_{\Gamma}{\boldsymbol{v}}-\beta\nabla_{\Gamma}\mathbf{Id}_{\Gamma}$ $=\delta u{\boldsymbol{n}}_{\Gamma},$
+$\partial^{\bullet}u+u\nabla_{\Gamma}\cdot{\boldsymbol{v}}-\Delta_{\Gamma}u$ $=f_{1}(u,w),$ (72a)
+$\partial^{\bullet}w+w\nabla_{\Gamma}\cdot{\boldsymbol{v}}-D_{c}\Delta_{\Gamma}w$ $=f_{2}(u,w),$ (72b)
+${\boldsymbol{v}}-\alpha\Delta_{\Gamma}{\boldsymbol{v}}-\beta\nabla_{\Gamma}\mathbf{Id}_{\Gamma}$ $=\delta u{\boldsymbol{n}}_{\Gamma},$ (72c)
 with the non-linear couplings
-(72) $f_{1}(u,w)=\gamma(a-u+u^{2}w),\quad f_{2}(u,w)=\gamma(b-u^{2}w).$
+$f_{1}(u,w)=\gamma(a-u+u^{2}w),\quad f_{2}(u,w)=\gamma(b-u^{2}w).$ (73)
 The parameters are set as
@@
 8 $10^{-3}$ 0.07 0 0.01 0.4 0.1 0.9 100 10
-As discussed in [67], the nonlinear system composed by (71a) and (71b) is solved until $t=5$ without coupling it to the mesh evolution, i.e.
+As discussed in Kovács et al.
+[2017], the nonlinear system composed by (72a) and (72b) is solved until $t=5$ without coupling it to the mesh evolution, i.e.
 ignoring (72c).
 After that, the full system is evolved until $t=8$.
-Implicit staggering is used.
-We thus have the following solvers:
+We study the following solvers:
   1. 1.
@@
     mcadrSolver 2: Implicit staggering of (47) and (71) and mesh redistribution.
-In Figure 12 we compare the mesh quality at the final time $t=8$. It is evident that the redistribution step keeps the mesh well-behaved while maintaining solution accuracy.
+In Figure 12 we compare the mesh quality at the final time $t=8$. It is evident that the redistribution step keeps the mesh well-behaved while maintaining the solution accuracy.
 (a) mcadrSolver 1
 (b) mcadrSolver 2
-Figure 12: Simulation at $t=8$ for the model (71) [67, Sec. 11.2] and the implicitly staggered schemes in List 8.1.
+Figure 12: Simulation at $t=8$ for the model (72) [Kovács et al., 2017, Sec. 11.2] and the implicitly staggered schemes in List 8.1.
+We can construct a bulk-surface problem by modifying the system in (72) with the following one
+$\partial^{\bullet}u+u\nabla\cdot{\boldsymbol{v}}-\Delta u$ $=f_{1}(u,w)\quad\text{ in }\Omega,$ (74a)
+$\partial^{\bullet}w+w\nabla\cdot{\boldsymbol{v}}-D_{c}\Delta w$ $=f_{2}(u,w)\quad\text{ in }\Omega,$ (74b)
+${\boldsymbol{v}}-\alpha\Delta_{\partial\Omega}{\boldsymbol{v}}-\beta\nabla_{\partial\Omega}\mathbf{Id}_{\partial\Omega}$ $=\delta u{\boldsymbol{n}}_{\partial\Omega}\quad\text{ on }\partial\Omega.$ (74c)
+The non-linear couplings and the parameters are kept the same as in (73) and Table 8.1.
+The initial geometry is a circle of unitary radius and the velocity prescribed by (74c) is extended in the bulk harmonically as described in Section 4.
+Results are presented in Figure 13, where it can be observed how the physics is correctly resolved while maintaining good mesh nodes distribution.
+(a) mcadrSolver 1
+(b) mcadrSolver 2
+Figure 13: Simulation at $t=8$ for the model (74) and the implicitly staggered schemes in List 8.1.
 8.2 Numerical results for coupled Helfrich flow and Cahn-Hilliard euqations
 We proceed now to test the coupling for more complicated systems and geometry evolution to explore the potential of our framework.
-The overdamped limit model for fluid deformable two-component membranes presented in [7, p.A41-33] is considered.
+The overdamped limit model for fluid deformable two-component membranes presented in [Bachini et al., 2023, p.A41-33] is considered.
 The discrete scheme (66) is augmented with a Lagrange multiplier to maintain the inextensibility constraint $\nabla_{\Gamma_{h}^{n}}\cdot{\boldsymbol{v}}_{h}^{n}=0$.
 The modified version reads: Given ${\boldsymbol{y}}^{n}_{h},{\boldsymbol{\kappa}}^{n}_{h}\in[V_{h}(\Gamma_{h}^{n})]^{d}$ find ${\boldsymbol{d}}^{n+1}_{h}\in[V_{h0}(\Gamma_{h}^{n})]^{d}$, ${\boldsymbol{y}}^{n+1}_{h}\in[V_{h}(\Gamma_{h}^{n})]^{d}$ and $\lambda_{h}^{n+1}\in V_{h}(\Gamma_{h}^{n})$ such that
-(73a) $\mathopen{}\left(\frac{{\boldsymbol{d}}_{h}^{n+1}}{\tau}\,,\,{\boldsymbol{\phi}}_{h}\mathclose{}\right)_{\Gamma_{h}^{n}}^{h}-(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{y}}^{n+1}_{h}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}+(\lambda_{h}^{n+1}{\boldsymbol{n}}_{\Gamma_{h}^{n}}\,,\,{\boldsymbol{\phi}}_{h})$
+$\mathopen{}\left(\frac{{\boldsymbol{d}}_{h}^{n+1}}{\tau}\,,\,{\boldsymbol{\phi}}_{h}\mathclose{}\right)_{\Gamma_{h}^{n}}^{h}-(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{y}}^{n+1}_{h}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}+(\lambda_{h}^{n+1}{\boldsymbol{n}}_{\Gamma_{h}^{n}}\,,\,{\boldsymbol{\phi}}_{h})$ (75a)
 $\qquad=~p_{h}({\boldsymbol{y}}_{h}^{n}\,,\,{\boldsymbol{\phi}}_{h}),$
-(73b) $\frac{1}{\gamma_{W}}({\boldsymbol{y}}^{n+1}_{h}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{d}}_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}$
+$\frac{1}{\gamma_{W}}({\boldsymbol{y}}^{n+1}_{h}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{d}}_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}$ (75b)
 $\qquad=-(\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}+(\kappa_{0}{\boldsymbol{n}}_{\Gamma_{h}^{n}}\,,\,{\boldsymbol{\psi}}_{h})_{\Gamma_{h}^{n}}^{h}+({\boldsymbol{\mu}}\,,\,{\boldsymbol{\psi}}_{h})_{\partial\Gamma^{n}},$
-(73c) $(\nabla_{\Gamma_{h}^{n}}\lambda_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}\mu_{h})_{\Gamma_{h}^{n}}+(|{\boldsymbol{\kappa}}_{h}^{n}|^{2}\lambda_{h}^{n+1}\,,\,\mu_{h})_{\Gamma_{h}^{n}}$
+$(\nabla_{\Gamma_{h}^{n}}\lambda_{h}^{n+1}\,,\,\nabla_{\Gamma_{h}^{n}}\mu_{h})_{\Gamma_{h}^{n}}+(|{\boldsymbol{\kappa}}_{h}^{n}|^{2}\lambda_{h}^{n+1}\,,\,\mu_{h})_{\Gamma_{h}^{n}}$ (75c)
 $\qquad=((\nabla_{\Gamma_{h}^{n}}{\boldsymbol{y}}^{n}_{h}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\kappa}}^{n}_{h})_{\Gamma_{h}^{n}}\,,\,\mu_{h})+(p_{h}({\boldsymbol{y}}_{h}^{n}\,,\,{\boldsymbol{\kappa}}_{h}^{n})\,,\,\mu_{h}),$
@@
 $\qquad-\kappa_{0}({\boldsymbol{\kappa}}^{n}_{h}\,,\,(\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})^{T}{\boldsymbol{n}}_{\Gamma_{h}^{n}})_{\Gamma_{h}^{n}}^{h}-\frac{1}{2}(\gamma_{W}|{\boldsymbol{\kappa}}^{n}_{h}-\kappa_{0}{\boldsymbol{n}}_{\Gamma_{h}^{n}}|^{2}\;\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}^{h}$
-(74) $\qquad+(({\boldsymbol{y}}^{n}_{h}\cdot{\boldsymbol{\kappa}}^{n}_{h})\;\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}^{h}+({\boldsymbol{f}}_{h}^{n+1}\,,\,{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}^{h}.$
+$\qquad+(({\boldsymbol{y}}^{n}_{h}\cdot{\boldsymbol{\kappa}}^{n}_{h})\;\mathbb{P}_{\Gamma_{h}^{n}}\,,\,\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}^{h}+({\boldsymbol{f}}_{h}^{n+1}\,,\,{\boldsymbol{\phi}}_{h})_{\Gamma_{h}^{n}}^{h}.$ (76)
 In (76), ${\boldsymbol{f}}_{h}^{n+1}$ is a general right-hand side. The system is composed of the following three steps:
@@
     Solver (75) with the right-hand side (recall (53)):
-    (75) ${\boldsymbol{f}}_{h}^{n+1}=\mathcal{E}_{CH}(u_{h}^{n+1}){\boldsymbol{\kappa}}_{h}^{n}-\sigma\epsilon\nabla_{\Gamma_{h}^{n}}u_{h}^{n+1}\mathbb{H}_{h}^{n}\nabla_{\Gamma_{h}^{n}}u_{h}^{n+1},$
-    where ${\boldsymbol{\omega}}_{\Gamma_{h}^{n}}$ is a certain vertex normal, see [11] for details, and $\mathbb{H}_{h}^{n}=-\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\omega}}_{\Gamma_{h}^{n}}$ is the discrete extended Weingarten map.
+    ${\boldsymbol{f}}_{h}^{n+1}=\mathcal{E}_{CH}(u_{h}^{n+1}){\boldsymbol{\kappa}}_{h}^{n}-\sigma\epsilon\nabla_{\Gamma_{h}^{n}}u_{h}^{n+1}\mathbb{H}_{h}^{n}\nabla_{\Gamma_{h}^{n}}u_{h}^{n+1},$ (77)
+    where ${\boldsymbol{\omega}}_{\Gamma_{h}^{n}}$ is a certain vertex normal, see Barrett et al.
+[2007] for details, and $\mathbb{H}_{h}^{n}=-\nabla_{\Gamma_{h}^{n}}{\boldsymbol{\omega}}_{\Gamma_{h}^{n}}$ is the discrete extended Weingarten map.
   2. 2.
     Solver (33) where the prescribed material velocity is set to
-    (76) ${\boldsymbol{v}}_{h}^{\top}=-\nabla_{\Gamma_{h}^{n}}\lambda_{h}^{n+1},\quad{\boldsymbol{v}}_{h}^{\perp}={\boldsymbol{d}}_{h}^{n+1}/\tau.$
+    ${\boldsymbol{v}}_{h}^{\top}=-\nabla_{\Gamma_{h}^{n}}\lambda_{h}^{n+1},\quad{\boldsymbol{v}}_{h}^{\perp}={\boldsymbol{d}}_{h}^{n+1}/\tau.$ (78)
   3. 3.
@@
 The common initial condition for the Cahn-Hilliard solvers is
-(77) $u_{0}({\boldsymbol{x}})=\cos(\pi x_{1})\cos(\pi x_{2})\cos(\pi x_{3}).$
-The simulation is run with variable elasticity modulus $\gamma_{W}\in\{0.5,~0.1,~0.02\}$ following the parameters in [7].
+$u_{0}({\boldsymbol{x}})=\cos(\pi x_{1})\cos(\pi x_{2})\cos(\pi x_{3}).$ (79)
+The simulation is run with variable elasticity modulus $\gamma_{W}\in\{0.5,~0.1,~0.02\}$ following the parameters in Bachini et al.
+[2023].
 The timestep is varied accordingly with values $\tau\in\{10^{-4},~10^{-4},2.5\cdot 10^{-5}\}$ respectively.
@@
 Results are presented in Figure 14, where it can be seen that the characteristic bulging dynamics is correctly recovered.
+As already noted in 7.2.1 and Figure 10, the pearlings’ mesh appears to be very coarse due to the extreme deformation.
+This is characteristic of the employed mesh redistribution algorithm presented in Section 4.
+As already noted in Garcke et al.
+[2025], alternative choices are possible that might lead to better results Sauer [2014], Hu and Li [2022], Porrmann et al.
+[2025], Sauer [2025], Zhang and Huang [2025], Duan [2025].
+Our setting is flexible enough to allow to switch between various options.
+We repeat the test in 7.2.1 substituting the surface mesh redistribution in Section 4 Duan and Li [2024] with the minimal deformation rate (MDR) redistribution of Hu and Li [2022] presented in (34).
+The results are shown in Figure 15.
+It can be seen how using the MDR technique leads, in this very specific case, to better results with respect to Figure 14.
+This said, we found the performance of the different redistributions to be case-specific.
+For example, one can notice how, despite the enhanced node distribution, the inextensibility constaint for Figure 15(c) is violated at an earlier time with respect to Figure 14(c).
 (b) $\gamma_{W}=0.5,~t=1$
@@
 (d) $\gamma_{W}=0.02,~t=0.2$
-Figure 13: (a) Initial setup for problem the problem 77.
+Figure 14: (a) Initial setup for the problem 79.
+(b)-(d) Mesh for different elasticity modulus.
+The simulation is shown at end time $t=1$ or at last timestep before inextensibility threshold is violated.
+(b) $\gamma_{W}=0.5,~t=1$
+(c) $\gamma_{W}=0.1,~t=0.45$
+(d) $\gamma_{W}=0.02,t=0.2375$
+Figure 15: (a) Initial setup for the problem 79.
 (b)-(d) Mesh for different elasticity modulus.
@@
 This work addresses a long-standing need for robust computational frameworks to tackle such problems.
-While other frameworks exist [85, 6, 7, 83, 13], our work targets flexibility of application while maintaining accuracy.
+While other frameworks exist Mokbel et al.
+[2024], Arroyo and DeSimone [2009], Bachini et al.
+[2023], Mackenzie et al.
+[2021], Barrett et al.
+[2017b], our work targets flexibility of application while maintaining accuracy.
 Algorithms that enjoy stability and convergence properties are tied to non-invasive postprocessing techniques so to favor complex applications.
@@
 The framework also couples recent mesh redistribution techniques for gradient flows with the ALE method.
-We showed that this union is efficient in handling complex reshaping dynamics without the need for remeshing.
+This union is showed to be efficient in handling complex reshaping dynamics without the need for remeshing.
 As a result, the algorithm is tunable to the needs of the user and does not modify the underlying domain evolution.
@@
 The framework developed here has strong potential to simulate problems with increasing biological complexity.
-To achieve this goal, we foresee the need to incorporate large systems of ADR equations, that are not discussed in the present work [50, 83, 82].
-Furthermore, fluid equations have also been shown to play an important role at certain timescales [94, 96, 2].
-Including Navier-Stokes type equations on both interfaces and bulk [85, 7] will broaden the applicability of the computational schemes presented here.
+To achieve this goal, we foresee the need to incorporate large systems of ADR equations, that are not discussed in the present work Francis et al.
+[2024], Mackenzie et al.
+[2021], MacDonald et al.
+[2016].
+Furthermore, fluid equations have also been shown to play an important role at certain timescales Seifert [1997], Shin and Brangwynne [2017], Alberti and Hyman [2021].
+Including Navier-Stokes type equations on both interfaces and bulk Mokbel et al.
+[2024], Bachini et al.
+[2023] will broaden the applicability of the computational schemes presented here.
 In summary, we presented a finite element framework able to handle a rich set of reshaping dynamics arising from bulk-surface coupled PDEs in biophysics.
@@
 Acknowledgments
-We acknowledge Dr. Emmet Francis for his valuable input on biologically relevant models for computational biology and for sharing the lessons learned from his programming experience [50].
+We acknowledge Dr.
+Emmet Francis for his valuable input on biologically relevant models for computational biology and for sharing the lessons learned from his programming experience Francis et al.
+[2024].
+All the simulations are performed using the software NgSolve Schöberl [2014], Gangl et al.
+[2021].
 Declaration of Interests
@@
 The terms of this arrangement have been reviewed and approved by the University of California, San Diego in accordance with its conflict-of-interest policies.
+Funding sources
+P.R. supported by an NIH grant R35 GM158446.